### v1

# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

Abstract
Satellite remote sensing (RS) enables a wide array of downstream Earth observation (EO) applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use. We present TESSERA, a novel Remote Sensing Foundation Model (RSFM) that uses Self-Supervised Learning (SSL) to generate global, robust representations at 10m scale from pixel-level satellite time series data. TESSERA combines information from only optical and SAR data streams using two parallel Transformer-based encoders: one dedicated to Sentinel-1 SAR polarizations and another to Sentinel-2 MSI data (10 selected spectral bands) to create representations that are then fused using a multilayer perceptron (MLP), resulting in a global representation map covering the years 2017 to 2024. Our precomputed representations set a new state-of-the-art performance benchmark and our open-source approach democratizes access to high-performance, high-resolution representations. We benchmark the performance of TESSERA in five diverse tasks, comparing our work with state-of-the-art task-specific models and other foundation models. Our results show that TESSERA outperforms both traditional RS baselines and the leading geospatial foundation models in these diverse downstream tasks.
Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical for addressing global challenges including food security, biodiversity loss, climate change, and disaster mitigation. Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future. However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training. Direct use of EO data also presents challenges as the data quality is affected by cloud cover, atmospheric effects, sensor biases, and non-uniform temporal sampling, preventing full utilization of this valuable resource.
Users of EO data usually mitigate cloudiness by annual or seasonal compositing and then use supervised training of a neural network to map from composited images to tasks-specific classification, such as crop identification. However, compositing removes the critical temporal signal and training a task-specific neural network is computationally onerous. More recent EO approaches tackle the second problem by adopting the paradigm of foundation models, creating general-purpose models pre-trained on a broad set of EO data. For example, Visual Foundation Models (VFMs) have been engineered to interpret a wide array of geospatial data types, including multispectral, hyperspectral, and SAR imagery. Their development follows several paths, including supervised pre-training on large labeled datasets for spatiotemporal analysis [13, 12, 83, 11], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself. These SSL techniques include generative methods, such as masked autoencoders [19, 55, 69, 32, 62, 48, 61, 28, 77, 74, 30, 72, 37, 79] and diffusion-based models [92, 42, 64], and contrastive methods which learn efficient representations by comparing data samples [7, 54, 5, 52, 43, 14, 29, 38, 85]. Subsequently, these general-purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution segmentation [84, 87, 88, 16, 91].
Whilst existing SSL approaches offer rich multi-dimensional ‘representations’ from unlabeled data through auxiliary tasks that reveal latent patterns [8, 51], they still do not address the problem of loss of the temporal signal during compositing. In recent work Lisaius et al [49] have demonstrated that the Barlow Twins method [90] is an elegant approach, grounded in redundancy reduction principles [10], to preserve the temporal signal in latent representations. Building on this insight, we introduce TESSERA, a foundation model for EO data that creates 128-dimensional latent representations at global scale to enable state-of-the-art performance across a diverse array of complex tasks. Our experiments demonstrate that TESSERA-learned representations outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California. Crucially, our model is open source, enabling ease of adoption and reproducibility. Our global representation map also enables users to keep their data private rather than having to upload it to a centralized repository.
1 Our Model
TESSERA processes unlabeled time series from Sentinel-1 SAR and Sentinel-2 Multispectral Instrument (MSI) data. It is a remote sensing foundation model designed for pixel-wise feature extraction from Sentinel-1 and Sentinel-2 imagery. It comprises two main components: a dual-branch encoder and a projector network (see Fig. 1 for a detailed diagram). For each 10-meter pixel globally, it generates a compact, 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics. A primary goal of this work is the production of global, annual, 10-meter resolution representation maps covering the years 2017 to 2024. These pre-computed maps significantly lower the barrier of entry for a wide array of downstream Earth observation applications, by providing readily usable, information-rich features. The core methodology is based on a self-supervised learning paradigm, leveraging the complementary information from optical and SAR data streams.
Figure 1: Architecture and data processing pipeline of the TESSERA foundation model. a, Overview of the end-to-end workflow. Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR are first extracted for each pixel location, forming modality-specific ‘d-pixels’. For self-supervised training, two distinct augmentations are created from each d-pixel and processed through a dual-branch architecture with shared weights within each branch. The resulting 128-dimensional representations from both modalities are fused by a multilayer perceptron MLP. The fused features are then expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn invariant and decorrelated representations. b, Detailed structure of a single encoder branch. A d-pixel, represented as a matrix of temporal steps versus spectral bands (e.g., 40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded. A DOY-based temporal positional encoding is added to these embeddings. The resulting sequence is then processed by an 8-block Transformer Encoder. Finally, a GRU pooling layer aggregates the temporal features to produce a single 128-dimensional representation for that modality, which is then passed to the fusion MLP shown in panel a.
The fundamental input unit for TESSERA is termed the “d-pixel”. For multi-spectral (from Sentinel-2) or SAR backscatter (from Sentinel-1) observations each d-pixel uniquely represents a single geographic 10-meter pixel by structuring the repeat observations from the complete annual time series of the modality into a two-dimensional array (i.e, timesteps by channels). This format explicitly preserves the temporal evolution and spectral/backscatter signatures inherent in the satellite data. The d-pixel design also intrinsically accommodates common observational gaps, such as those due to cloud cover, by masking invalid data points, which are then handled during subsequent processing stages. See Appendix A.4 and Supplementary Fig. 1.
1.1 Dual-encoder
TESSERA employs a multi-modal architecture to process and integrate information from the Sentinel-1 and Sentinel-2 observations. The model consists of specialized encoders for each modality followed by a projection head for self-supervised learning.
The model features two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data (VV and VH polarizations) and another for Sentinel-2 MSI data (10 selected spectral bands). Each encoder processes the temporal sequence of a d-pixel, utilizing multi-head self-attention mechanisms to capture complex temporal dependencies and patterns specific to each data modality. Positional encodings derived from the Day-of-Year (DOY) are incorporated to provide temporal context. An attention-pooling layer within each encoder aggregates the temporal features into a fixed-size vector representing the annual signature for that modality. See Appendix A.7 and Supplementary Figure 1 for architectural details.
1.2 Projector
The modality-specific representations derived from the dual encoders are first fused using a multi-layer perceptron (MLP). This fused representation, which serves as the final 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network. This projector, a significantly larger MLP, expands the dimensionality of the fused representation [^1] This is a critical step inspired by Barlow Twins [90] to facilitate effective redundancy reduction during the self-supervised learning phase.
1.3 Model Training
The TESSERA model learns its representations through a self-supervised pretraining process on a large-scale, unlabeled global dataset. It is trained by optimizing a modified Barlow Twins loss function [90]. The core principle is to learn embeddings that are invariant to differing or incomplete views of the same pixel’s data (e.g., different sets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions. This is achieved by generating two distinct augmented views for each input d-pixel, primarily through independent sparse temporal sampling of annual Sentinel-1 and Sentinel-2 observations, rather than applying artificial distortions or augmentations. The model then learns to reconcile these partial views. The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, augmented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Appendix A.7.2 for the full loss function, Equation 12).
The pretraining phase utilizes a vast global dataset comprising approximately 0.8 billion d-pixels, sampled from Sentinel-1 and Sentinel-2 imagery. A crucial element for successful generalization was a specific multi-level data sampling, shuffling, and batching strategy. This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features. The model was trained for one epoch over this dataset using distributed training techniques.
1.4 Model Inference and Global Representation Maps
Following pretraining, the TESSERA dual-encoder (with frozen weights and excluding the projector) is used to generate the final 128-dimensional representation for every 10-meter pixel globally, for each year from 2017 to 2024. This involves processing the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA representation maps. These maps are designed as readily usable, multi-channel geospatial data layers. This “Model-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, pre-computed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself.
2 Downstream Tasks
The effectiveness and generalizability of the TESSERA representations are rigorously evaluated across a diverse range of downstream remote sensing tasks. These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height estimation), and patch-based dense predictions (e.g., semantic segmentation and land change detection). For all such evaluations, the pretrained TESSERA encoder acts as a fixed feature extractor. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed representations. This approach typically requires substantially less labeled data and computational resources compared to training deep models from scratch, highlighting the transfer learning capabilities of TESSERA. See Appendix A.9 and Supplementary Fig. 9 for the general downstream application workflow.
2.1 Crop Type Classification Evaluation
To validate the effectiveness of TESSERA representations for agriculture, we performed a rigorous evaluation on a crop type classification task. We used the official Austrian INVEKOS dataset for the 2021-2022 growing season [1], which contains 17 aggregated crop classes.
Our evaluation methodology centered on comparing the performance of a lightweight classification head trained on TESSERA representations against two key baselines: (1) a traditional Random Forest classifier trained on engineered temporal features from raw satellite data, and (2) the PRESTO foundation model [79], where we applied the same lightweight head to its embeddings for a direct comparison. We assessed performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labeled data down to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Further details on data processing, model architectures, and training protocols are provided in Section B.
2.2 Crop classification
Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [86]. Crop type classification inform vital agricultural decisions by allowing accurate estimates of crop area and yield, and facilitating low-cost monitoring of diseases and pests over vast regions [31]. However, this task is notably more complex than many other land cover classification challenges, as it requires discerning subtle spectral and temporal variations that differentiate various crop classes [31]. For years, classifiers like Random Forest (RF) applied to hand-crafted composite time-series data have served as a standard baseline [31, 63].
Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific training, offer generalizable representations that can be applied universally [4, 41]. Here, we demonstrate that TESSERA’s pre-computed representations set a new state-of-the-art. We compare TESSERA against two critical benchmarks: the traditional RF baseline and the leading pixel-based foundation model, PRESTO (Pretrained Remote Sensing Transformer) [79], a transformer-based model utilizing multi-modal inputs. Our results show TESSERA not only substantially outperforms both, but also offers a more streamlined and efficient application workflow.
To evaluate performance, we used the INVEKOS Austrian crop dataset from the 2021–2022 growing season, an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on phenology and data availability [1]. TESSERA consistently surpasses both baselines across all training regimes, especially in low-data settings crucial for operational deployment where labeled data is scarce (Fig. 2). In pixel-wise classification tasks with training data splits from 1% to 30%, a simple multilayer perceptron (MLP) trained on TESSERA representations achieves higher average and balanced F1 scores than both RF and an identical MLP trained on PRESTO embeddings, with performance gains frequently exceeding 10% and 30%, respectively (Fig. 2a). TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig. 2b). Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models across various patch sizes (Fig. 2c).
These performance gains are rooted in the superior quality of the learned representations. A UMAP analysis of the embedding space reveals that TESSERA’s representations form tighter, more semantically coherent clusters by crop class compared to those from PRESTO (Fig. 2d). This improved separability highlights TESSERA’s strength as a foundational tool for agricultural applications. Notably, using TESSERA’s compact, pre-computed 128-dimensional representations eliminates the need for the complex data preparation and feature engineering required by traditional methods like RF, democratizing access to high-performance crop classification.
Figure 2: TESSERA representations achieve state-of-the-art performance in crop type classification. The evaluation was conducted on the 2022 Austrian INVEKOS dataset. a, Average and Balanced F1 scores for pixel-wise classification as a function of the training data ratio. TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model trained on raw time-series data and the PRESTO foundation model. b, Overall F1 score in few-shot learning scenarios. TESSERA maintains a significant performance advantage even with very few training samples per class. Error bars represent the standard deviation over multiple runs. c, Patch-based semantic segmentation performance comparison. TESSERA representations, used with a UPerNet head, achieve higher mean Intersection over Union (mIoU) and Balanced F1 scores than other foundation models across various patch sizes and training data ratios. d, UMAP visualization of the embedding spaces for TESSERA and PRESTO for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering, as supported by superior Silhouette and Davies-Bouldin scores, indicating a more semantically meaningful representation space.
2.3 Canopy height estimation
Canopy height is a key structural attribute of forests, closely linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks like REDD+ [57, 26]. It reflects ecosystem function [56], forest age, and successional stage, while also serving as a proxy for habitat quality and vertical complexity—important determinants of biodiversity patterns and species richness [20, 15, 73]. Accurate, scalable canopy height models (CHMs) enable detection of forest degradation and disturbance arising from selective logging, disease outbreaks and storm damage.
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. GEDI, a spaceborne LiDAR system, provides spatially sparse measurements of canopy height that are used to train models for generating maps from optical and radar data; however, GEDI does not sample high latitudes [21], and produces unreliable estimates in mountainous terrain [27]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to isolate canopy height accurately. Optical data, while widely available, lacks inherent sensitivity to vertical structure and requires indirect proxies or fusion with information from active sensors  [82, 50] - yet none of the three most widely used global CHM models currently implement sensor fusion [46, 68, 78]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., $>500$ features in Potapov et al. [68]), which may generalize poorly across ecosystems. These models tend to under-predict tall canopies, especially in tropical forests where the majority of above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [45], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, with current products underestimating local variability and structural extremes [58]. Most CHMs are also static, failing to capture seasonal or interannual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [59, 65].
We evaluated TESSERA representations to predict the height of the airborne LiDAR-derived canopy at 10 m resolution within a 5 × 6 km area of tall old-growth tropical forest in the Danum Valley, Borneo[40]. To ensure spatial independence, we conducted four-fold spatial cross-validation, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 64 × 64 pixel patches of the representations as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. Accuracy metrics for the foundation models are reported as the average across all 12 runs. TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of $-0.62$ m.
We benchmarked TESSERA’s performance against state-of-the-art approaches through three comparisons (Fig. 3). First, we compared predictions against three global canopy height products: Potapov et al. [68] (‘GLAD’ map), Lang et al. [46] (‘ETH’ map), and Tolan et al. [78] (‘Meta’ map). Although local models generally outperform global products in regional evaluations [70], the pronounced saturation of these global models at much lower canopy heights—combined with RMSEs exceeding their reported global averages [68, 46, 78]—underscores the challenge of accurately mapping structurally complex, tall-canopy regions such as Danum Valley. Second, we compared TESSERA to the PRESTO [79] foundation model, using the same model architecture and training procedure but replacing the inputs of TESSERA with the representations of PRESTO. Third, we compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines generated by Lang et al. [47], which used deep convolutional neural networks trained on Sentinel-2 imagery and GEDI LiDAR reference data for regional canopy height and biomass estimation.
In all comparisons, TESSERA outperformed competing methods. Moreover, its representations do not require preprocessing, greatly enhancing reproducibility. Instead, only the region and year of interest are needed to replicate the input data pipeline. By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional representations encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
Figure 3: TESSERA representations outperform global and regional models in canopy height estimation. Top row: Density scatter plots compare predicted canopy heights from the TESSERA (left) and PRESTO (middle) models against airborne LiDAR-derived heights. TESSERA achieves higher accuracy than PRESTO, demonstrating superior performance in capturing structural forest attributes. The rightmost panel shows model bias across 10m canopy height bins, with LiDAR values binned on the x-axis and model predictions on the y-axis. TESSERA more closely follows the 1:1 line, particularly in mid- to high-canopy regimes, while PRESTO underestimates heights above  40m. Histogram above indicates pixel count distribution per bin. Bottom row: Comparison with three global canopy height products— ETH Global CHM, GLAD (Potapov) Global CHM, Meta (Tolan) Global CHM, and ETH Regional CHM—evaluated on the same test region. All four products exhibit substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14m$), including negative correlations for three out of four models. These results underscore the difficulty of accurately estimating tall tropical canopies using global models trained without local data. Rightmost column: Spatial comparison between true canopy height (LiDAR) and TESSERA-predicted canopy height for a representative test patch, illustrating the model’s capacity to capture fine-scale structural variation while still saturating for taller canopy values.
2.4 Burned Area Detection
Wildfire frequency and severity are projected to increase in many parts of the world due to climate change [3, 39, 66]. Accurate maps of burned areas (BAs) and burn severity are critical for monitoring wildfire trends [6], predicting wildfire occurrence [17], assessing ecosystem recovery [89], and establishing effective fire management strategies [33]. Remote sensing is widely used to detect burned areas and assess severity regionally and globally [44].
Most existing BA mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [44, 67]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), though many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [44]. The majority of existing methods use optical sensors [44], which can detect fire-specific spectral signatures but are susceptible to cloud cover [34], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [35]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [44, 76, 81]. Recent reviews emphasize the need for sensor fusion, combining passive and active sensors to improve accuracy and robustness [44, 18]. Sentinel sensors in particular, when combined together at 10 m resolution, have shown promise in detecting BA from small fires that are often left out of coarse, global BA products [76, 81, 71]. TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of abstract, data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig. 4).
Figure 4: TESSERA representations capture wildfire extent, timing, and severity across two burned regions. (a–c) Burn Area A, which experienced two distinct fires in June and August 2021. (a) Time of fire map based on MTBS, showing unburned areas (blue), the June fire (orange), and the August fire (red). (b) Burn severity classification for the same region, with unburned (blue), low (green), moderate (orange), and high (red) severity. (c) UMAP projection of TESSERA embeddings from 2021, with colors corresponding to burn severity. Embeddings form well-separated clusters for unburned, June-, and August-burned areas, with a visible internal gradient reflecting severity within August fire. (d–f) Burn Area B, which experienced a single large fire in September 2021. (d) Time of fire map showing unburned (blue) and burned (green) areas. (e) Burn severity classification showing heterogeneous severity across the landscape. (f) UMAP projection for Burn Area B, revealing fuzzier boundaries between burned and unburned pixels, though the severity gradient remains evident.
We tested whether TESSERA’s temporal representations capture fire-induced changes in spectral and radar signals. We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 × 15 km regions from the Burn Scar Benchmarking dataset, each affected by wildfires in California during 2021. TESSERA representations for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels. Fire boundaries and burn severity data were sourced from the Monitoring Trends in Burn Severity (MTBS) dataset[2]. The resulting UMAP projections demonstrated strong separability in two-dimensional space, visually highlighting TESSERA’s capacity to detect disturbances at the pixel level. Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single, large fire affecting mostly forested areas, resulting in heterogeneous burn severity. The UMAP revealed separability across three domains (Fig. 4):
  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned ones, with clearer separation in Burn Area A.
  • Temporal fire differentiation: In Burn Area A, embeddings separated pixels burned in June vs. August, suggesting temporal sensitivity to disturbance events.
  • Burn severity gradient: In both regions, a gradient structure in UMAP space corresponded to burn severity, indicating that TESSERA encodes disturbance magnitude.
Because TESSERA encodes annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images. This capacity would support a simplified, scalable pipeline for large-area disturbance monitoring.
2.5 Above-ground biomass estimation
Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [36]. Accurate estimation of above-ground biomass (AGB) is essential for quantifying forest carbon stocks and monitoring temporal changes due to deforestation, degradation, and recovery [22].
We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset. Firstly, we considered the geospatial foundation model SpectralGPT [37], which was found in the Pangaea benchmark [53] to outperform other tested foundation models on the BioMassters dataset [60]. Secondly, as the state-of-the-art bespoke model, we considered the winner of the Biomassters competition, a UNet with a temporal attention encoder [60]. The BioMassters dataset comprises 11,462 image patches representing above-ground biomass across Finnish forests from 2017 to 2021. Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey. Each patch covers a 2560×2560 m area, with 10×10 m spatial resolution per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
A comparison of RMSE across varying training label fractions (Fig. 5a) shows that a UNet trained on the TESSERA representations consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set. These results underscore the robustness of TESSERA in low-label regimes. Moreover, while the best task-specific model trained on the full dataset outperforms TESSERA, the gap is moderate, even with low label availability. In general, bespoke models are inherently difficult to surpass in fully supervised settings. It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
Figure 5: TESSERA representations demonstrate robustness under limited label availability. a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGBT (RMSE values from Ref. [53]). Following the evaluation protocol of Ref. [53], RMSE is computed on a fixed test set of 2773 patches, with the remaining data randomly split into training and validation sets (80:20). The UNet model is trained on the TESSERA representations for 80 epochs, with the checkpoint achieving the lowest validation loss selected as the final model. For limited labels testing, a subset of labels is randomly chosen from the training set. The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability. b, Predicted AGB map for an example patch using 10% of the training labels. c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset. d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top) and Biomassters winner (bottom).
2.6 Stocking indices in voluntary carbon markets
The most widely applied standard in the voluntary carbon market (VCM) for reforestation and afforestation projects of degraded pastures requires project proponents to demonstrate that their chosen remotely-sensed “stocking index” is correlated with the AGB (REF). Upon acceptance, project developers can use the index to quantify project additionality and to assess project performance in-between the 5-yearly ground-truth assessments. There is considerable freedom in choosing a stocking index, which can be either a pre-existing or proprietary canopy height or AGB product. Here, we evaluate TESSERA’s performance against five widely used, globally available canopy height and AGB products often used in the VCM. The in situ data used to compare the different stocking indices was collected by ICRAF in 38 sites in Para State, Brazil, that were degraded pastureland until being converted into agroforestry systems during 1980-2010 (Atzberger et al., 2025). To provide a fair comparison, stocking indices were rescaled using a linear transformation to minimize the RMSE with respect to the in situ data (details in the SI).
Figure 6: TESSERA outperforms commonly used stocking indices in the Voluntary Carbon Market. Comparison of in-situ aboveground biomass measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products. Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using X * 15.502 + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA: AGB converted using 0.407 X + 34; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X above ground biomass prediction converted using 0.3X + 54.3. Details of conversation functions provided in SI
3 Conclusion
In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that achieves worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level satellite time series data. The representations generated by TESSERA establish new performance standards, while our open source methodology ensures widespread access to superior, high-resolution representations. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. The findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
We next provide supplementary materials and methods about the details of how TESSERA is implemented.
Appendix A Earth Observation data
Our input data consists of Sentinel-2 optical satellite imagery (Level-2A bottom-of-atmosphere) and Sentinel-1 Synthetic Aperture Radar (SAR, Radiometrically Terrain Corrected).
A.1 Data Representation
We consider remote sensing data with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
$R_{t}\in\mathbb{R}^{W\times H\times C}$ (1)
where $W$ is the width (longitude dimension), $H$ is the height (latitude dimension), and $C$ is the number of spectral channels.
Each tile is accompanied by a corresponding binary mask $V_{t}$ of dimensions:
$V_{t}\in\{0,1\}^{W\times H}$ (2)
where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
A.2 Temporal Data Stacking
We stack spatially aligned tiles over a time period spanning $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (3)
$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (4)
A.3 Time Series Extraction and d-pixel Definition
For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (5)
We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
$P_{i,j}=S(i,j)$ (6)
In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time. Note that d-pixels are potentially sparse and have an accompanying mask $m_{i,j}$ that indicates the timesteps in which there is valid data:
$m_{i,j}=[v_{i,j,0},v_{i,j,1},\ldots,v_{i,j,T-1}]$ (7)
Variable Dimensions Description
$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
$V_{t}$ $W\times H$ Binary mask for tile at time $t$ (1=valid, 0=invalid/clouded)
$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
$W,H$ scalar Spatial dimensions (width, height)
$C$ scalar Number of spectral channels
$T$ scalar Number of time steps
Table 1: Variable definitions for remote sensing data structure
A.4 The d-pixel Representation
We represent the rich information contained within satellite image time series at the level of individual pixels using the aforementioned “d-pixel”. For a single 10-meter geographic pixel location, it consolidates its complete annual multi-spectral (from Sentinel-2) and SAR backscatter (from Sentinel-1) time series. This structure explicitly preserves the inherent spectral and temporal dimensions of the data (Supplementary Fig. 1 a illustrates the d-pixel concept).
In this array, each row corresponds to a distinct observation date throughout the year, ordered chronologically. Each column corresponds to a specific spectral band or polarization. The value at the intersection of a given row and column thus represents the measured reflectance or backscatter for that specific band/polarization on that particular date for that single pixel.
A critical feature of the d-pixel is its capacity to accommodate the frequent data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked within the d-pixel array. These missing entries are explicitly handled by downstream model components, for instance, by being ignored during attention computation or managed during the temporal sampling process described in Appendix A.7.3.
A.5 Dual-Encoder Architecture
Given the distinct nature of Sentinel-1 SAR and Sentinel-2 MSI data, TESSERA employs two separate, parallel Transformer-based encoder branches.
  • Sentinel-2 MSI Encoder: This branch processes time series of 10 spectral bands from Sentinel-2. We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
  • Sentinel-1 SAR Encoder: This branch processes time series of 2 polarizations from Sentinel-1 (VV and VH).
Each encoder begins by linearly embedding the input features (spectral bands or polarizations) for each time step. To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings. The core of each encoder consists of a stack of 8 standard Transformer blocks [80], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
To derive a single vector summarizing the entire time series for each modality, an attention-pooling layer weighs the importance of different time steps before aggregation. The resulting modality-specific representations (one from the S1 encoder, one from the S2 encoder) are then fused using a multi-layer perceptron.
A.6 Projector Network
The fused representation from the dual-encoder stage is subsequently fed into a large projector network. This projector is a five-layer MLP, with each hidden layer having 16,384 dimensions. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [90], to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector for each input d-pixel is an embedding, which is then used in the loss calculation. For downstream tasks, we typically use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel representation. The TESSERA encoder (up to the fusion MLP) has approximately 40 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
A.7 Self-Supervised Training
A.7.1 Augmented View Generation
The TESSERA model is trained using a modified Barlow Twins objective function [90]. For this objective, two distorted views, denoted as $Y_{A}$ and $Y_{B}$, are generated for each input d-pixel. In TESSERA, these views are created by independently running the temporal sampling and preprocessing pipeline twice for the Sentinel-1 and Sentinel-2 data associated with a given d-pixel. This process involves:
  1. 1.
    For each view, independently sampling a fixed number of valid observation dates from the annual Sentinel-2 time series (10 spectral bands).
  2. 2.
    For each view, independently sampling a fixed number of valid observation dates from the annual Sentinel-1 time series (2 polarizations).
These views represent different, valid, but inherently incomplete glimpses of the pixel’s true temporal-spectral evolution, akin to observing the same location through intermittent cloud cover or from different satellite passes at different times. The model learns by reconciling these partial views. The inherent differences between the Sentinel-1 SAR and Sentinel-2 MSI modalities further provide diverse perspectives on the same underlying physical processes. Thus, our augmentations are fundamentally about sampling from the available, inherently incomplete information streams, rather than artificially distorting a complete input.
A.7.2 Loss Function
The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as [90]:
$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (8)
Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$. The indices $i$ and $j$ iterate over the dimensions of the embedding vectors. The first term (invariance term) encourages similar representations for different views of the same input ($C_{ii}\to 1$). The second term (redundancy reduction term) promotes informational efficiency by minimizing correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
To further enhance model robustness and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term, $\mathcal{L}_{MIX}$, inspired by Bandara et al. [9]. This involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating mixed views $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$, where $\alpha_{mix}\sim\text{Beta}(\beta_{p},\beta_{p})$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained. The mix-up loss penalizes deviations from the assumption that a linear interpolation in input space corresponds to a linear interpolation in embedding space:
$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (9)
$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (10)
$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (11)
where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model’s outputs. The total loss function optimized during the training of TESSERA is a weighted sum:
$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (12)
where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{BT}=5\times 10^{-3}$ and $\lambda_{mix}=1.0$ to be effective.
A.7.3 Pretraining Details
The TESSERA model (Alpha version) was pretrained using approximately 0.8 billion d-pixel samples derived from 3,012 globally distributed Military Grid Reference System (MGRS) tiles. For pretraining, these d-pixels were generated from Sentinel-1 and Sentinel-2 data that were spatially downsampled by a factor of 400 from their native 10-meter resolution. Each d-pixel for Sentinel-2 contained time series for 10 spectral bands, and for Sentinel-1, 2 polarizations (VV and VH). The sequence length for the Transformer encoders was fixed at 40 timesteps for both modalities.
For each pixel location, after filtering invalid observations (e.g., due to cloud cover for Sentinel-2), we performed sparse temporal sampling. This involves randomly selecting a fixed number of 40 valid observation dates from the year’s data. If fewer than 40 valid dates exist, sampling is performed with replacement. This strategy standardizes the input sequence length and serves as a key data augmentation mechanism, building invariance to data gaps and teaching the model that the underlying signal persists regardless of the specific dates observed. The temporal context of each sampled observation was encoded by transforming its normalized Day-of-Year (DOY) into sine and cosine features, which were then concatenated with the corresponding spectral or backscatter measurements. Finally, these values were standardized using global statistics to stabilize training.
Figure 7: Geographical distribution and data density of the training dataset. TESSERA was trained on over 3,000 MGRS tiles distributed globally from 2017 to 2024. The color of each hexagon in the map corresponds to the number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate color legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
The model was trained for 1 epoch over the entire dataset, which corresponded to approximately 3,000 GPU hours on 8 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. The AdamW optimizer was employed with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule included a linear warmup over the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay. The global batch size was 8192.
A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a vast and geographically diverse dataset. Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch. To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$0.8 billion training samples, which constituted over 2TB of initial d-pixel data. This pipeline is illustrated conceptually in Fig. 8a.
The impact of this approach on training stability is empirically demonstrated in Fig. 8b. Compared to a conventional, localized shuffling strategy which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
Figure 8: Overview of the data shuffling pipeline and its impact on training stability. a, Schematic of the data shuffling and loading process. D-pixels (colored squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
Operationally, the process begins with the aggregation of d-pixels from all MGRS tiles into a single, comprehensive pool. A global shuffling operation is performed on this pool, a critical step to break the spatial contiguity of data from individual tiles and ensure each training batch contains a diverse mix of geographical and environmental contexts. Following this global shuffle, the data augmentation required by the Barlow Twins framework is applied. As detailed in Appendix 5.1, this involves generating two distinct augmented views (e.g., $Y^{A}$ and $Y^{B}$) for each d-pixel.
To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed the pipeline as a custom Rust binary. This high-performance binary handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation. The resulting pairs of augmented d-pixels are then serialized into a compact, pickle-like file format. These files are organized into manageable chunks and loaded by PyTorch DataLoader workers, which stream the data and assemble the final training batches. This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
A.8 Global Representation Map Generation
A primary output of the TESSERA project is the generation of annual global representation maps at 10-meter resolution for the years 2017-2024.
A.8.1 Model Inference
To generate these maps, the pretrained and frozen TESSERA dual-encoder (excluding the projector) is used. For each 10-meter pixel on the globe and for each year:
  1. 1.
    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pretraining, no spatial downsampling is performed at this stage.
  2. 2.
    A fixed number of 40 timesteps are sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
  3. 3.
    These sampled time series are fed into their respective frozen TESSERA encoders.
  4. 4.
    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional representation vector for that pixel for that year.
This process is repeated for all land pixels globally to create an annual representation map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
A.8.2 Data Product and Accessibility
The resulting product consists of eight such global representation maps, one for each year from 2017 to 2024. These maps are intended to be released publicly, alongside the open-sourced TESSERA model parameters and generation code. This “Model-as-Data” approach significantly lowers the entry barrier for users, who can treat these representation maps as conventional multi-channel images. They can be directly ingested by downstream models without requiring users to process raw satellite data or run the TESSERA model themselves, thus democratizing access to advanced RSFM capabilities.
A.9 Downstream Task Application Methodology
A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature representations that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data. Having pretrained TESSERA on large unlabeled datasets, we evaluated the utility of its learned representations across different remote sensing applications. The evaluation methodology involves using the pretrained TESSERA encoders as fixed feature extractors.
The pipeline for applying TESSERA representations to downstream tasks is as follows (illustrated in Supplementary Fig. 9):
Figure 9: Illustration of RSFM training and downstream task applications. (b) shows the training process of a typical RSFM, where both the encoder and the projector (or decoder) are updated. (a) depicts a common application method: the RSFM encoder is extracted and combined with a task-specific head to form a new model for downstream tasks. Users can then fine-tune the entire model or just the head. (c) presents the emerging paradigm championed by TESSERA, where pre-generated representations from the RSFM encoder are used directly as input. This results in extremely lightweight downstream models that consist only of a task-specific head.
  1. 1.
    Load Pretrained Encoder: The weights from the saved pretraining checkpoint are loaded into the TESSERA dual-encoder architecture. The parameters of these encoders are frozen and are not updated during downstream model training. This ensures that downstream performance directly reflects the quality of the fixed representations.
  2. 2.
    Prepare Labeled Downstream Data: The specific labeled dataset for the target task (e.g., pixel-level crop type labels, canopy height measurements, or land change polygons) is prepared.
  3. 3.
    Extract Representations: For each input sample (e.g., pixel, object, or patch) in the labeled dataset, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year undergo the same preprocessing and d-pixel creation pipeline used during TESSERA inference (i.e., using 10m resolution data and temporal sampling to 40 timesteps). The preprocessed sequences are passed through the respective frozen encoders, and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA representation for that sample. To enhance stability, this extraction can be repeated multiple times with different random temporal samplings, and the resulting representations averaged, although for many tasks, a single extraction is sufficient.
  4. 4.
    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA representations as input.
    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (e.g., 1-3 layers) ending in a softmax output layer.
    • For pixel-wise regression (e.g., canopy height regression), the head is usually an MLP ending in a single linear output neuron.
    • For tasks requiring spatial context from the representations (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA representations (e.g., $64\times 64\times 128$). The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions. For land change detection, a simple approach involves computing the dot product between representations from two different years.
  5. 5.
    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA representations as input features and the corresponding labels. Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used. This training typically requires significantly less labeled data and computational power compared to training a deep model from scratch.
  6. 6.
    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA representations for the test samples and passing them through the trained head. Performance is evaluated using standard metrics relevant to the task.
This standardized workflow allows for robust assessment of TESSERA representations across diverse applications, demonstrating their value as foundational features for geospatial analysis.
Appendix B Downstream Task: Austrian Crop Classification
The following section details the experimental setup for the crop classification task presented in the main text.
B.1 Dataset and Preprocessing
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [1]. The dataset originally contained 154 crop types, which we grouped into 17 broader classes (e.g., merging different varieties of wheat) based on phenological similarity and sample availability to ensure robust training and evaluation. For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA representation from our generated 2022 global representation map.
B.2 Pixel-wise Classification Baselines
To provide a comprehensive performance comparison, we implemented three distinct models for pixel-wise classification:
  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA representations were used as input to a simple MLP. The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
  • PRESTO + MLP: For a direct and fair comparison with the closest foundation model, we used the official pre-trained PRESTO model [79] to generate its pixel embeddings. These embeddings were then fed into an identical MLP head as the one used for TESSERA.
  • Random Forest: As a traditional baseline, we trained a Random Forest classifier directly on raw time-series data. For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year. The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector. The RF model consisted of 200 trees, with other parameters set to standard values.
For the experiments shown in the main text’s figure, we trained these models on randomly selected subsets of the data (from 1% to 30% for panel a; specified samples-per-class for panel b), using a fixed validation set for hyperparameter tuning and a held-out test set for final evaluation.
B.3 Patch-wise Semantic Segmentation
To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
  • For pixel-based foundation models like TESSERA and PRESTO, which do not explicitly model spatial context in their representations, we first constructed representation patches (e.g., of size $64\times 64\times 128$) from the pixel-wise embeddings. These patches were then fed into a standard UNet architecture, which learns to model the spatial relationships between the representations to produce a segmentation map.
  • For other foundation models that are inherently patch-based (e.g., Prithvi [75], Satlas [11]), their encoders already process image patches. Therefore, for these models, we only needed to attach and train a UPerNet decoder head to their frozen encoders to generate the final segmentation outputs.
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores.
B.4 Embedding Space Analysis
The 2D visualizations shown in the main text were generated by applying the Uniform Manifold Approximation and Projection (UMAP) algorithm to the representations of every pixel within the Austrian study area. Specifically, the entire representation map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as direct input to UMAP. This was performed for both TESSERA and PRESTO representations.
To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
  • The Silhouette score, $s(i)$, for a single data point $i$ measures how similar it is to its own cluster compared to other clusters. It is defined as:
    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (13)
    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighboring cluster. The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
  • The Davies-Bouldin Index (DBI) evaluates clustering quality by computing the ratio of within-cluster scatter to between-cluster separation. For a set of $k$ clusters, it is defined as:
    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (14)
    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$. Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where clusters are compact and well-separated.
Acknowledgments
We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
Funding:
Z. F. and S. K were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge. S. J was funded by a charitable donation from John Bernstein. J. B received support from the UKRI CRCRM scheme (grant MR/Y024354). S. S. received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
Author contributions:
System design: CA, ZF, SJ, SK
System implementation: ZF, RY,
Data curation: ZF, RY, JK, SJ, AM, DAC, SK
Downstream tasks: ML, CA, SS, DAC, CA, MI, ZF, JK
Paper conceptualization: SK, AM, DAC, AB, SJ
Paper draft: ZF, ML, RY, SS, SK, JB, JK
Paper review and refinement: AB, AM, DAC, SJ, SK
Competing interests:
There are no competing interests to declare.
Data and materials availability:
Our framework integrates data from the Copernicus Sentinel-2 (S2) [24] optical and Sentinel-1 (S1) [23] SAR missions, leveraging the strengths of each sensor. We utilized S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands  [^2] We used red, blue, green, NIR, NIR8, red-edge1, red-edge2, red-edge3, SWIR16, and SWIR22.. A challenge with optical data is frequent cloud contamination; we filtered the time series using the associated Sen2Cor scene classification map (SCL) to mask cloudy observations. For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [25].

### v2

@@
 Abstract
-Satellite remote sensing (RS) enables a wide array of downstream Earth observation (EO) applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use.
-We present TESSERA, a novel Remote Sensing Foundation Model (RSFM) that uses Self-Supervised Learning (SSL) to generate global, robust representations at 10m scale from pixel-level satellite time series data.
-TESSERA combines information from only optical and SAR data streams using two parallel Transformer-based encoders: one dedicated to Sentinel-1 SAR polarizations and another to Sentinel-2 MSI data (10 selected spectral bands) to create representations that are then fused using a multilayer perceptron (MLP), resulting in a global representation map covering the years 2017 to 2024.
-Our precomputed representations set a new state-of-the-art performance benchmark and our open-source approach democratizes access to high-performance, high-resolution representations.
-We benchmark the performance of TESSERA in five diverse tasks, comparing our work with state-of-the-art task-specific models and other foundation models.
-Our results show that TESSERA outperforms both traditional RS baselines and the leading geospatial foundation models in these diverse downstream tasks.
-Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical for addressing global challenges including food security, biodiversity loss, climate change, and disaster mitigation.
-Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future.
-However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training.
-Direct use of EO data also presents challenges as the data quality is affected by cloud cover, atmospheric effects, sensor biases, and non-uniform temporal sampling, preventing full utilization of this valuable resource.
-Users of EO data usually mitigate cloudiness by annual or seasonal compositing and then use supervised training of a neural network to map from composited images to tasks-specific classification, such as crop identification.
-However, compositing removes the critical temporal signal and training a task-specific neural network is computationally onerous.
-More recent EO approaches tackle the second problem by adopting the paradigm of foundation models, creating general-purpose models pre-trained on a broad set of EO data.
-For example, Visual Foundation Models (VFMs) have been engineered to interpret a wide array of geospatial data types, including multispectral, hyperspectral, and SAR imagery.
-Their development follows several paths, including supervised pre-training on large labeled datasets for spatiotemporal analysis [13, 12, 83, 11], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
-These SSL techniques include generative methods, such as masked autoencoders [19, 55, 69, 32, 62, 48, 61, 28, 77, 74, 30, 72, 37, 79] and diffusion-based models [92, 42, 64], and contrastive methods which learn efficient representations by comparing data samples [7, 54, 5, 52, 43, 14, 29, 38, 85].
-Subsequently, these general-purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution segmentation [84, 87, 88, 16, 91].
-Whilst existing SSL approaches offer rich multi-dimensional ‘representations’ from unlabeled data through auxiliary tasks that reveal latent patterns [8, 51], they still do not address the problem of loss of the temporal signal during compositing.
-In recent work Lisaius et al [49] have demonstrated that the Barlow Twins method [90] is an elegant approach, grounded in redundancy reduction principles [10], to preserve the temporal signal in latent representations.
-Building on this insight, we introduce TESSERA, a foundation model for EO data that creates 128-dimensional latent representations at global scale to enable state-of-the-art performance across a diverse array of complex tasks.
-Our experiments demonstrate that TESSERA-learned representations outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California.
-Crucially, our model is open source, enabling ease of adoption and reproducibility.
-Our global representation map also enables users to keep their data private rather than having to upload it to a centralized repository.
-1 Our Model
-TESSERA processes unlabeled time series from Sentinel-1 SAR and Sentinel-2 Multispectral Instrument (MSI) data.
-It is a remote sensing foundation model designed for pixel-wise feature extraction from Sentinel-1 and Sentinel-2 imagery.
+Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use.
+However, satellite time series are voluminous, often corrupted by sensor noise, clouds, and atmospheric conditions, and unevenly spaced in time, making them challenging to use.
+We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time series data.
+TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-60 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps.
+We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines.
+We believe that TESSERA’s ease of use, openness, computation-, label-, and data-efficiency, and high performance will prove transformative in a wide range of ecological applications.
+Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [118, 183, 24].
+Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future [176].
+However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training [95].
+The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and nonuniform temporal sampling, preventing full and easy use of this valuable resource [193, 60, 130].
+EO data users usually mitigate these issues by annual or seasonal compositing [34], then using supervised machine learning to map from composited images to task-specific classification, such as crop identification [8].
+However, compositing removes the critical temporal signal arising from phenology and training a task-specific ML model is computationally onerous.
+Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data [178].
+Their development follows several paths, including supervised pre-training on large labeled datasets for spatio-temporal analysis [21, 20, 160, 18], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
+These SSL techniques include generative methods, such as masked autoencoders [32, 104, 126, 61, 113, 94, 112, 53, 144, 140, 134, 67, 149], diffusion-based models [192, 80, 115], and contrastive methods that learn efficient representations by comparing data samples [14, 102, 7, 99, 81, 23, 59, 70, 169].
+Subsequently, these general purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution (semantic) segmentation [161, 177, 184, 28, 191].
+Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabeled data through auxiliary tasks that reveal latent patterns [15, 97], they still do not address the loss of the temporal signal during compositing.
+In recent work, Lisaius et al [95] have demonstrated that the Barlow Twins loss [187], grounded in redundancy reduction principles [17], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series.
+Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks.
+Our experiments demonstrate that these TESSERA-derived embeddings outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California.
+Importantly, our outputs are ‘ready to use’: Users need neither to fine-tune the model, nor download and process voluminous satellite data, nor perform inferencing other than what is needed to perform a downstream task.
+This makes TESSERA frugal in terms of both data handline and computation.
+They can also keep their data private instead of having to upload it to a centralized repository.
+This removes a crucial barrier to use.
+We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training are needed to achieve state-of-the-art performance.
+Moreover, our model is open source, enabling extensibility and scientific reproducibility.
+We have made a gap-free, global, annual, 10-meter resolution embedding map for 2024 openly available through the Geotessera Python library [152] and aim to eventually provide global maps for the entire period 2017-2024.
+These information-rich precomputed maps greatly simplify a wide range of downstream EO applications.
+We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see [1] for details).
+To our knowledge, TESSERA is unprecedented in its ease of use, scale, and accuracy: no other foundation model provides analysis-ready outputs, is open, and provides global, annual coverage at 10m resolution using only spectral-temporal features at pixel level (Supplementary Text:  LABEL:tab:sup_comparison).
+1 Overview of the TESSERA System
+TESSERA uses self-supervised learning to create per-pixel embeddings from the unlabeled time-series of observations from Sentinel-1 (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams.
 It comprises two main components: a dual-branch encoder and a projector network (see Fig.
-1 for a detailed diagram).
-For each 10-meter pixel globally, it generates a compact, 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics.
-A primary goal of this work is the production of global, annual, 10-meter resolution representation maps covering the years 2017 to 2024.
-These pre-computed maps significantly lower the barrier of entry for a wide array of downstream Earth observation applications, by providing readily usable, information-rich features.
-The core methodology is based on a self-supervised learning paradigm, leveraging the complementary information from optical and SAR data streams.
+1 for details).
+TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps).
+This format preserves the temporal evolution of spectral/backscatter signatures inherent in the satellite data.
+Our design accommodates common observational gaps and irregularities, such as those due to cloud cover or orbit overlap, by masking invalid rows, which are handled during subsequent processing stages.
+Model training and inferencing is done by sparse temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig.
+1 for details).
+For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics.
+Under typical mid-latitude observation conditions, this compresses input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
 Figure 1: Architecture and data processing pipeline of the TESSERA foundation model.
@@
 Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR are first extracted for each pixel location, forming modality-specific ‘d-pixels’.
-For self-supervised training, two distinct augmentations are created from each d-pixel and processed through a dual-branch architecture with shared weights within each branch.
-The resulting 128-dimensional representations from both modalities are fused by a multilayer perceptron MLP.
+For self-supervised training, two distinct augmentations are created from each d-pixel–each augmentation is a random cloud-free sample from an annual time series of Sentinel-2 observations along with the corresponding Sentinel-1 data–processed through a dual-branch architecture with weights shared between the branches.
+The resulting 128-dimensional embeddings from both modalities are fused by a multilayer perceptron (MLP) to create the output embeddings of the system.
 The fused features are then expanded to 16,384 dimensions by a large projector network.
-A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn invariant and decorrelated representations.
+A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn augmentation-invariant and decorrelated embeddings that are internally decorrelated, and fairly invariant to obscured days in the Sentinel-2 time series.
 b, Detailed structure of a single encoder branch.
-A d-pixel, represented as a matrix of temporal steps versus spectral bands (e.g., 40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded.
-A DOY-based temporal positional encoding is added to these embeddings.
-The resulting sequence is then processed by an 8-block Transformer Encoder.
-Finally, a GRU pooling layer aggregates the temporal features to produce a single 128-dimensional representation for that modality, which is then passed to the fusion MLP shown in panel a.
-The fundamental input unit for TESSERA is termed the “d-pixel”.
-For multi-spectral (from Sentinel-2) or SAR backscatter (from Sentinel-1) observations each d-pixel uniquely represents a single geographic 10-meter pixel by structuring the repeat observations from the complete annual time series of the modality into a two-dimensional array (i.e, timesteps by channels).
-This format explicitly preserves the temporal evolution and spectral/backscatter signatures inherent in the satellite data.
-The d-pixel design also intrinsically accommodates common observational gaps, such as those due to cloud cover, by masking invalid data points, which are then handled during subsequent processing stages.
-See Appendix A.4 and Supplementary Fig.
-1.
-1.1 Dual-encoder
-TESSERA employs a multi-modal architecture to process and integrate information from the Sentinel-1 and Sentinel-2 observations.
-The model consists of specialized encoders for each modality followed by a projection head for self-supervised learning.
-The model features two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data (VV and VH polarizations) and another for Sentinel-2 MSI data (10 selected spectral bands).
+A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded.
+A Day-of-Year-based temporal encoding is added to these embeddings.
+The resulting sequence is then processed by an 4-block Transformer Encoder.
+Finally, a Gated Recurrent Unit (GRU) pooling layer processes the sequence of temporal features; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
+TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared).
 Each encoder processes the temporal sequence of a d-pixel, utilizing multi-head self-attention mechanisms to capture complex temporal dependencies and patterns specific to each data modality.
 Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context.
-An attention-pooling layer within each encoder aggregates the temporal features into a fixed-size vector representing the annual signature for that modality.
-See Appendix A.7 and Supplementary Figure 1 for architectural details.
-1.2 Projector
-The modality-specific representations derived from the dual encoders are first fused using a multi-layer perceptron (MLP).
-This fused representation, which serves as the final 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network.
-This projector, a significantly larger MLP, expands the dimensionality of the fused representation [^1] This is a critical step inspired by Barlow Twins [90] to facilitate effective redundancy reduction during the self-supervised learning phase.
+An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [1].
+See Appendix A.7 and Figure 1 for architectural details.
+Modality-specific embeddings derived from dual encoders are first fused using a multilayer perceptron (MLP) which is trained alongside the encoders.
+This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network.
+This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate effective redundancy reduction during the self-supervised learning phase.
+A detailed breakdown of the model’s architectural parameters is provided in Appendix A.6 (Table 2).
 1.1 Model Training
-The TESSERA model learns its representations through a self-supervised pretraining process on a large-scale, unlabeled global dataset.
-It is trained by optimizing a modified Barlow Twins loss function [90].
-The core principle is to learn embeddings that are invariant to differing or incomplete views of the same pixel’s data (e.g., different sets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions.
-This is achieved by generating two distinct augmented views for each input d-pixel, primarily through independent sparse temporal sampling of annual Sentinel-1 and Sentinel-2 observations, rather than applying artificial distortions or augmentations.
+TESSERA learns the mapping of inputs to embeddings through a self-supervised pretraining process on a large-scale, unlabeled global dataset.
+It is trained by optimizing a modified Barlow Twins loss function [187, 95].
+The core principle is to learn embeddings that are invariant to differing views of the same pixel’s optical-band data (e.g., different subsets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions.
+This is achieved by generating two distinct augmented views for each input d-pixel, through independent sparse temporal samples.
+This avoids creating augmentations by applying artificial (e.g.
+hand-crafted) distortions.
 The model then learns to reconcile these partial views.
-The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, augmented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Appendix A.7.2 for the full loss function, Equation 12).
-The pretraining phase utilizes a vast global dataset comprising approximately 0.8 billion d-pixels, sampled from Sentinel-1 and Sentinel-2 imagery.
-A crucial element for successful generalization was a specific multi-level data sampling, shuffling, and batching strategy.
+The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and intersample interaction during training (see Appendix A.7.2 for the full loss function, Equation 11).
+In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
+The pretraining phase utilizes a vast global dataset comprising approximately 800 million d-pixels, down-sampled from Sentinel-1 and Sentinel-2 imagery (we use a downsampling ratio of 400:1).
+To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed.
 This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features.
-The model was trained for one epoch over this dataset using distributed training techniques.
-1.4 Model Inference and Global Representation Maps
-Following pretraining, the TESSERA dual-encoder (with frozen weights and excluding the projector) is used to generate the final 128-dimensional representation for every 10-meter pixel globally, for each year from 2017 to 2024.
-This involves processing the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders.
-The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA representation maps.
-These maps are designed as readily usable, multi-channel geospatial data layers.
-This “Model-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, pre-computed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself.
+The model was trained for one epoch on this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy and optimization).
+The key hyperparameters used during pretraining are summarized in Appendix A.7.3 (Table 3).
+1.2 Model Inference and Global Embeddings Maps
+Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10-meter pixel globally.
+This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders.
+To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations.
+The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA embedding maps.
+These maps are designed as readily usable multichannel geospatial data layers.
+This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Appendix A.8).
+We are generating global embeddings for 2024 (approx.
+250 TB) and will progressively extend coverage backward year by year until 2017.
 2 Downstream Tasks
-The effectiveness and generalizability of the TESSERA representations are rigorously evaluated across a diverse range of downstream remote sensing tasks.
-These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height estimation), and patch-based dense predictions (e.g., semantic segmentation and land change detection).
-For all such evaluations, the pretrained TESSERA encoder acts as a fixed feature extractor.
-Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed representations.
-This approach typically requires substantially less labeled data and computational resources compared to training deep models from scratch, highlighting the transfer learning capabilities of TESSERA.
-See Appendix A.9 and Supplementary Fig.
+The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks.
+These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height and AGB estimation) and patch-based dense predictions (for example, semantic segmentation).
+In these evaluations, the pre-trained TESSERA encoder acts as a fixed feature extractor; given adequate computing resources, fine-tuning of the model is also possible.
+Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed embeddings.
+This approach typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state of the art.
+See Appendix D and Supplementary Fig.
 10 for the general downstream application workflow.
-2.1 Crop Type Classification Evaluation
-To validate the effectiveness of TESSERA representations for agriculture, we performed a rigorous evaluation on a crop type classification task.
-We used the official Austrian INVEKOS dataset for the 2021-2022 growing season [1], which contains 17 aggregated crop classes.
-Our evaluation methodology centered on comparing the performance of a lightweight classification head trained on TESSERA representations against two key baselines: (1) a traditional Random Forest classifier trained on engineered temporal features from raw satellite data, and (2) the PRESTO foundation model [79], where we applied the same lightweight head to its embeddings for a direct comparison.
-We assessed performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labeled data down to one-shot learning.
+We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [1].
+2.1 Crop Type Classification
+Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [175].
+The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [60].
+However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [60].
+In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [60, 114].
+Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [6, 79].
+Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
+To evaluate performance, we used the INVEKOS Austrian crop dataset [3] from the 2021–2022 growing season.
+This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3].
+We compared the performance of a lightweight classification head trained on TESSERA embeddings against three key baselines: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [149], and (3) Google Satellite Embedding (GSE) [56].
+For the foundation models, we applied the same lightweight head to their respective embeddings for a direct comparison.
+We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning.
 We also evaluated performance on a patch-based semantic segmentation task against several other foundation models.
-Further details on data processing, model architectures, and training protocols are provided in Section B.
-2.2 Crop classification
-Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [86].
-Crop type classification inform vital agricultural decisions by allowing accurate estimates of crop area and yield, and facilitating low-cost monitoring of diseases and pests over vast regions [31].
-However, this task is notably more complex than many other land cover classification challenges, as it requires discerning subtle spectral and temporal variations that differentiate various crop classes [31].
-For years, classifiers like Random Forest (RF) applied to hand-crafted composite time-series data have served as a standard baseline [31, 63].
-Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific training, offer generalizable representations that can be applied universally [4, 41].
-Here, we demonstrate that TESSERA’s pre-computed representations set a new state-of-the-art.
-We compare TESSERA against two critical benchmarks: the traditional RF baseline and the leading pixel-based foundation model, PRESTO (Pretrained Remote Sensing Transformer) [79], a transformer-based model utilizing multi-modal inputs.
-Our results show TESSERA not only substantially outperforms both, but also offers a more streamlined and efficient application workflow.
-To evaluate performance, we used the INVEKOS Austrian crop dataset from the 2021–2022 growing season, an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on phenology and data availability [1].
-TESSERA consistently surpasses both baselines across all training regimes, especially in low-data settings crucial for operational deployment where labeled data is scarce (Fig.
+Details on data processing, model architectures, and training protocols are provided in Appendix E.
+TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labeled data are scarce (Fig.
 2).
-In pixel-wise classification tasks with training data splits from 1% to 30%, a simple multilayer perceptron (MLP) trained on TESSERA representations achieves higher average and balanced F1 scores than both RF and an identical MLP trained on PRESTO embeddings, with performance gains frequently exceeding 10% and 30%, respectively (Fig.
+In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig.
 2a).
@@
 2b).
-Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models across various patch sizes (Fig.
+Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig.
 2c).
-These performance gains are rooted in the superior quality of the learned representations.
-A UMAP analysis of the embedding space reveals that TESSERA’s representations form tighter, more semantically coherent clusters by crop class compared to those from PRESTO (Fig.
+These performance gains are due to the high quality of the learned embeddings.
+A UMAP analysis [103] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from both PRESTO and GSE (Fig.
 2d).
-This improved separability highlights TESSERA’s strength as a foundational tool for agricultural applications.
-Notably, using TESSERA’s compact, pre-computed 128-dimensional representations eliminates the need for the complex data preparation and feature engineering required by traditional methods like RF, democratizing access to high-performance crop classification.
-Figure 2: TESSERA representations achieve state-of-the-art performance in crop type classification.
-The evaluation was conducted on the 2022 Austrian INVEKOS dataset.
-a, Average and Balanced F1 scores for pixel-wise classification as a function of the training data ratio.
-TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model trained on raw time-series data and the PRESTO foundation model.
-b, Overall F1 score in few-shot learning scenarios.
+This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
+Figure 2: TESSERA embeddings achieve state-of-the-art performance in crop type classification.
+The evaluation was conducted on the 2022 Austrian crop dataset.
+a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio.
+TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google Satellite Embedding (GSE).
+b, Patch-based semantic segmentation performance comparison.
+TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios.
+c, Weighted F1 score in few-shot learning scenarios.
 TESSERA maintains a significant performance advantage even with very few training samples per class.
 Error bars represent the standard deviation over multiple runs.
-c, Patch-based semantic segmentation performance comparison.
-TESSERA representations, used with a UPerNet head, achieve higher mean Intersection over Union (mIoU) and Balanced F1 scores than other foundation models across various patch sizes and training data ratios.
-d, UMAP visualization of the embedding spaces for TESSERA and PRESTO for 17 crop classes.
-TESSERA’s embeddings exhibit clearer separation and more coherent clustering, as supported by superior Silhouette and Davies-Bouldin scores, indicating a more semantically meaningful representation space.
+d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes.
+TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
 2.2 Canopy height estimation
-Canopy height is a key structural attribute of forests, closely linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks like REDD+ [57, 26].
-It reflects ecosystem function [56], forest age, and successional stage, while also serving as a proxy for habitat quality and vertical complexity—important determinants of biodiversity patterns and species richness [20, 15, 73].
-Accurate, scalable canopy height models (CHMs) enable detection of forest degradation and disturbance arising from selective logging, disease outbreaks and storm damage.
+Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [106, 49].
+It is correlated with ecosystem function [105], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [37, 25, 136].
+Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
 Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability.
-GEDI, a spaceborne LiDAR system, provides spatially sparse measurements of canopy height that are used to train models for generating maps from optical and radar data; however, GEDI does not sample high latitudes [21], and produces unreliable estimates in mountainous terrain [27].
-Radar-based CHMs (e.g., TanDEM-X) achieve global coverage but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to isolate canopy height accurately.
-Optical data, while widely available, lacks inherent sensitivity to vertical structure and requires indirect proxies or fusion with information from active sensors  [82, 50] - yet none of the three most widely used global CHM models currently implement sensor fusion [46, 68, 78].
-Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., $>500$ features in Potapov et al.
+GEDI, a spaceborne LiDAR system, can provide reasonably accurate canopy height measurements to train models to generate CHMs from optical and radar data; however, it does not provide wall-to-wall maps, and does not sample high latitudes [41], and produces unreliable estimates in mountainous terrain [52].
+Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [131].
+Optical data, while widely available, lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [158, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 121, 148].
+Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al.
 [121]), which may generalize poorly across ecosystems.
-These models tend to under-predict tall canopies, especially in tropical forests where the majority of above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [45], and inconsistencies in input quality.
-As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, with current products underestimating local variability and structural extremes [58].
-Most CHMs are also static, failing to capture seasonal or interannual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
+These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality.
+As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, and current products underestimate local variability and structural extremes [107].
+Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
 These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [108, 116].
-We evaluated TESSERA representations to predict the height of the airborne LiDAR-derived canopy at 10 m resolution within a 5 × 6 km area of tall old-growth tropical forest in the Danum Valley, Borneo[40].
-To ensure spatial independence, we conducted four-fold spatial cross-validation, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation.
-A 30-million-parameter UNet was trained using 64 × 64 pixel patches of the representations as input.
+We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [74] with openly available ground-truth data [33].
+To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation.
+A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input.
 To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set.
-Accuracy metrics for the foundation models are reported as the average across all 12 runs.
-TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of $-0.62$ m.
-We benchmarked TESSERA’s performance against state-of-the-art approaches through three comparisons (Fig.
+The accuracy metrics for the foundation models are reported as the average across the 12 runs.
+TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
+We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig.
 3).
-First, we compared predictions against three global canopy height products: Potapov et al.
-[68] (‘GLAD’ map), Lang et al.
-[46] (‘ETH’ map), and Tolan et al.
-[78] (‘Meta’ map).
-Although local models generally outperform global products in regional evaluations [70], the pronounced saturation of these global models at much lower canopy heights—combined with RMSEs exceeding their reported global averages [68, 46, 78]—underscores the challenge of accurately mapping structurally complex, tall-canopy regions such as Danum Valley.
-Second, we compared TESSERA to the PRESTO [79] foundation model, using the same model architecture and training procedure but replacing the inputs of TESSERA with the representations of PRESTO.
-Third, we compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines generated by Lang et al.
-[47], which used deep convolutional neural networks trained on Sentinel-2 imagery and GEDI LiDAR reference data for regional canopy height and biomass estimation.
+First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.
+[121], the ETH map by Lang et al.
+[84], and the Meta map by Tolan et al.
+[148].
+Although local models generally outperform global products in regional evaluations [127], the pronounced saturation of these global models at much lower canopy heights - combined with RMSEs that exceed their reported global averages [121, 84, 148]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
+We also compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
+[85], which used deep convolutional neural networks trained on Sentinel-2 time series and GEDI LiDAR to estimate regional canopy height and biomass.
+Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [149] and Google Satellite Embedding (GSE) [56], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
+Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
 In all comparisons, TESSERA outperformed competing methods.
-Moreover, its representations do not require preprocessing, greatly enhancing reproducibility.
+Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility.
 Instead, only the region and year of interest are needed to replicate the input data pipeline.
-By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional representations encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
-Figure 3: TESSERA representations outperform global and regional models in canopy height estimation.
-Top row: Density scatter plots compare predicted canopy heights from the TESSERA (left) and PRESTO (middle) models against airborne LiDAR-derived heights.
-TESSERA achieves higher accuracy than PRESTO, demonstrating superior performance in capturing structural forest attributes.
-The rightmost panel shows model bias across 10m canopy height bins, with LiDAR values binned on the x-axis and model predictions on the y-axis.
-TESSERA more closely follows the 1:1 line, particularly in mid- to high-canopy regimes, while PRESTO underestimates heights above  40m.
-Histogram above indicates pixel count distribution per bin.
-Bottom row: Comparison with three global canopy height products— ETH Global CHM, GLAD (Potapov) Global CHM, Meta (Tolan) Global CHM, and ETH Regional CHM—evaluated on the same test region.
-All four products exhibit substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14m$), including negative correlations for three out of four models.
-These results underscore the difficulty of accurately estimating tall tropical canopies using global models trained without local data.
-Rightmost column: Spatial comparison between true canopy height (LiDAR) and TESSERA-predicted canopy height for a representative test patch, illustrating the model’s capacity to capture fine-scale structural variation while still saturating for taller canopy values.
+By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
+Figure 3: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation.
+Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo.
+Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure.
+TESSERA achieves the highest coefficient of determination and the lowest RMSE (8.88±0.988.88$\pm$ 0.988.88 ± 0.98 m), indicating superior performance in structurally complex, tall tropical canopies.
+Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM.
+All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations.
+These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration.
+A 1:1 reference line (blue) is shown in each panel for comparison.
+2.3 Above-ground biomass estimation
+Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [66].
+Accurate estimation of above-ground biomass (AGB) is essential to quantify forest carbon stocks and monitor temporal changes due to deforestation, degradation, and recovery [45].
+We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset.
+Firstly, we considered the geospatial foundation model SpectralGPT [67] combined with a UPerNet regression head – a setup found in the Pangaea benchmark [101] to outperform other foundation models tested on the BioMassters dataset [111].
+As the preprocessing pipeline is not yet available in the codebase of Ref.
+[67] at the time of our analysis, we adopted the implementation from Ref.
+[101], which uses the preprocessed satellite imagery from the BioMassters dataset as input.
+Secondly, we used the winning model from the BioMassters competition as our state-of-the-art supervised baseline – a UNet with a temporal attention encoder [111], which achieved the lowest RMSE among over 1,000 submissions.
+The BioMassters dataset comprises 11,462 image patches that represent above-ground biomass in Finnish forests from 2017 to 2021.
+Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey.
+Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel.
+The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
+A comparison of RMSE across varying training label fractions (Fig.
+4a) shows that a UNet trained on the TESSERA embeddings consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set.
+These results underscore the robustness of TESSERA in sparse-label regime.
+Moreover, TESSERA achieves results comparable to those of the supervised baseline trained on the full dataset, even with limited label availability.
+In general, bespoke models are inherently difficult to surpass in fully supervised settings.
+It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input.
+However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
+Figure 4: TESSERA embeddings demonstrate robustness for AGB mapping under limited label availability.
+a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGPT.
+Following the evaluation protocol in Ref.
+[111], RMSE is computed on a fixed test set of 2773 patches and reported as the average RMSE per patch.
+For limited labels testing, a subset of labels is randomly chosen from the remaining data, and randomly split into training and validation sets (80:20).
+The figure shows the average RMSE computed over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
+The UNet model is trained on the TESSERA embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
+The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability.
+b, Predicted AGB map for an example patch using 10% of the training labels.
+c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset.
+The top panel depicts the histogram of the ground-truth AGB values.
+d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top), the supervised baseline (middle) and SpectralGPT+UPerNet (bottom).
 2.4 Burned Area Detection
-Wildfire frequency and severity are projected to increase in many parts of the world due to climate change [3, 39, 66].
-Accurate maps of burned areas (BAs) and burn severity are critical for monitoring wildfire trends [6], predicting wildfire occurrence [17], assessing ecosystem recovery [89], and establishing effective fire management strategies [33].
+The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [5, 71, 117].
+Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [10], predicting wildfire occurrence [29], assessing ecosystem recovery [186], and establishing effective fire management strategies [63].
 Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82].
-Most existing BA mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [44, 67].
+Most existing burn area mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 119].
 These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), though many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82].
-The majority of existing methods use optical sensors [44], which can detect fire-specific spectral signatures but are susceptible to cloud cover [34], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [35].
+Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [64], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [65].
 SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 143, 154].
 Recent reviews emphasize the need for sensor fusion, combining passive and active sensors to improve accuracy and robustness [82, 30].
-Sentinel sensors in particular, when combined together at 10 m resolution, have shown promise in detecting BA from small fires that are often left out of coarse, global BA products [76, 81, 71].
-TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of abstract, data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig.
-4).
-Figure 4: TESSERA representations capture wildfire extent, timing, and severity across two burned regions.
+Sentinel-1 and 2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [143, 154, 129].
+TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig.
+5).
+Figure 5: TESSERA embeddings capture wildfire extent, timing, and severity across two burned regions.
 (a–c) Burn Area A, which experienced two distinct fires in June and August 2021.
@@
 (f) UMAP projection for Burn Area B, revealing fuzzier boundaries between burned and unburned pixels, though the severity gradient remains evident.
-We tested whether TESSERA’s temporal representations capture fire-induced changes in spectral and radar signals.
-We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 × 15 km regions from the Burn Scar Benchmarking dataset, each affected by wildfires in California during 2021.
-TESSERA representations for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels.
-Fire boundaries and burn severity data were sourced from the Monitoring Trends in Burn Severity (MTBS) dataset[2].
+We tested whether TESSERA’s temporal embeddings capture fire-induced changes in spectral and radar signals.
+We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 $\times$ 15 km regions from the Monitoring Trends in Burn Severity (MTBS) dataset [4], each affected by wildfires in California during 2021.
+TESSERA embeddings for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels.
+Fire boundaries and burn severity data were sourced from the dataset.
 The resulting UMAP projections demonstrated strong separability in two-dimensional space, visually highlighting TESSERA’s capacity to detect disturbances at the pixel level.
-Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single, large fire affecting mostly forested areas, resulting in heterogeneous burn severity.
+Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single large fire that affected mostly forested areas, resulting in a heterogeneous severity of the burn.
 The UMAP revealed separability across three domains (Fig.
-4):
-  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned ones, with clearer separation in Burn Area A.
+5):
+  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned, with a clearer separation in Burn Area A.
   • Temporal fire differentiation: In Burn Area A, embeddings separated pixels burned in June vs. August, suggesting temporal sensitivity to disturbance events.
-  • Burn severity gradient: In both regions, a gradient structure in UMAP space corresponded to burn severity, indicating that TESSERA encodes disturbance magnitude.
-Because TESSERA encodes annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images.
-This capacity would support a simplified, scalable pipeline for large-area disturbance monitoring.
-2.5 Above-ground biomass estimation
-Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [36].
-Accurate estimation of above-ground biomass (AGB) is essential for quantifying forest carbon stocks and monitoring temporal changes due to deforestation, degradation, and recovery [22].
-We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset.
-Firstly, we considered the geospatial foundation model SpectralGPT [37], which was found in the Pangaea benchmark [53] to outperform other tested foundation models on the BioMassters dataset [60].
-Secondly, as the state-of-the-art bespoke model, we considered the winner of the Biomassters competition, a UNet with a temporal attention encoder [60].
-The BioMassters dataset comprises 11,462 image patches representing above-ground biomass across Finnish forests from 2017 to 2021.
-Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey.
-Each patch covers a 2560×2560 m area, with 10×10 m spatial resolution per pixel.
-The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
-A comparison of RMSE across varying training label fractions (Fig.
-5a) shows that a UNet trained on the TESSERA representations consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set.
-These results underscore the robustness of TESSERA in low-label regimes.
-Moreover, while the best task-specific model trained on the full dataset outperforms TESSERA, the gap is moderate, even with low label availability.
-In general, bespoke models are inherently difficult to surpass in fully supervised settings.
-It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input.
-However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
-Figure 5: TESSERA representations demonstrate robustness under limited label availability.
-a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGBT (RMSE values from Ref.
-[53]).
-Following the evaluation protocol of Ref.
-[53], RMSE is computed on a fixed test set of 2773 patches, with the remaining data randomly split into training and validation sets (80:20).
-The UNet model is trained on the TESSERA representations for 80 epochs, with the checkpoint achieving the lowest validation loss selected as the final model.
-For limited labels testing, a subset of labels is randomly chosen from the training set.
-The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability.
-b, Predicted AGB map for an example patch using 10% of the training labels.
-c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset.
-d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top) and Biomassters winner (bottom).
+  • Burn severity gradient: In both regions, a gradient structure in the UMAP space corresponded to burn severity, indicating that TESSERA encodes the magnitude of the disturbance.
+Because TESSERA encodes an annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images.
+This capacity supports a simplified, scalable pipeline for monitoring large-area disturbances.
 2.5 Stocking indices in voluntary carbon markets
-The most widely applied standard in the voluntary carbon market (VCM) for reforestation and afforestation projects of degraded pastures requires project proponents to demonstrate that their chosen remotely-sensed “stocking index” is correlated with the AGB (REF).
-Upon acceptance, project developers can use the index to quantify project additionality and to assess project performance in-between the 5-yearly ground-truth assessments.
-There is considerable freedom in choosing a stocking index, which can be either a pre-existing or proprietary canopy height or AGB product.
-Here, we evaluate TESSERA’s performance against five widely used, globally available canopy height and AGB products often used in the VCM.
-The in situ data used to compare the different stocking indices was collected by ICRAF in 38 sites in Para State, Brazil, that were degraded pastureland until being converted into agroforestry systems during 1980-2010 (Atzberger et al., 2025).
-To provide a fair comparison, stocking indices were rescaled using a linear transformation to minimize the RMSE with respect to the in situ data (details in the SI).
-Figure 6: TESSERA outperforms commonly used stocking indices in the Voluntary Carbon Market.
-Comparison of in-situ aboveground biomass measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products.
-Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using X * 15.502 + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA: AGB converted using 0.407 X + 34; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X above ground biomass prediction converted using 0.3X + 54.3.
-Details of conversation functions provided in SI
+The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration while minimizing leakage liabilities [110].
+To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
+This is a costly and inefficient undertaking if it is done only by field sampling.
+To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [155]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers merely need to demonstrate that the chosen stocking index is correlated with the AGB.
+Upon acceptance, project developers can use the index to quantify the additionality of the project and to assess the performance of the project.
+Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
+To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise RF regression following the methodology outlined in Atzberger et al.
+[13].
+Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [1] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil.
+This rescaling ensures a fair comparison between the various products.
+A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Figure 6).
+As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
+Figure 6: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market.
+Comparison of in-situ AGB measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details).
+Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 X + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA AGB converted using 0.407 X + 34.0; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X AGB prediction converted using 0.3 X + 54.3.
+Details on the conversion functions are provided in Supplemental section I.
 3 Conclusion
-In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that achieves worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level satellite time series data.
-The representations generated by TESSERA establish new performance standards, while our open source methodology ensures widespread access to superior, high-resolution representations.
+In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
+TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing.
 We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models.
-The findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
-We next provide supplementary materials and methods about the details of how TESSERA is implemented.
-Appendix A Earth Observation data
+Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
+Acknowledgments
+We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
+Funding:
+ZF and SK were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge.
+SJ was funded by a charitable donation from John Bernstein.
+SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
+Author contributions:
+TESSERA conceptualization: CA
+System design: CA, ZF, SJ, SK
+System implementation: ZF, RY, MCL, AM
+Data curation: ZF, RY, JK, SJ, AM, DAC, SK
+Downstream tasks: MCL, CA, SS, DAC, CA, MI, ZF, JK
+Paper conceptualization: SK, AM, DAC, AB, SJ
+Paper draft: ZF, MCL, RY, SS, SK, JK
+Paper review and refinement: AB, CA, DAC, SJ, SK
+Competing interests:
+There are no competing interests to declare.
+Data and materials availability:
+Our framework integrates openly available data from the Copernicus Sentinel-2 (S2) [47] optical and Sentinel-1 (S1) [46] SAR missions, leveraging the strengths of each sensor.
+We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22).
+To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations.
+For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [48].
+The Google Satellite Embedding V1 dataset [56] is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
+PRESTO embeddings can be created on demand in Google Earth Engine [57] using VertexAI [58].
+Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code [152].
+To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository [124], which is the official third-party software repository for Python.
+This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer [62] that can be passed directly into downstream tasks.
+The embeddings themselves are checksummed and indexed using Pooch [151] to check data integrity and handle retrieval.
+The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [3].
+The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) datasest [4].
+The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [33].
+The AGB downstream task uses the open BioMassters dataset [111].
+The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners, hence data access will require a Data Transfer Agreement.
+Appendix A Model details
 Our input data consists of Sentinel-2 optical satellite imagery (Level-2A bottom-of-atmosphere) and Sentinel-1 Synthetic Aperture Radar (SAR, Radiometrically Terrain Corrected).
 A.1 Data Representation
-We consider remote sensing data with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
+We consider remote sensing data [46, 47] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
 $R_{t}\in\mathbb{R}^{W\times H\times C}$ (1)
@@
 $V_{t}\in\{0,1\}^{W\times H}$ (2)
-where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
+[…]
 A.2 Temporal Data Stacking
-We stack spatially aligned tiles over a time period spanning $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
+We stack spatially aligned tiles over a time period that spans $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
 $\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (3)
@@
 A.3 Time Series Extraction and d-pixel Definition
-For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
+[…]
 $S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (5)
-We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
-$P_{i,j}=S(i,j)$ (6)
+[…]
+$P_{i,j}(c)=S(i,j,c)$ (6)
 In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time.
-Note that d-pixels are potentially sparse and have an accompanying mask $m_{i,j}$ that indicates the timesteps in which there is valid data:
-$m_{i,j}=[v_{i,j,0},v_{i,j,1},\ldots,v_{i,j,T-1}]$ (7)
+Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there is valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
+Table 1: Variable definitions for remote sensing data structure
 Variable Dimensions Description
@@
 $\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
-$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
-$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
-$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
+[…]
+[…]
+[…]
 $W,H$ Scalar Spatial dimensions (width, height)
-$C$ scalar Number of spectral channels
+$C$ Scalar Number of spectral channels + polarizations
 $T$ Scalar Number of time steps
-Table 1: Variable definitions for remote sensing data structure
 A.4 The d-pixel Representation
-We represent the rich information contained within satellite image time series at the level of individual pixels using the aforementioned “d-pixel”.
-For a single 10-meter geographic pixel location, it consolidates its complete annual multi-spectral (from Sentinel-2) and SAR backscatter (from Sentinel-1) time series.
-This structure explicitly preserves the inherent spectral and temporal dimensions of the data (Supplementary Fig.
-1 a illustrates the d-pixel concept).
-In this array, each row corresponds to a distinct observation date throughout the year, ordered chronologically.
-Each column corresponds to a specific spectral band or polarization.
-The value at the intersection of a given row and column thus represents the measured reflectance or backscatter for that specific band/polarization on that particular date for that single pixel.
-A critical feature of the d-pixel is its capacity to accommodate the frequent data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities.
-Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked within the d-pixel array.
-These missing entries are explicitly handled by downstream model components, for instance, by being ignored during attention computation or managed during the temporal sampling process described in Appendix A.7.3.
+We represent each 10-m pixel in the time series of images from annual multi-spectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
+1 a).
+In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization.
+This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities.
+Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored during attention computation or ignored in the temporal sampling process.
 A.5 Dual-Encoder Architecture
 Given the distinct nature of Sentinel-1 SAR and Sentinel-2 MSI data, TESSERA employs two separate, parallel Transformer-based encoder branches.
-  • Sentinel-2 MSI Encoder: This branch processes time series of 10 spectral bands from Sentinel-2.
+  • Sentinel-2 MSI Encoder: This branch processes a time series of 10 spectral bands (excluding the 60 m bands used for the detection of water vapor and clouds) from Sentinel-2.
 We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
-  • Sentinel-1 SAR Encoder: This branch processes time series of 2 polarizations from Sentinel-1 (VV and VH).
+  • Sentinel-1 SAR Encoder: This branch processes a time series of 2 polarizations from Sentinel-1 (VV and VH).
 Each encoder begins by linearly embedding the input features (spectral bands or polarizations) for each time step.
 To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings.
-The core of each encoder consists of a stack of 8 standard Transformer blocks [80], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
+The core of each encoder consists of a stack of 4 standard Transformer blocks [153], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
 To derive a single vector summarizing the entire time series for each modality, an attention-pooling layer weighs the importance of different time steps before aggregation.
-The resulting modality-specific representations (one from the S1 encoder, one from the S2 encoder) are then fused using a multi-layer perceptron.
+The resulting modality-specific embeddings (one from the S1 encoder, one from the S2 encoder) are then fused using a multi-layer perceptron.
 A.6 Projector Network
-The fused representation from the dual-encoder stage is subsequently fed into a large projector network.
+The fused embedding from the dual-encoder stage is subsequently fed into a large projector network.
 This projector is a five-layer MLP, with each hidden layer having 16,384 dimensions.
@@
 The final output of the projector for each input d-pixel is an embedding, which is then used in the loss calculation.
-For downstream tasks, we typically use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel representation.
-The TESSERA encoder (up to the fusion MLP) has approximately 40 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
+For downstream tasks, we use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding.
+The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
+Table 2: Architectural parameters of the TESSERA v1.0 model.
+Component Parameters Percentage of Total
+Encoder (Feature Extractor) 45,697,026 3.29%
+   Sentinel-1 Encoder 22,846,465 1.64%
+   Sentinel-2 Encoder 22,850,561 1.64%
+Fusion MLP 131,200 $<$0.01%
+Projector (for Training only) 1,344,700,421 96.71%
+Total Parameters 1,390,528,647 100.00%
 A.7 Self-Supervised Training
@@
 The indices $i$ and $j$ iterate over the dimensions of the embedding vectors.
-The first term (invariance term) encourages similar representations for different views of the same input ($C_{ii}\to 1$).
-The second term (redundancy reduction term) promotes informational efficiency by minimizing correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
-To further enhance model robustness and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term, $\mathcal{L}_{MIX}$, inspired by Bandara et al.
+The first term (invariance term) encourages similar embeddings for different views of the same input ($C_{ii}\to 1$).
+The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
+We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [187].
+To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al.
 [16].
@@
 The embeddings $Z_{M}$ and $Z_{S}$ are obtained.
-The mix-up loss penalizes deviations from the assumption that a linear interpolation in input space corresponds to a linear interpolation in embedding space:
+The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
 $C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (8)
@@
 $\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
-where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{BT}=5\times 10^{-3}$ and $\lambda_{mix}=1.0$ to be effective.
-A.7.3 Pretraining Details
-The TESSERA model (Alpha version) was pretrained using approximately 0.8 billion d-pixel samples derived from 3,012 globally distributed Military Grid Reference System (MGRS) tiles.
-For pretraining, these d-pixels were generated from Sentinel-1 and Sentinel-2 data that were spatially downsampled by a factor of 400 from their native 10-meter resolution.
-Each d-pixel for Sentinel-2 contained time series for 10 spectral bands, and for Sentinel-1, 2 polarizations (VV and VH).
+where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [16]) to be effective based on our experiments.
+A.7.3 Pre-training Details
+The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles.
+d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400, i.e, selecting one pixel from each 20$\times$20 patch.
+Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
+For each pixel location, after removing invalid observations (e.g., due to cloud cover for Sentinel-2), we performed sparse temporal sampling.
+This involves randomly selecting a fixed number of 40 valid observation dates from the year’s data.
 The sequence length for the Transformer encoders was fixed at 40 timesteps for both modalities.
-For each pixel location, after filtering invalid observations (e.g., due to cloud cover for Sentinel-2), we performed sparse temporal sampling.
-This involves randomly selecting a fixed number of 40 valid observation dates from the year’s data.
-If fewer than 40 valid dates exist, sampling is performed with replacement.
+If there are fewer than 40 cloud-free dates, replacement sampling is performed.
 This strategy standardizes the input sequence length and serves as a key data augmentation mechanism, building invariance to data gaps and teaching the model that the underlying signal persists regardless of the specific dates observed.
+Finally, these values were standardized using global statistics to stabilize training.
 The temporal context of each sampled observation was encoded by transforming its normalized Day-of-Year (DOY) into sine and cosine features, which were then concatenated with the corresponding spectral or backscatter measurements.
-Finally, these values were standardized using global statistics to stabilize training.
-Figure 7: Geographical distribution and data density of the training dataset.
-TESSERA was trained on over 3,000 MGRS tiles distributed globally from 2017 to 2024.
+Table 3: Pre-training hyperparameters for the TESSERA v1.0 model.
+Description Symbol Value
+Optimizer & Schedule
+Optimizer - AdamW
+Batch Size (Per-GPU) $B_{local}$ 2,048
+Batch Size (Global) $B_{global}$ 32,768 (16 GPUs $\times$ 2,048)
+Epochs - 1
+Base Learning Rate $\eta$ 0.002
+Weight Decay $\lambda_{wd}$ $1\times 10^{-6}$
+Learning Rate Schedule - Linear warmup (10%), plateau (20%), cosine decay
+Gradient Clipping Norm - 2.0
+Loss Function
+Barlow Twins Lambda $\lambda_{BT}$ $5\times 10^{-3}$
+Mixup Lambda $\lambda_{mix}$ 1.0
+Mixup Beta Parameter $\beta_{p}$ 1.0
+Architecture & Data
+Sequence Length $L$ 40
+Representation Dimension $D_{repr}$ 128
+Projector Output Dimension $D_{proj}$ 16,384
+Quantization Bits (QAT) - 8
+Figure 7: TESSERA is trained on globally-distributed training data with greater data density in the northern hemisphere.
+TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024.
 The color of each hexagon in the map corresponds to the number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate color legend.
 This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
-The model was trained for 1 epoch over the entire dataset, which corresponded to approximately 3,000 GPU hours on 8 AMD MI300X GPUs (192GB memory each).
+As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each).
 We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled.
-The AdamW optimizer was employed with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$.
-The learning rate schedule included a linear warmup over the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay.
-The global batch size was 8192.
-A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a vast and geographically diverse dataset.
+AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$.
+The learning rate schedule included a linear warming during the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay.
+The global batch size was 32768.
+A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
 Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch.
-To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$0.8 billion training samples, which constituted over 2TB of initial d-pixel data.
-This pipeline is illustrated conceptually in Fig.
+To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2TB of initial d-pixel data.
+This pipeline is conceptually illustrated in Fig.
 8a.
@@
 8b.
-Compared to a conventional, localized shuffling strategy which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
+Compared to a conventional, localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
 This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
-Figure 8: Overview of the data shuffling pipeline and its impact on training stability.
+Figure 8: Data shuffling greatly improves training stability.
 a, Schematic of the data shuffling and loading process.
 D-pixels (colored squares) from thousands of MGRS tiles are first aggregated into a global pool.
-A custom Rust binary performs a global shuffle on this multi-terabyte dataset before applying augmentations.
+A custom Rust binary [152] performs a global shuffle on this multi-terabyte dataset before applying augmentations.
 The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches.
@@
 As detailed in Appendix 5.1, this involves generating two distinct augmented views (e.g., $Y^{A}$ and $Y^{B}$) for each d-pixel.
-To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed the pipeline as a custom Rust binary.
-This high-performance binary handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation.
+To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed a custom high-performance pipeline that handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation.
 The resulting pairs of augmented d-pixels are then serialized into a compact, pickle-like file format.
@@
 This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
-A.8 Global Representation Map Generation
-A primary output of the TESSERA project is the generation of annual global representation maps at 10-meter resolution for the years 2017-2024.
-A.8.1 Model Inference
-To generate these maps, the pretrained and frozen TESSERA dual-encoder (excluding the projector) is used. For each 10-meter pixel on the globe and for each year:
+A.8 Global Embedding Map Generation
+A primary output of the TESSERA project is the generation of annual global embedding maps with 10-meter resolution for the years 2017-2024.
+To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used.
+For each 10-meter pixel on the globe and for each year:
   1. 1.
-    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pretraining, no spatial downsampling is performed at this stage.
+    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
   2. 2.
-    A fixed number of 40 timesteps are sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
+    A fixed number of 40 timesteps is sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
   3. 3.
@@
   4. 4.
-    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional representation vector for that pixel for that year.
-This process is repeated for all land pixels globally to create an annual representation map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
-A.8.2 Data Product and Accessibility
-The resulting product consists of eight such global representation maps, one for each year from 2017 to 2024.
-These maps are intended to be released publicly, alongside the open-sourced TESSERA model parameters and generation code.
-This “Model-as-Data” approach significantly lowers the entry barrier for users, who can treat these representation maps as conventional multi-channel images.
-They can be directly ingested by downstream models without requiring users to process raw satellite data or run the TESSERA model themselves, thus democratizing access to advanced RSFM capabilities.
-A.9 Downstream Task Application Methodology
-A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature representations that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data.
-Having pretrained TESSERA on large unlabeled datasets, we evaluated the utility of its learned representations across different remote sensing applications.
-The evaluation methodology involves using the pretrained TESSERA encoders as fixed feature extractors.
-The pipeline for applying TESSERA representations to downstream tasks is as follows (illustrated in Supplementary Fig. 9):
-Figure 9: Illustration of RSFM training and downstream task applications.
-(b) shows the training process of a typical RSFM, where both the encoder and the projector (or decoder) are updated.
-(a) depicts a common application method: the RSFM encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
-Users can then fine-tune the entire model or just the head.
-(c) presents the emerging paradigm championed by TESSERA, where pre-generated representations from the RSFM encoder are used directly as input.
-This results in extremely lightweight downstream models that consist only of a task-specific head.
+    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional embedding vector for that pixel for that year.
+This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
+Appendix B Scaling with data and network size
+To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task.
+Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
+9).
+Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
+We found that using 4 transformer layers, each with 4 attention heads, provided a strong baseline without incurring excessive computational overhead during inference.
+This makes the feature-extracting part of the model computationally efficient.
+In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning.
+We performed a parameter sweep over the projector’s depth (number of layers) and width (number of neurons per layer), with results shown in Fig.
+9a.
+As discussed by Zbontar et al.
+[187], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task.
+The highest validation score was achieved with the largest configuration tested (8 layers deep, 32,768 neurons wide).
+However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure.
+For the final TESSERA v1.0 model, we selected the configuration with 4 layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
+Furthermore, we analyzed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
+9b.
+The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
+This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks.
+The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
+Figure 9: Model performance improves with increasing projector size and additional computation.
+a, Hyperparameter sweep for the projector network’s architecture.
+Each point represents a configuration, with its position determined by depth (x-axis) and width (y-axis).
+The color and size of the point indicate the validation F1 score and the number of parameters, respectively.
+Performance generally increases with both width and depth.
+The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints.
+The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency.
+b, Scaling laws for TESSERA.
+The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters).
+A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used.
+The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
+Appendix C Ablation study
+To validate the contribution of each key component within the TESSERA framework, we conducted an ablation study.
+We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
+Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score [54].
+The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions.
+Given a batch of embeddings represented by the matrix $Z\in\mathbb{R}^{N\times D}$, where $N$ is the batch size and $D$ is the embedding dimension, we first compute its singular values $s_{1},s_{2},\ldots,s_{D}$.
+These are normalized to form a probability distribution $p_{i}=s_{i}/\sum_{j}s_{j}$.
+The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
+$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (12)
+A score closer to 1 indicates that the embeddings are of higher effective rank, utilizing the full dimensionality of the feature space more effectively to create richer, less redundant representations.
+The results of the ablation study are summarized in Table 4.
+Our analysis reveals several key insights into the model’s architecture and training strategy.
+The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches.
+Removing mixup regularization also severely degrades results.
+When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable, robust representations.
+With removal of Sentinel-1 data, task performance decreases moderately, demonstrating some gain from the fusion of optical and SAR data yields.
+However, performance is still relatively good even with only Sentinel-2 data.
+This highlights both TESSERA’s overall robustness and its continued efficacy in regions or historical periods where Sentinel-1 data are sparse or unavailable, which is not uncommon in practice.
+Finally, the effect of quantization is particularly noteworthy.
+We employ Quantization-Aware Training (QAT) to convert the embeddings from 32-bit floating-point (float32) to 8-bit integer (int8) format.
+Removing this quantization step (i.e.
+using full float32 precision) results in a marginal performance improvement.
+This demonstrates the success of our QAT strategy: we achieve a nearly four-fold reduction in the storage and bandwidth requirements for our final data products, with a negligible impact on embedding quality.
+This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
+Table 4: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance.
+We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset.
+The baseline represents the full model.
+Each subsequent row shows the result of removing the specified component.
+Performance is measured by validation F1 score and the RankMe score.
+Arrows ($\uparrow$) indicate that higher values are better.
+The results demonstrate that all components contribute positively to the model’s overall performance.
+Configuration Validation F1 Score ($\uparrow$) RankMe Score ($\uparrow$)
+Baseline (Full Model) 0.773 0.963
+w/o Global Data Shuffling 0.681 0.847
+w/o Mixup Regularization 0.662 0.857
+w/o Sentinel-1 Data 0.742 0.931
+w/o Shuffling & Mixup 0.626 0.867
+w/o Embedding Quantization 0.779 0.972
+Appendix D Downstream Task Application Methodology
+A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data.
+Having pre-trained TESSERA on large unlabeled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications.
+The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
+The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary Fig. 10):
+Figure 10: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications.
+Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
+Users can then fine-tune the entire model or only the head.
+Another approach (b) is where both the encoder and the projector (or decoder) are typically updated.
+In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input.
+This results in lightweight downstream models that consist only of a task-specific head.
   1. 1.
-    Load Pretrained Encoder: The weights from the saved pretraining checkpoint are loaded into the TESSERA dual-encoder architecture.
-The parameters of these encoders are frozen and are not updated during downstream model training.
-This ensures that downstream performance directly reflects the quality of the fixed representations.
+    Download Embeddings for the Region of Interest: The Geotessera Python library [152] allows users to download embeddings for a desired region and year in the form of a numpy array.
   2. 2.
-    Prepare Labeled Downstream Data: The specific labeled dataset for the target task (e.g., pixel-level crop type labels, canopy height measurements, or land change polygons) is prepared.
+    Prepare Labeled Downstream Data: The labeled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
   3. 3.
-    Extract Representations: For each input sample (e.g., pixel, object, or patch) in the labeled dataset, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year undergo the same preprocessing and d-pixel creation pipeline used during TESSERA inference (i.e., using 10m resolution data and temporal sampling to 40 timesteps).
-The preprocessed sequences are passed through the respective frozen encoders, and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA representation for that sample.
-To enhance stability, this extraction can be repeated multiple times with different random temporal samplings, and the resulting representations averaged, although for many tasks, a single extraction is sufficient.
+    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA embeddings as input.
+    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (1-3 layers) ending in a softmax output layer.
+    • For pixel-wise regression (e.g., canopy height regression), the head is usually an MLP ending in a single linear output neuron.
+    • For tasks requiring spatial context from the embeddings (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA embeddings (e.g., 64×64×12864$\times$ 64$\times$ 12864 × 64 × 128).
+The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions.
   4. 4.
-    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA representations as input.
-    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (e.g., 1-3 layers) ending in a softmax output layer.
-    • For pixel-wise regression (e.g., canopy height regression), the head is usually an MLP ending in a single linear output neuron.
-    • For tasks requiring spatial context from the representations (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA representations (e.g., $64\times 64\times 128$).
-The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions.
-For land change detection, a simple approach involves computing the dot product between representations from two different years.
-  5. 5.
-    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA representations as input features and the corresponding labels.
+    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA embeddings as input features and the corresponding labels.
 Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used.
 This training typically requires significantly less labeled data and computational power compared to training a deep model from scratch.
-  6. 6.
-    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA representations for the test samples and passing them through the trained head.
+  5. 5.
+    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA embeddings for the test samples and passing them through the trained head.
 Performance is evaluated using standard metrics relevant to the task.
-This standardized workflow allows for robust assessment of TESSERA representations across diverse applications, demonstrating their value as foundational features for geospatial analysis.
-Appendix B Downstream Task: Austrian Crop Classification
-The following section details the experimental setup for the crop classification task presented in the main text.
-B.1 Dataset and Preprocessing
+This workflow allows the use of TESSERA embeddings in a range of diverse applications, demonstrating its role as a foundational model for geospatial analysis.
+TESSERA users who would like to generate their own embeddings can load the frozen weights from the saved pre-training checkpoint into the TESSERA dual encoder architecture.
+For each input sample (e.g., pixel, object, or patch) in the region of interest, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year must be downloaded and processed using the d-pixel creation pipeline.
+These d-pixels are passed through the frozen encoders and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA embedding for that sample.
+To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
+Appendix E Downstream Task: Austrian Crop Classification
+E.1 Dataset and Preprocessing
 We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [3].
-The dataset originally contained 154 crop types, which we grouped into 17 broader classes (e.g., merging different varieties of wheat) based on phenological similarity and sample availability to ensure robust training and evaluation.
-For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA representation from our generated 2022 global representation map.
-B.2 Pixel-wise Classification Baselines
+The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation.
+For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
+E.2 Pixel-wise Classification Baselines
 To provide a comprehensive performance comparison, we implemented three distinct models for pixel-wise classification:
-  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA representations were used as input to a simple MLP.
+  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA embeddings were used as input to a simple MLP.
 The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
-  • PRESTO + MLP: For a direct and fair comparison with the closest foundation model, we used the official pre-trained PRESTO model [79] to generate its pixel embeddings.
-These embeddings were then fed into an identical MLP head as the one used for TESSERA.
-  • Random Forest: As a traditional baseline, we trained a Random Forest classifier directly on raw time-series data.
+  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model [149] to generate its pixel embeddings.
+These embeddings were then fed into an MLP head identical to the one used for TESSERA.
+  • GSE + MLP: To include another key foundation model, we used the publicly available Google Satellite Embedding (GSE) V1 dataset [56].
+These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
+  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data.
 For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year.
@@
 For the experiments shown in the main text’s figure, we trained these models on randomly selected subsets of the data (from 1% to 30% for panel a; specified samples-per-class for panel b), using a fixed validation set for hyperparameter tuning and a held-out test set for final evaluation.
-B.3 Patch-wise Semantic Segmentation
+E.3 Patch-wise Semantic Segmentation
 To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
-  • For pixel-based foundation models like TESSERA and PRESTO, which do not explicitly model spatial context in their representations, we first constructed representation patches (e.g., of size $64\times 64\times 128$) from the pixel-wise embeddings.
-These patches were then fed into a standard UNet architecture, which learns to model the spatial relationships between the representations to produce a segmentation map.
+  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [149], and GSE [56], a distinct workflow was used.
+We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$).
+These patches were then fed into a standard UNet architecture [128], which learns to model the spatial relationships between the embeddings to produce a segmentation map.
+It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do.
+However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
   • For other foundation models that are inherently patch-based (e.g., Prithvi [141], Satlas [18]), their encoders already process image patches.
-Therefore, for these models, we only needed to attach and train a UPerNet decoder head to their frozen encoders to generate the final segmentation outputs.
-Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores.
-B.4 Embedding Space Analysis
-The 2D visualizations shown in the main text were generated by applying the Uniform Manifold Approximation and Projection (UMAP) algorithm to the representations of every pixel within the Austrian study area.
-Specifically, the entire representation map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as direct input to UMAP.
-This was performed for both TESSERA and PRESTO representations.
+Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output.
+We did not freeze the encoders during the finetuning.
+Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 11.
+E.4 Embedding Space Analysis
+The 2D visualizations shown in the main text (Figure 2 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [103] algorithm to the embeddings of every pixel within the Austrian study area.
+UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology.
+It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure.
+UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
+In this work, the entire embeddings map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for TESSERA, PRESTO, and GSE embeddings.
 To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
@@
     where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$.
-Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where clusters are compact and well-separated.
-Acknowledgments
-We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
-Funding:
-Z.
-F.
-and S.
-K were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge.
-S.
-J was funded by a charitable donation from John Bernstein.
-J.
-B received support from the UKRI CRCRM scheme (grant MR/Y024354).
-S.
-S.
-received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
-Author contributions:
-System design: CA, ZF, SJ, SK
-System implementation: ZF, RY,
-Data curation: ZF, RY, JK, SJ, AM, DAC, SK
-Downstream tasks: ML, CA, SS, DAC, CA, MI, ZF, JK
-Paper conceptualization: SK, AM, DAC, AB, SJ
-Paper draft: ZF, ML, RY, SS, SK, JB, JK
-Paper review and refinement: AB, AM, DAC, SJ, SK
-Competing interests:
-There are no competing interests to declare.
-Data and materials availability:
-Our framework integrates data from the Copernicus Sentinel-2 (S2) [24] optical and Sentinel-1 (S1) [23] SAR missions, leveraging the strengths of each sensor.
-We utilized S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands  [^2] We used red, blue, green, NIR, NIR8, red-edge1, red-edge2, red-edge3, SWIR16, and SWIR22..
-A challenge with optical data is frequent cloud contamination; we filtered the time series using the associated Sen2Cor scene classification map (SCL) to mask cloudy observations.
-For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [25].
+Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where clusters are compact and well separated.
+Figure 11: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
+Each row corresponds to a different patch size (32x32, 64x64, 128x128).
+The first column shows the ground truth segmentation.
+For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labeled ’Embedding’) and the final predicted segmentation map (labeled ’Predicted’).
+Appendix F Downstream Task: Canopy height estimation
+F.1 Study Areas and Ground Truth Datasets
+Canopy height prediction was evaluated in a 5$\times$6 km area in the Danum Valley, Borneo.
+This region is covered in tropical rainforest dominated by the fast-growing dipterocarp family, which hosts some of the tallest trees in the tropics [135, 120].
+Ground truth data was derived from airborne laser scanning (ALS) performed in February 2020 [33].
+ALS data provides detailed 3D point clouds of vegetation and terrain structure, which were processed using the  lspikefree algorithm [51] to estimate top-of-canopy height at 1 m resolution.
+These canopy height rasters were then downsampled using mean pooling to 10 m to match the resolution of the TESSERA representations.
+F.2 Modeling Methods
+To predict the canopy height from the different embeddings, we used a standard U-Net architecture [128].
+It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations.
+For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube.
+The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values.
+The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches.
+At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation.
+In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training.
+Model training was done in Python using PyTorch.
+To ensure spatial generalization and avoid overfitting local patterns, the reference data were divided into training, validation, and test sets that are spatially disjoint.
+Each study region was divided into four spatial folds based on cardinal directions (bottom, top, left, right), each fold covering 50% of the area.
+For each fold, the remaining half of the region was used for training and validation.
+To account for inherent model randomness, we trained three independent models per fold, resulting in 12 trained models for each foundation model.
+This strategy captures variability due to both spatial heterogeneity and training randomness.
+The final performance of the model was evaluated on the test data held using the following metrics: coefficient of determination (R²), root mean square error (RMSE), and mean bias.
+F.3 Baselines
+TESSERA, PRESTO [149], and GSE [56] embeddings were compared by training identical models using each representation as input, with all other settings held constant.
+Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
+To benchmark performance against traditional remote sensing approaches, we compared the predictions against three global canopy height products: Potapov et al.
+[2021][121] (GLAD), Lang et al.
+[2023][84] (ETH), and Tolan et al.
+[2024][148] (Meta) as well as a regional map produced by Lang et al.
+[2021][85].
+These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
+Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region.
+Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
+This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery.
+Cloud masking and compositing were performed in GEE, resulting in 12 channel input images (band XYZ excluded), with one channel for each Sentinel-2 spectral band.
+Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
+Appendix G Downstream Task: Burned area detection
+G.1 Study Areas and Ground Truth Datasets
+For fire disturbance detection, two 15 × 15 km tiles in California were selected at random from the pool of training samples in the burn scars benchmarking dataset [76], referred to as Burn Area A and Burn Area B.
+Burn Area A comprises a heterogeneous landscape of forest, shrubland, agricultural land, and built infrastructure.
+In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
+Burn Area B experienced a single large fire ’Glass’ ($\sim$27,500 ha) in September 2020 that primarily affected forested land.
+In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
+To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [2].
+The dataset from the Monitoring Trends in Burn Severity (MTBS) program [4] was used to identify burned areas and burn severity.
+Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
+Burn severity–defined as the measure of ecological damage caused by fire [55]–was derived from categorical and continuous severity metrics provided by MTBS.
+To obtain burn severity categories of low, medium and high, as well as continuous burn severity, we used MTBS burn severity annualized dataset as well as individual severity maps for each fire.
+MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
+Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [55].
+Since MTBS datasets are derived from Landsat, they are provided at 30-meter resolution, and were upsampled to 10-meters to match the resolution of TESSERA representations.
+The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [57].
+G.2 Dimensionality Reduction
+We used UMAP (Uniform Manifold Approximation and Projection) [103] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires.
+We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections.
+To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs.
+unburned), land cover class, or burn severity.
+These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
+Each Burn Area contains over two million 10-meter pixels.
+To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
+UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space.
+UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
+Appendix H Downstream Task: Above-ground biomass estimation
+H.1 Dataset
+For the AGB regression task, we used the ground truth data from the Biomassters dataset [111].
+Patch coordinates and collection years [111] were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE) [56].
+For the SpectralGPT baseline, we used the preprocessed satellite data from the Biomassters dataset as input, since the codebase for SpectralGPT [67] did not include a preprocessing pipeline at the time of our analysis.
+H.2 UNet model architecture
+To obtain AGB estimates, we trained a convolutional neural network based on a UNet architecture [128], using TESSERA or GSE embeddings  [56] as input representations.
+The model received 128$\times$256$\times$256 and 64$\times$256$\times$256 input patches for TESSERA and GSE, respectively, where the first dimension denotes the number of channels.
+The network followed a standard UNet structure with two encoder and decoder stages.
+The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling.
+A 1024-channel bottleneck block was used prior to upsampling.
+The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double convolution blocks.
+A final 1×1 convolution produced the single-channel regression output.
+The model had approximately 30 million trainable parameters.
+The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
+Training was performed with the Adam optimizer (learning rate = 1e-4, batch size = 4).
+Finetuning on the full dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours.
+H.3 Baselines
+In addition to GSE, we compared TESSERA’s performance to two other baselines: the Biomassters winner [111] and the SpectralGPT foundation model [67] combined with an UPerNet [179].
+For SpectralGPT, we used the implementation provided in Ref.
+[101], with the Biomassters satellite data as input and an UPerNet as the regression head.
+This setup was found in Ref.
+[101] to outperform other tested foundation models.
+To ensure consistency in the limited-label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
+Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [101] as it resulted in slightly better performance.
+Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
+To evaluate the Biomassters winning model, we downloaded the pre-trained model weights and ran inference on the test set.
+Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model.
+According to the Biomassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days – highlighting the greater time efficiency of TESSERA, which required only five hours.
+Appendix I Downstream Task: Stocking indices in voluntary carbon markets
+I.1 Relative height modeling
+Participants in the voluntary carbon market can use a remotely sensed ”stocking index” to quantify project additionality.
+The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
+To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings.
+This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
+To model a suitable stocking index from the 128-dimensional TESSERA embeddings, 110,000 relative heights RH95 and RH10 from 2022 were extracted from the Sentinel-2 tile 22MCT in Para Sate, Brazil.
+From the 10 m pixel falling at the center of the full-waveform footprint, the 128 TESSERA embeddings were extracted as predictors.
+After data extraction, the relative height differences (dRH = RH95-RH10) were used as the target variable and the 128 TESSERA embeddings as predictors.
+A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
+To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
+The TerraClass Amazonia 2022 land cover map [146] was used to distinguish between forest and non-forest.
+For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
+Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
+I.2 Stocking index baselines
+For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [84], CTrees canopy tree height (5 m) [156], ESA-CCI AGB (100 m) [132], Meta CHM (1 m)  [148] and GTDX AGB (25 m) [42].
+Using the same shapefiles delineating the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
+To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
+After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and R2) directly comparable.
+Appendix J Interactive Habitat Mapping Tool
+Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure 12.
+The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles.
+To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
+The core of the tool is the interactive labeling workflow.
+A user can define custom habitat classes and provide training data by clicking on the map to place labeled points.
+These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier.
+The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view.
+This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise.
+We also provide a batch processing tool for larger regions of interest.
+The tools and their open source code are freely available in our github repository [152].
+Figure 12: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
+We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
+Appendix K Supplementary Table
+The table below compares Tessera along several dimensions with a comprehensive list of remote sensing foundation models [73].
+Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy.
+Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10m resolution (10m).
+It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS).
+Codes: $\checkmark$= Yes, $\times$= No, ?
+= Unknown, p= Potentially true, ..
+= N/A.
+Foundation Model R O G A 10 m ST P CI CR LO MS
+TESSERA (2025) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+GeoKR (2021) [87] $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+CVPRW (2021) [139] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
+GASSL (2021) [14] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
+SeCo (2021) [102] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+DINO-MM (2022) [167] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
+SatMAE (2022) [32] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
+RS-BYOL (2022) [75] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
+GeCo (2022) [88] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
+RingMo (2022) [140] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+RVSA (2022) [163] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
+RSP (2022) [160] $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
+MATTER (2022) [7] $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+CSPT (2022) [190] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+CVPRW (2022) [133] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+BFM (2023) [26] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+TOV (2023) [145] ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
+CMID (2023) [109] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+RingMo-Sense (2023) [185] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
+lal-SimCLR (2023) [122] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+CACo (2023) [100] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SatLas (2023) [18] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+GFM (2023) [104] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Scale-MAE (2023) [126] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+DINO-MC (2023) [174] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+CROMA (2023) [53] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Cross-Scale MAE (2023) [144] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+DeCUR (2023) [166] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
+Presto (2023) [149] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+CtxMIM (2023) [189] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+FG-MAE (2023) [170] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Prithvi (2023) [134] $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+RingMo-Lite (2023) [172] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+IGARSS (2023) [50] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+EarthPTi (2023) [137] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+USat (2023) [72] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+AIEarth (2023) [182] p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Clay (2023) [31] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Hydro (2023) [35] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+U-Barn (2023) [44] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
+GeRSP (2023) [69] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SwiMDiff (2023) [147] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+OFA-Net (2023) [181] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+SML-FR (2023) [40] p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Spectral-GP (2024) [67] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+S2MAE (2024) [90] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SatMAE++ (2024) [113] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+msGFM (2024) [61] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SkySense (2024) [59] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+MTP (2024) [162] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+DOFA (2024) [180] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+MMEarth (2024) [112] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+LeMeViT (2024) [78] $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
+SoftCon (2024) [168] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+RS-DFM (2024) [173] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+A2MAE (2024) [188] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+HyperSIGM (2024) [159] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SelectiveMAE (2024) [164] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+OmniSat (2024) [11] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
+MM-VSE (2024) [125] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
+MA3E (2024) [94] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SpectralEarth (2024) [93] $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SenPa-MAE (2024) [123] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+RingMo-Ae (2024) [38] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SAR-JEPA (2024) [89] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
+PIS (2024) [9] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
+OReole-FM (2024) [39] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+PIEVIT (2024) [98] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SatVisionTOA (2024) [138] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+RS-vHeat (2024) [68] $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
+Prithvi-EO-2.0 (2024) [141] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+AnySat (2024) [12] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+WildSAT (2024) [36] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SeaMo (2024) [92] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+FoMo (2025) [22] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+SatMamba (2025) [43] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Galileo (2025) [150] p p $\checkmark$ ? $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+SatDiFuser (2025)[77] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+RoMA (2025) [165] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Panopticon (2025) [157] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+FedSense (2025) [142] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Copernicus-FM (2025) [171] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+HyperFree (2025) [86] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+DynamicVis (2025) [27] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+FlexiMo (2025) [91] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Google Satellite (2025)[56] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+RingMoE (2025) [19] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$

### v3

@@
 Abstract
-Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including climate modeling, carbon accounting, and strategies for conservation and sustainable land use.
+Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including habitat mapping, carbon accounting, and strategies for conservation and sustainable land use.
 However, satellite time series are voluminous, often corrupted by sensor noise, clouds, and atmospheric conditions, and unevenly spaced in time, making them challenging to use.
-We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time series data.
-TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-60 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps.
+We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time-series data.
+TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-20 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps.
 We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines.
-We believe that TESSERA’s ease of use, openness, computation-, label-, and data-efficiency, and high performance will prove transformative in a wide range of ecological applications.
+We believe that TESSERA’s ease of use, state-of-the-art performance, openness, and computation- and labelled data-efficiency will prove transformative in a wide range of ecological applications.
 Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [117, 185, 25].
-Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise exponential growth in the future [176].
-However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labeled datasets required for supervised training [95].
-The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and nonuniform temporal sampling, preventing full and easy use of this valuable resource [193, 60, 130].
+Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise continued growth in the future [178].
+However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labelled datasets required for supervised training [95].
+The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling, preventing full and easy use of this valuable resource [196, 61, 131].
 EO data users usually mitigate these issues by annual or seasonal compositing [35], then using supervised machine learning to map from composited images to task-specific classification, such as crop identification [9].
-However, compositing removes the critical temporal signal arising from phenology and training a task-specific ML model is computationally onerous.
+However, compositing removes the critical temporal signal arising from phenology (i.e.
+timing of biological events such as flowering) and training a task-specific ML model is computationally onerous.
 Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data [180].
-Their development follows several paths, including supervised pre-training on large labeled datasets for spatio-temporal analysis [21, 20, 160, 18], and various self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
-These SSL techniques include generative methods, such as masked autoencoders [32, 104, 126, 61, 113, 94, 112, 53, 144, 140, 134, 67, 149], diffusion-based models [192, 80, 115], and contrastive methods that learn efficient representations by comparing data samples [14, 102, 7, 99, 81, 23, 59, 70, 169].
-Subsequently, these general purpose models are adapted and fine-tuned for specific remote sensing tasks such as high-resolution (semantic) segmentation [161, 177, 184, 28, 191].
-Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabeled data through auxiliary tasks that reveal latent patterns [15, 97], they still do not address the loss of the temporal signal during compositing.
-In recent work, Lisaius et al [95] have demonstrated that the Barlow Twins loss [187], grounded in redundancy reduction principles [17], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series.
+These include both supervised pre-training on large labelled datasets for spatio-temporal analysis [22, 21, 162, 19], and self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
+Common SSL techniques include generative methods, such as masked autoencoders [33, 136, 68, 151], diffusion-based models  [195, 81, 114], and contrastive methods [15, 60, 71, 171].
+General purpose foundation models must be adapted (or ‘fine-tuned’) for specific downstream tasks such as habitat mapping and semantic segmentation [163, 179, 186, 29, 193].
+Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabelled data through auxiliary tasks that reveal latent patterns [16, 97], they still do not address the loss of the temporal signal during compositing.
+In recent work, Lisaius et al.
+[95] have demonstrated that the Barlow Twins loss [189], grounded in redundancy reduction principles [18], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series.
 Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks.
-Our experiments demonstrate that these TESSERA-derived embeddings outperform baseline methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, and identifying fire scars in California.
+Our experiments demonstrate that these TESSERA-derived embeddings outperform state-of-the-art methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, mapping above-ground biomass in Finland, and identifying fire scars in California.
 Importantly, our outputs are ‘ready to use’: users need neither to fine-tune the model, nor download and process voluminous satellite data, nor perform inferencing other than what is needed to perform a downstream task.
-This makes TESSERA frugal in terms of both data handline and computation.
+This makes TESSERA frugal in terms of both data handling and computation.
 They can also keep their data private instead of having to upload it to a centralized repository.
 This removes a crucial barrier to use.
-We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training are needed to achieve state-of-the-art performance.
+We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training is needed to achieve state-of-the-art performance.
 Moreover, our model is open source, enabling extensibility and scientific reproducibility.
-We have made a gap-free, global, annual, 10-meter resolution embedding map for 2024 openly available through the Geotessera Python library [152] and aim to eventually provide global maps for the entire period 2017-2024.
+We have made a gap-free, global, annual, 10 m resolution embedding map for 2024 openly available through the Geotessera Python library [154] and aim to eventually provide global maps for the entire period 2017-2024.
 These information-rich precomputed maps greatly simplify a wide range of downstream EO applications.
@@
 1 Overview of the TESSERA System
-TESSERA uses self-supervised learning to create per-pixel embeddings from the unlabeled time-series of observations from Sentinel-1 (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams.
-It comprises two main components: a dual-branch encoder and a projector network (see Fig.
+TESSERA uses self-supervised learning to create per-pixel embeddings from a global dataset of unlabelled time-series of observations from Sentinel-1 Synthetic Aperture Radar (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams.
+It is trained by optimizing a modified Barlow Twins loss function [189].
+The original Barlow Twins approach seeks to create embeddings that are a) invariant to augmentations of its input, where these augmentations are typically rotations, reflections, and cropping of an image and b) whose dimensions are as decorrelated as possible.
+Inspired by the work by Lisaius et al.
+[95], we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions.
+This is achieved by randomly sampling non-cloudy days from the Sentinel-2 annual observation time series so that model then learns to reconcile these partial views, interpolating missing data.
+The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Appendix A.7.2 for details, Equation 11).
+We now discuss the system architecture.
+TESSERA comprises two main components: a dual-branch encoder and a projector network (see Fig.
 1 for details).
 TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps).
-This format preserves the temporal evolution of spectral/backscatter signatures inherent in the satellite data.
-Our design accommodates common observational gaps and irregularities, such as those due to cloud cover or orbit overlap, by masking invalid rows, which are handled during subsequent processing stages.
-Model training and inferencing is done by sparse temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig.
+This format preserves the temporal evolution of the spectral/backscatter signatures inherent in the satellite data, while allowing observational gaps and irregularities, such as those due to cloud cover or orbit overlap.
+Rows that are cloud-corrupted or otherwise invalid are indicated by a binary mask and ignored when creating augmentations.
+Model training and inferencing is done by random temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig.
 1 for details).
+In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
 For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics.
-Under typical mid-latitude observation conditions, this compresses input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
+Under typical mid-latitude observation conditions, this compresses the input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
 Figure 1: Architecture and data processing pipeline of the TESSERA foundation model.
 a, Overview of the end-to-end workflow.
-Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR are first extracted for each pixel location, forming modality-specific ‘d-pixels’.
-For self-supervised training, two distinct augmentations are created from each d-pixel–each augmentation is a random cloud-free sample from an annual time series of Sentinel-2 observations along with the corresponding Sentinel-1 data–processed through a dual-branch architecture with weights shared between the branches.
-The resulting 128-dimensional embeddings from both modalities are fused by a multilayer perceptron (MLP) to create the output embeddings of the system.
-The fused features are then expanded to 16,384 dimensions by a large projector network.
-A modified Barlow Twins loss ($\mathcal{L}_{BT}+\mathcal{L}_{MIX}$) is calculated on the cross-correlation matrix of these projected features, optimizing the model to learn augmentation-invariant and decorrelated embeddings that are internally decorrelated, and fairly invariant to obscured days in the Sentinel-2 time series.
-b, Detailed structure of a single encoder branch.
-A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded.
-A Day-of-Year-based temporal encoding is added to these embeddings.
-The resulting sequence is then processed by an 4-block Transformer Encoder.
+Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are first extracted for each pixel location, to form mode-specific ‘d-pixels’ (shown as vertical rectangle, with cloud-corrupted rows indicated in gray).
+To allow self-supervised training, two augmentations are created from each d-pixel.
+Each augmentation is a cloud-free random sample from the corresponding d-pixel.
+Augmentations are processed through a dual-branch architecture with weights shared between the branches, which is the key feature of Barlow Twin SSL.
+The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings.
+During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network.
+A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of these projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked (i.e, cloud-corrupted) days in the Sentinel-2 d-pixel.
+k b, Detailed structure of a single encoder branch.
+A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded using a single-layer MLP (not shown).
+A Day-of-Year-based temporal encoding is added to this initial embedding.
+The resulting sequence is processed by an 4-block Transformer Encoder.
 Finally, a Gated Recurrent Unit (GRU) pooling layer processes the sequence of temporal features; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
 TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared).
-Each encoder processes the temporal sequence of a d-pixel, utilizing multi-head self-attention mechanisms to capture complex temporal dependencies and patterns specific to each data modality.
+Each encoder processes a d-pixel, utilizing multi-head self-attention mechanisms to capture the temporal dependencies and patterns specific to each data modality.
 Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context.
@@
 See Appendix A.7 and Figure 1 for architectural details.
-Modality-specific embeddings derived from dual encoders are first fused using a multilayer perceptron (MLP) which is trained alongside the encoders.
+Modality-specific embeddings derived from dual encoders are fused using a multilayer perceptron (MLP) which is trained alongside the encoders.
 This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network.
-This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate effective redundancy reduction during the self-supervised learning phase.
+This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate redundancy reduction during the self-supervised learning phase.
 A detailed breakdown of the model’s architectural parameters is provided in Appendix A.6 (Table 2).
+We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [1].
 1.1 Model Training
-TESSERA learns the mapping of inputs to embeddings through a self-supervised pretraining process on a large-scale, unlabeled global dataset.
-It is trained by optimizing a modified Barlow Twins loss function [187, 95].
-The core principle is to learn embeddings that are invariant to differing views of the same pixel’s optical-band data (e.g., different subsets of cloud-free observation dates) while minimizing redundancy between the learned feature dimensions.
-This is achieved by generating two distinct augmented views for each input d-pixel, through independent sparse temporal samples.
-This avoids creating augmentations by applying artificial (e.g.
-hand-crafted) distortions.
-The model then learns to reconcile these partial views.
-The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and intersample interaction during training (see Appendix A.7.2 for the full loss function, Equation 11).
-In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
-The pretraining phase utilizes a vast global dataset comprising approximately 800 million d-pixels, down-sampled from Sentinel-1 and Sentinel-2 imagery (we use a downsampling ratio of 400:1).
+The pretraining phase utilizes a large global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series.
 To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed.
 This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features.
-The model was trained for one epoch on this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy and optimization).
+The model was trained for one epoch in this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy, and optimization).
 The key hyperparameters used during pretraining are summarized in Appendix A.7.3 (Table 3).
 1.2 Model Inference and Global Embeddings Maps
-Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10-meter pixel globally.
+Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally.
 This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders.
 To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations.
-The primary output of this inference stage is a set of annual, global, 10-meter resolution TESSERA embedding maps.
+The primary output of this inference stage is a set of annual, global, 10 m resolution TESSERA embedding maps.
 These maps are designed as readily usable multichannel geospatial data layers.
@@
 250 TB) and will progressively extend coverage backward year by year until 2017.
+Visualizing the first three dimensions of the embeddings as R,G, and B channels reveals rich, detailed textures corresponding to diverse land cover and land use patterns.
+These visualizations highlight the model’s ability to capture information across multiple scales, from individual agricultural fields to entire continents (Fig.
+2).
+N
+Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales.
+The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization.
+a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns.
+b, A multi-scale view of a complex agricultural landscape.
+b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields.
+c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
 2 Downstream Tasks
 The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks.
-These include, but are not limited to, pixel-wise classification (e.g., crop type mapping), pixel-wise regression (e.g., canopy height and AGB estimation) and patch-based dense predictions (for example, semantic segmentation).
-In these evaluations, the pre-trained TESSERA encoder acts as a fixed feature extractor; given adequate computing resources, fine-tuning of the model is also possible.
-Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these fixed embeddings.
-This approach typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state of the art.
-See Appendix D and Supplementary Fig.
-10 for the general downstream application workflow.
-We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [1].
+These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation.
+(See Appendix D and Supplementary Fig.
+12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings.
+Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling.
+This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
 2.1 Crop Type Classification
@@
 This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3].
-We compared the performance of a lightweight classification head trained on TESSERA embeddings against three key baselines: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [149], and (3) Google Satellite Embedding (GSE) [56].
-For the foundation models, we applied the same lightweight head to their respective embeddings for a direct comparison.
+We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [151], and (3) Google Satellite Embedding (GSE) [57].
+For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison.
 We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning.
@@
 Details on data processing, model architectures, and training protocols are provided in Appendix E.
-TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labeled data are scarce (Fig.
-2).
+TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labelled data are scarce (Fig.
+3).
 In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig.
-2a).
+3a).
 TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig.
-2b).
+3b).
 Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig.
-2c).
+3c).
 These performance gains are due to the high quality of the learned embeddings.
-A UMAP analysis [103] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from both PRESTO and GSE (Fig.
-2d).
+A UMAP analysis [102] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig.
+3d).
 This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
@@
 Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability.
-GEDI, a spaceborne LiDAR system, can provide reasonably accurate canopy height measurements to train models to generate CHMs from optical and radar data; however, it does not provide wall-to-wall maps, and does not sample high latitudes [41], and produces unreliable estimates in mountainous terrain [52].
+Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [23].
+GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [42], and does not produce unreliable estimates in mountainous terrain [53].
 Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [132].
-Optical data, while widely available, lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [158, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 121, 148].
+Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [160, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 120, 150].
 Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al.
@@
 These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality.
-As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, and current products underestimate local variability and structural extremes [107].
+As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [106].
 Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
@@
 The accuracy metrics for the foundation models are reported as the average across the 12 runs.
-TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
+TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
 We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig.
-3).
+4).
 First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.
@@
 [150].
-Although local models generally outperform global products in regional evaluations [127], the pronounced saturation of these global models at much lower canopy heights - combined with RMSEs that exceed their reported global averages [121, 84, 148]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
-We also compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
-[85], which used deep convolutional neural networks trained on Sentinel-2 time series and GEDI LiDAR to estimate regional canopy height and biomass.
+Although local regression models, such as ours, generally outperform global products in regional evaluations [128], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [120, 84, 150]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
+We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
+[85], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass.
 Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [151] and Google Satellite Embedding (GSE) [57], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
@@
 Instead, only the region and year of interest are needed to replicate the input data pipeline.
-By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
+By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
 Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation.
@@
 A 1:1 reference line (blue) is shown in each panel for comparison.
-2.3 Above-ground biomass estimation
-Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [66].
-Accurate estimation of above-ground biomass (AGB) is essential to quantify forest carbon stocks and monitor temporal changes due to deforestation, degradation, and recovery [45].
-We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset.
-Firstly, we considered the geospatial foundation model SpectralGPT [67] combined with a UPerNet regression head – a setup found in the Pangaea benchmark [101] to outperform other foundation models tested on the BioMassters dataset [111].
-As the preprocessing pipeline is not yet available in the codebase of Ref.
-[67] at the time of our analysis, we adopted the implementation from Ref.
-[101], which uses the preprocessed satellite imagery from the BioMassters dataset as input.
-Secondly, we used the winning model from the BioMassters competition as our state-of-the-art supervised baseline – a UNet with a temporal attention encoder [111], which achieved the lowest RMSE among over 1,000 submissions.
-The BioMassters dataset comprises 11,462 image patches that represent above-ground biomass in Finnish forests from 2017 to 2021.
-Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey.
+2.3 Above-ground forest biomass estimation
+Forests play a central role in the global carbon cycle [67].
+This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [46].
+A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions.
+Thus, there has been a long-standing effort to link field measurements to global satellite observations [194].
+Recently, with the creation of the BioMassters benchmark dataset [110] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [5].
+We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98.
+In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [68] and GSE [57].
+First, a summary of the BioMassters dataset.
+This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [126, 127] informed by extensive field plots.
 Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel.
-The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
-A comparison of RMSE across varying training label fractions (Fig.
-4a) shows that a UNet trained on the TESSERA embeddings consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set.
-These results underscore the robustness of TESSERA in sparse-label regime.
-Moreover, TESSERA achieves results comparable to those of the supervised baseline trained on the full dataset, even with limited label availability.
-In general, bespoke models are inherently difficult to surpass in fully supervised settings.
-It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input.
+The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig.
+5b) that would not appear in coarser resolution data.
+To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
+We compared a UNet-based regression model [129] trained on TESSERA embeddings with SpectralGPT embeddings [68] combined with a lightweight spatial-temporal encoder [133] for temporal aggregation and a UPerNet [181] regression head, because this combination has been found in the Pangaea benchmark [100] to outperform other foundation models [110, 100].
+Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase.
+Thus, we adopted the Pangaea implementation of their model [100], which uses preprocessed satellite imagery from the BioMassters dataset.
+We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline.
+This model uses a UNet with a temporal attention encoder [110].
+Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input.
 However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
-Figure 4: TESSERA embeddings demonstrate robustness for AGB mapping under limited label availability.
-a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGPT.
-Following the evaluation protocol in Ref.
-[111], RMSE is computed on a fixed test set of 2773 patches and reported as the average RMSE per patch.
-For limited labels testing, a subset of labels is randomly chosen from the remaining data, and randomly split into training and validation sets (80:20).
-The figure shows the average RMSE computed over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
-The UNet model is trained on the TESSERA embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
-The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability.
-b, Predicted AGB map for an example patch using 10% of the training labels.
-c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset.
-The top panel depicts the histogram of the ground-truth AGB values.
-d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top), the supervised baseline (middle) and SpectralGPT+UPerNet (bottom).
+Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
+Our performance metrics are the mean RMSE, the coefficient of determination ($R^{2}$), and the mean bias (MB) in AGB across the test patches as a function of the fraction of labelled data used to train the regressor (Fig.
+5a).
+We make three significant findings.
+First, although custom models developed using supervised learning are typically difficult to outperform, even with a small fraction of training labels, our regressor nearly matches the performance of the BioMassters competition winner, which uses a supervised learning approach.
+Second, our regressor trained on 1% of the labelled data outperforms a UPerNet-based regressor trained on the entire training set using SpectralGPT embeddings.
+Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
+Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings.
+a, TESSERA achieves lower RMSE values than SpectralGPT+UPerNet and results similar to the state-of-the-art supervised baseline and Google Satellite Embedding foundation model (top row).
+The coefficient of determination ($R^{2}$) and mean bias (MB) are shown in the middle and bottom row as functions of label availability.
+b, Mean predicted AGB values show increasing bias at higher ground-truth AGB values, which, however, are sparse in the dataset (the y-axis in the upper panel is on a logarithmic scale).
+Mean predictions are calculated for ground-truth AGB values grouped into 5 t/ha intervals.
+c, Predicted AGB map for an example patch using 4% of labels in training.
+d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (left) and the supervised baseline (right).
+The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
 2.4 Burned Area Detection
@@
 Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82].
-Most existing burn area mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 119].
-These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), though many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82].
+Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 118].
+These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82].
 Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [65], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [66].
 SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 145, 156].
-Recent reviews emphasize the need for sensor fusion, combining passive and active sensors to improve accuracy and robustness [82, 30].
-Sentinel-1 and 2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [143, 154, 129].
-TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig.
-5).
-Figure 5: TESSERA embeddings capture wildfire extent, timing, and severity across two burned regions.
-(a–c) Burn Area A, which experienced two distinct fires in June and August 2021.
-(a) Time of fire map based on MTBS, showing unburned areas (blue), the June fire (orange), and the August fire (red).
-(b) Burn severity classification for the same region, with unburned (blue), low (green), moderate (orange), and high (red) severity.
-(c) UMAP projection of TESSERA embeddings from 2021, with colors corresponding to burn severity.
-Embeddings form well-separated clusters for unburned, June-, and August-burned areas, with a visible internal gradient reflecting severity within August fire.
-(d–f) Burn Area B, which experienced a single large fire in September 2021.
-(d) Time of fire map showing unburned (blue) and burned (green) areas.
-(e) Burn severity classification showing heterogeneous severity across the landscape.
-(f) UMAP projection for Burn Area B, revealing fuzzier boundaries between burned and unburned pixels, though the severity gradient remains evident.
-We tested whether TESSERA’s temporal embeddings capture fire-induced changes in spectral and radar signals.
-We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 $\times$ 15 km regions from the Monitoring Trends in Burn Severity (MTBS) dataset [4], each affected by wildfires in California during 2021.
-TESSERA embeddings for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels.
-Fire boundaries and burn severity data were sourced from the dataset.
-The resulting UMAP projections demonstrated strong separability in two-dimensional space, visually highlighting TESSERA’s capacity to detect disturbances at the pixel level.
-Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single large fire that affected mostly forested areas, resulting in a heterogeneous severity of the burn.
-The UMAP revealed separability across three domains (Fig.
-5):
-  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned, with a clearer separation in Burn Area A.
-  • Temporal fire differentiation: In Burn Area A, embeddings separated pixels burned in June vs. August, suggesting temporal sensitivity to disturbance events.
-  • Burn severity gradient: In both regions, a gradient structure in the UMAP space corresponded to burn severity, indicating that TESSERA encodes the magnitude of the disturbance.
-Because TESSERA encodes an annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images.
-This capacity supports a simplified, scalable pipeline for monitoring large-area disturbances.
+Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [82, 31].
+Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [145, 156, 130].
+TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig.
+6).
+Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California.
+a, UMAP projection of 2020 TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events.
+A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance.
+b, Burned vs.
+unburned classification performance using random forest classifiers trained on different input representations.
+TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google Satellite Embeddings (GSE).
+c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas.
+As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
+We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6).
+Our analysis focused on two randomly selected 15$\times$15 km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [4].
+UMAP projections [102] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters.
+In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a).
+To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels.
+We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites.
+The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b).
+Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency.
+These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
+To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps.
+These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Figure 6C).
+The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
 2.5 Stocking indices in voluntary carbon markets
-The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration while minimizing leakage liabilities [110].
+The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [109].
 To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
 This is a costly and inefficient undertaking if it is done only by field sampling.
-To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [155]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers merely need to demonstrate that the chosen stocking index is correlated with the AGB.
-Upon acceptance, project developers can use the index to quantify the additionality of the project and to assess the performance of the project.
+To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [157]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB.
+Upon acceptance, project developers can use the index to assess the performance of the project.
 Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
-To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise RF regression following the methodology outlined in Atzberger et al.
+To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.
 [14].
@@
 Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market.
-Comparison of in-situ AGB measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details).
+Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details).
 Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 X + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA AGB converted using 0.407 X + 34.0; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X AGB prediction converted using 0.3 X + 54.3.
@@
 3 Conclusion
-In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
+In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
 TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing.
@@
 Acknowledgments
-We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
+We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
 Funding:
@@
 Author contributions:
-TESSERA conceptualization: CA
+Initial conceptualization: CA
 System design: CA, ZF, SJ, SK
@@
 Paper draft: ZF, MCL, RY, SS, SK, JK
-Paper review and refinement: AB, CA, DAC, SJ, SK
+Paper review and refinement: All
 Competing interests:
@@
 The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [3].
-The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) datasest [4].
+The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset [4].
 The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [34].
 The AGB downstream task uses the open BioMassters dataset [110].
-The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners, hence data access will require a Data Transfer Agreement.
+The code for the winning model of the BioMassters competition is described in [110] and freely available on the corresponding GitHub page.
+The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark [100].
+The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners; hence, data access will require a Data Transfer Agreement.
 Appendix A Model details
@@
 In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time.
-Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there is valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
+Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there are valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
 Table 1: Variable definitions for remote sensing data structure
@@
 A.4 The d-pixel Representation
-We represent each 10-m pixel in the time series of images from annual multi-spectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
+We represent each 1- m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
 1 a).
@@
 This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities.
-Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored during attention computation or ignored in the temporal sampling process.
+Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored in the temporal sampling process.
 A.5 Dual-Encoder Architecture
 Given the distinct nature of Sentinel-1 SAR and Sentinel-2 MSI data, TESSERA employs two separate, parallel transformer-based encoder branches.
-  • Sentinel-2 MSI Encoder: This branch processes a time series of 10 spectral bands (excluding the 60 m bands used for the detection of water vapor and clouds) from Sentinel-2.
+  • Sentinel-2 MSI Encoder: This branch processes a time series of 10 spectral bands (excluding the 60 m bands used for the detection of water vapour and clouds) from Sentinel-2.
 We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
@@
 The fused embedding from the dual-encoder stage is subsequently fed into a large projector network.
-This projector is a five-layer MLP, with each hidden layer having 16,384 dimensions.
+This projector is a six-layer MLP.
+Its architecture comprises an input layer mapping the fused embeddings to 16,384 dimensions, followed by four hidden layers of 16,384 dimensions each, and a final linear output layer.
+Each of the first five layers is a fully-connected linear layer followed by Batch Normalization and a ReLU non-linear activation function, making the network deep and highly non-linear.
 This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [189], to enable effective redundancy reduction during the self-supervised loss computation.
-The final output of the projector for each input d-pixel is an embedding, which is then used in the loss calculation.
-For downstream tasks, we use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding.
+The final output of the projector is then used in the loss calculation.
+For downstream tasks, we used the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding.
 The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
@@
   1. 1.
-    For each view, independently sampling a fixed number of valid observation dates from the annual Sentinel-2 time series (10 spectral bands).
+    For each view, independent sampling of a fixed number of valid observation dates from the annual Sentinel-2 time series (10 spectral bands).
   2. 2.
-    For each view, independently sampling a fixed number of valid observation dates from the annual Sentinel-1 time series (2 polarizations).
+    For each view, independent sampling of a fixed number of valid observation dates from the annual Sentinel-1 time series (2 polarizations).
 These views represent different, valid, but inherently incomplete glimpses of the pixel’s true temporal-spectral evolution, akin to observing the same location through intermittent cloud cover or from different satellite passes at different times.
@@
 [17].
-This involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating mixed views $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$, where $\alpha_{mix}\sim\text{Beta}(\beta_{p},\beta_{p})$.
-The embeddings $Z_{M}$ and $Z_{S}$ are obtained.
+For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$.
+The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution on the unit interval, $\alpha_{mix}\sim U(0,1)$.
+The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views.
 The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
@@
 $\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (10)
-where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model’s outputs.
-The total loss function optimized during the training of TESSERA is a weighted sum:
+where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model’s outputs. The total loss function optimized during training is a weighted sum:
 $\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
@@
 The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles.
-d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400, i.e, selecting one pixel from each 20$\times$20 patch.
+d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400.
+This was achieved by systematically taking every 20th pixel along the horizontal and vertical axes.
 Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
@@
 Weight Decay $\lambda_{wd}$ $1\times 10^{-6}$
-Learning Rate Schedule - Linear warmup (10%), plateau (20%), cosine decay
+Learning Rate Schedule - Linear warmup (10%) followed by cosine decay
 Gradient Clipping Norm - 2.0
@@
 Mixup Lambda $\lambda_{mix}$ 1.0
-Mixup Beta Parameter $\beta_{p}$ 1.0
 Architecture & Data
@@
 Quantization Bits (QAT) - 8
-Figure 7: TESSERA is trained on globally-distributed training data with greater data density in the northern hemisphere.
+Figure 8: TESSERA is trained on globally-distributed training data.
 TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024.
-The color of each hexagon in the map corresponds to the number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate color legend.
+The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend.
 This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
@@
 AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$.
-The learning rate schedule included a linear warming during the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay.
+The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training.
 The global batch size was 32768.
+Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary Fig.
+9.
+Figure 9: Training dynamics curves show that TESSERA rapidly reaches its asymptotes.
+a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss.
+b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay.
+c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction.
+d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant.
+e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
 A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
@@
 This pipeline is conceptually illustrated in Fig.
-8a.
+10a.
 The impact of this approach on training stability is empirically demonstrated in Fig.
-8b.
-Compared to a conventional, localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
+10b.
+Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
 This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
@@
 a, Schematic of the data shuffling and loading process.
-D-pixels (colored squares) from thousands of MGRS tiles are first aggregated into a global pool.
+D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool.
 A custom Rust binary [154] performs a global shuffle on this multi-terabyte dataset before applying augmentations.
@@
 A.8 Global Embedding Map Generation
-A primary output of the TESSERA project is the generation of annual global embedding maps with 10-meter resolution for the years 2017-2024.
+A primary output of the TESSERA project is the generation of annual global embedding maps with 1- meter resolution for the years 2017-2024.
 To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used.
-For each 10-meter pixel on the globe and for each year:
+For each 1- meter pixel on the globe and for each year:
   1. 1.
-    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
+    The full Sentinel-1 and Sentinel-2 time series data at 1- meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
   2. 2.
@@
     The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional embedding vector for that pixel for that year.
-This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
+This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 1- meter grid.
 Appendix B Scaling with data and network size
@@
 Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
-9).
+11).
 Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
@@
 In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning.
-We performed a parameter sweep over the projector’s depth (number of layers) and width (number of neurons per layer), with results shown in Fig.
-9a.
+As detailed in Section A.6, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers.
+Our parameter sweep focused on this computationally intensive hidden block.
+We varied the number of these hidden layers (referred to as ’depth’ in Fig.
+11a) and their width (number of neurons per layer), with results shown in Fig.
+11a.
 As discussed by Zbontar et al.
 [189], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task.
-The highest validation score was achieved with the largest configuration tested (8 layers deep, 32,768 neurons wide).
+The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide).
 However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure.
-For the final TESSERA v1.0 model, we selected the configuration with 4 layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
-Furthermore, we analyzed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
-9b.
+For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
+Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
+11b.
 The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
@@
 a, Hyperparameter sweep for the projector network’s architecture.
-Each point represents a configuration, with its position determined by depth (x-axis) and width (y-axis).
-The color and size of the point indicate the validation F1 score and the number of parameters, respectively.
+Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis).
+The colour and size of the point indicate the validation F1 score and the number of parameters, respectively.
 Performance generally increases with both width and depth.
@@
 Appendix C Ablation study
-To validate the contribution of each key component within the TESSERA framework, we conducted an ablation study.
+To validate the contribution of each key component within the TESSERA framework, we performed an ablation study.
 We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
@@
 Removing mixup regularization also severely degrades results.
-When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable, robust representations.
+When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
 With removal of Sentinel-1 data, task performance decreases moderately, demonstrating some gain from the fusion of optical and SAR data yields.
@@
 Appendix D Downstream Task Application Methodology
-A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data.
-Having pre-trained TESSERA on large unlabeled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications.
+A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data.
+Having pre-trained TESSERA on large unlabelled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications.
 The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
@@
   2. 2.
-    Prepare Labeled Downstream Data: The labeled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
+    Prepare labelled Downstream Data: The labelled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
   3. 3.
@@
 Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used.
-This training typically requires significantly less labeled data and computational power compared to training a deep model from scratch.
+This training typically requires significantly less labelled data and computational power compared to training a deep model from scratch.
   5. 5.
@@
 The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector.
-The RF model consisted of 200 trees, with other parameters set to standard values.
+The RF model was configured with 100 trees, while all other hyperparameters retained their default values as implemented in the scikit-learn library.
 For the experiments shown in the main text’s figure, we trained these models on randomly selected subsets of the data (from 1% to 30% for panel a; specified samples-per-class for panel b), using a fixed validation set for hyperparameter tuning and a held-out test set for final evaluation.
@@
 Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output.
-We did not freeze the encoders during the finetuning.
+We did not freeze the encoders during the fine-tuning.
 Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 13.
@@
 UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
-In this work, the entire embeddings map (e.g., an array of shape $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for TESSERA, PRESTO, and GSE embeddings.
+In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
 To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
@@
     $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (13)
-    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighboring cluster.
+    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighbouring cluster.
 The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
@@
     where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$.
-Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where clusters are compact and well separated.
+Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
 Figure 13: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
@@
 The first column shows the ground truth segmentation.
-For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labeled ’Embedding’) and the final predicted segmentation map (labeled ’Predicted’).
+For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
 Appendix F Downstream Task: Canopy height estimation
@@
 Canopy height prediction was evaluated in a 5$\times$6 km area in the Danum Valley, Borneo.
-This region is covered in tropical rainforest dominated by the fast-growing dipterocarp family, which hosts some of the tallest trees in the tropics [135, 120].
+This region is covered in tropical rainforest dominated by species in the Dipterocarpaceae family, which hosts some of the tallest trees in the tropics [137, 119].
 Ground truth data was derived from airborne laser scanning (ALS) performed in February 2020 [34].
@@
 These canopy height rasters were then downsampled using mean pooling to 10 m to match the resolution of the TESSERA representations.
-F.2 Modeling Methods
+F.2 Modelling methods
 To predict the canopy height from the different embeddings, we used a standard U-Net architecture [129].
@@
 Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
-To benchmark performance against traditional remote sensing approaches, we compared the predictions against three global canopy height products: Potapov et al.
+To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al.
 [2021][120] (GLAD), Lang et al.
@@
 These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
-Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region.
+Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
 Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
 This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery.
-Cloud masking and compositing were performed in GEE, resulting in 12 channel input images (band XYZ excluded), with one channel for each Sentinel-2 spectral band.
+Cloud masking and compositing were performed in GEE, resulting in 12 channel input images, with one channel for each Sentinel-2 spectral band.
 Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
@@
 G.1 Study Areas and Ground Truth Datasets
-For fire disturbance detection, two 15 × 15 km tiles in California were selected at random from the pool of training samples in the burn scars benchmarking dataset [76], referred to as Burn Area A and Burn Area B.
-Burn Area A comprises a heterogeneous landscape of forest, shrubland, agricultural land, and built infrastructure.
+For fire disturbance detection, two 15$\times$15 km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [77], referred to as Burn Area A and Burn Area B.
+Each Burn Area contains over two million 1 m pixels.
+Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure.
 In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
-Burn Area B experienced a single large fire ’Glass’ ($\sim$27,500 ha) in September 2020 that primarily affected forested land.
+Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land.
 In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
 To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [2].
-The dataset from the Monitoring Trends in Burn Severity (MTBS) program [4] was used to identify burned areas and burn severity.
+The dataset from the Monitoring Trends in Burn Severity (MTBS) program [4] was used to identify burned areas and severity of the burn.
 Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
-Burn severity–defined as the measure of ecological damage caused by fire [55]–was derived from categorical and continuous severity metrics provided by MTBS.
-To obtain burn severity categories of low, medium and high, as well as continuous burn severity, we used MTBS burn severity annualized dataset as well as individual severity maps for each fire.
+Burn severity, defined as the measure of ecological damage caused by fire [56], was derived from categorical and continuous severity metrics provided by MTBS.
+To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire.
 MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
 Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [56].
-Since MTBS datasets are derived from Landsat, they are provided at 30-meter resolution, and were upsampled to 10-meters to match the resolution of TESSERA representations.
+Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations.
 The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [58].
+For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high).
+The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
 G.2 Dimensionality Reduction
@@
 These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
-Each Burn Area contains over two million 10-meter pixels.
 To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
@@
 UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
+G.3 Label Efficiency Analysis and Baselines
+We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available.
+The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set.
+We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%.
+The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance.
+Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
+We compared three types of input embeddings for the classification task.
+The first was TESSERA, using embeddings for 2020.
+The second was a task-specific composite baseline, constructed using prior knowledge of the fire event.
+This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector.
+The third type of input was Google Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [57].
+GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
 Appendix H Downstream Task: Above-ground biomass estimation
@@
 For the AGB regression task, we used the ground truth data from the BioMassters dataset [110].
-Patch coordinates and collection years [111] were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE) [56].
-For the SpectralGPT baseline, we used the preprocessed satellite data from the Biomassters dataset as input, since the codebase for SpectralGPT [67] did not include a preprocessing pipeline at the time of our analysis.
+This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a 2,560$\times$2,560  m area at 10 meter pixel resolution.
+Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre.
+Patch coordinates and collection years–provided by the authors of Ref.
+[110] upon request–were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE) [57].
 H.2 UNet model architecture
-To obtain AGB estimates, we trained a convolutional neural network based on a UNet architecture [128], using TESSERA or GSE embeddings  [56] as input representations.
-The model received 128$\times$256$\times$256 and 64$\times$256$\times$256 input patches for TESSERA and GSE, respectively, where the first dimension denotes the number of channels.
+To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [129], using TESSERA or GSE embeddings  [57] as input.
+The model received 256$\times$256$\times$128 and 256$\times$256$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
 The network followed a standard UNet structure with two encoder and decoder stages.
@@
 A 1024-channel bottleneck block was used prior to upsampling.
-The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double convolution blocks.
+The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double-convolution blocks.
 A final 1×1 convolution produced the single-channel regression output.
@@
 The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
-Training was performed with the Adam optimizer (learning rate = 1e-4, batch size = 4).
-Finetuning on the full dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours.
+Following the evaluation protocol in Ref.
+[110], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch.
+For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20).
+The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
+Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
+Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
+Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
 H.3 Baselines
-In addition to GSE, we compared TESSERA’s performance to two other baselines: the Biomassters winner [111] and the SpectralGPT foundation model [67] combined with an UPerNet [179].
+In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [110] and the SpectralGPT foundation model [68] combined with a UPerNet [181].
 For SpectralGPT, we used the implementation provided in Ref.
-[101], with the Biomassters satellite data as input and an UPerNet as the regression head.
-This setup was found in Ref.
-[101] to outperform other tested foundation models.
-To ensure consistency in the limited-label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
-Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [101] as it resulted in slightly better performance.
+[100], with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE [133] for temporal aggregation, and a UPerNet as the regression head.
+This setup was found in [100] to outperform other tested foundation models on the BioMassters dataset.
+To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
+Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [100] because it resulted in slightly better performance.
 Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
-To evaluate the Biomassters winning model, we downloaded the pre-trained model weights and ran inference on the test set.
+For evaluating the supervised baseline, we downloaded the pre-trained weights of the BioMassters winning model and ran inference on the test set, applying a post-processing step to clip predictions to a non-negative range.
 Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model.
-According to the Biomassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days – highlighting the greater time efficiency of TESSERA, which required only five hours.
+According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
 Appendix I Downstream Task: Stocking indices in Voluntary Carbon Markets
-I.1 Relative height modeling
-Participants in the voluntary carbon market can use a remotely sensed ”stocking index” to quantify project additionality.
+I.1 Relative height modelling
+Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events.
 The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
@@
 This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
-To model a suitable stocking index from the 128-dimensional TESSERA embeddings, 110,000 relative heights RH95 and RH10 from 2022 were extracted from the Sentinel-2 tile 22MCT in Para Sate, Brazil.
-From the 10 m pixel falling at the center of the full-waveform footprint, the 128 TESSERA embeddings were extracted as predictors.
-After data extraction, the relative height differences (dRH = RH95-RH10) were used as the target variable and the 128 TESSERA embeddings as predictors.
+To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [42] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil.
+From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables.
 A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
+The TerraClass Amazonia 2022 land cover map [148] was used to distinguish between forest and non-forest.
 To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
-The TerraClass Amazonia 2022 land cover map [146] was used to distinguish between forest and non-forest.
+I.2 Stocking index baselines
 For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
 Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
-I.2 Stocking index baselines
-For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [84], CTrees canopy tree height (5 m) [156], ESA-CCI AGB (100 m) [132], Meta CHM (1 m)  [148] and GTDX AGB (25 m) [42].
-Using the same shapefiles delineating the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
+For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [84], CTrees canopy tree height (5 m) [158], ESA-CCI AGB (100 m) [134], Meta CHM (1 m)  [150], and GTDX AGB (25 m) [43].
+Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
 To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
@@
 To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
-The core of the tool is the interactive labeling workflow.
-A user can define custom habitat classes and provide training data by clicking on the map to place labeled points.
+The core of the tool is the interactive labelling workflow.
+A user can define custom habitat classes and provide training data by clicking on the map to place labelled points.
 These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier.
@@
 Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy.
-Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10m resolution (10m).
+Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m).
 It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS).
@@
 SatMamba (2025) [44] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Galileo (2025) [150] p p $\checkmark$ ? $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+Galileo (2025) [152] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
 SatDiFuser (2025)[78] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$

### v4

@@
 Abstract
-Satellite remote sensing from repeated observations and multiple sensors enables a wide range of downstream applications, including habitat mapping, carbon accounting, and strategies for conservation and sustainable land use.
-However, satellite time series are voluminous, often corrupted by sensor noise, clouds, and atmospheric conditions, and unevenly spaced in time, making them challenging to use.
+Satellite remote sensing enables a wide range of downstream applications, including habitat mapping, carbon accounting, and strategies for conservation and sustainable land use.
+However, satellite time series are voluminous and often corrupted, making them challenging to use.
 We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time-series data.
-TESSERA uses two parallel Transformer-based encoders to combine optical data from ten Sentinel-2 spectral bands at 10-20 m spatial resolution and two Sentinel-1 synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are subsequently fused with a multilayer perceptron to create annual global embedding maps.
+TESSERA uses two encoders to combine optical data with synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are fused with a multilayer perceptron to create annual global embedding maps.
 We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines.
 We believe that TESSERA’s ease of use, state-of-the-art performance, openness, and computation- and labelled data-efficiency will prove transformative in a wide range of ecological applications.
-Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [117, 185, 25].
-Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise continued growth in the future [178].
-However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labelled datasets required for supervised training [95].
-The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling, preventing full and easy use of this valuable resource [196, 61, 131].
-EO data users usually mitigate these issues by annual or seasonal compositing [35], then using supervised machine learning to map from composited images to task-specific classification, such as crop identification [9].
+Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation  (?, ?, ?).
+Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise continued growth in the future  (?).
+However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labelled datasets required for supervised training  (?).
+The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling, preventing full and easy use of this valuable resource  (?, ?, ?).
+EO data users usually mitigate these issues by annual or seasonal compositing  (?), then using supervised machine learning to map from composited images to task-specific classification, such as crop identification  (?).
 However, compositing removes the critical temporal signal arising from phenology (i.e.
 timing of biological events such as flowering) and training a task-specific ML model is computationally onerous.
-Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data [180].
-These include both supervised pre-training on large labelled datasets for spatio-temporal analysis [22, 21, 162, 19], and self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
-Common SSL techniques include generative methods, such as masked autoencoders [33, 136, 68, 151], diffusion-based models  [195, 81, 114], and contrastive methods [15, 60, 71, 171].
-General purpose foundation models must be adapted (or ‘fine-tuned’) for specific downstream tasks such as habitat mapping and semantic segmentation [163, 179, 186, 29, 193].
-Although existing SSL approaches create rich multidimensional ‘embeddings’ from unlabelled data through auxiliary tasks that reveal latent patterns [16, 97], they still do not address the loss of the temporal signal during compositing.
+Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data  (?).
+These include both supervised pre-training on large labelled datasets for spatio-temporal analysis  (?, ?, ?, ?), and self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
+Common SSL techniques include generative methods, such as masked autoencoders  (?, ?, ?, ?), diffusion-based models   (?, ?, ?), and contrastive methods  (?, ?, ?, ?).
+General purpose foundation models must be adapted (or ‘fine-tuned’) for specific downstream tasks such as habitat mapping and semantic segmentation  (?, ?, ?, ?, ?).
+Although several existing SSL approaches create rich multidimensional ‘embeddings’ from unlabelled data through auxiliary tasks that reveal latent patterns  (?, ?), they still do not address the loss of the temporal signal during compositing.
 In recent work, Lisaius et al.
-[95] have demonstrated that the Barlow Twins loss [189], grounded in redundancy reduction principles [18], can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series.
+(?) have demonstrated that the Barlow Twins loss  (?), grounded in redundancy reduction principles  (?), can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series.
 Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks.
@@
 This makes TESSERA frugal in terms of both data handling and computation.
-They can also keep their data private instead of having to upload it to a centralized repository.
+Users can also keep their data private instead of having to upload it to a centralized repository.
 This removes a crucial barrier to use.
@@
 Moreover, our model is open source, enabling extensibility and scientific reproducibility.
-We have made a gap-free, global, annual, 10 m resolution embedding map for 2024 openly available through the Geotessera Python library [154] and aim to eventually provide global maps for the entire period 2017-2024.
+We have made a gap-free, global, annual, 10 m resolution embedding map for 2024 openly available through the Geotessera Python library  (?) and aim to eventually provide global maps for the entire period 2017-2024.
 These information-rich precomputed maps greatly simplify a wide range of downstream EO applications.
-We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see [1] for details).
-To our knowledge, TESSERA is unprecedented in its ease of use, scale, and accuracy: no other foundation model provides analysis-ready outputs, is open, and provides global, annual coverage at 10 m resolution using only spectral-temporal features at pixel level (Supplementary Text:  LABEL:tab:sup_comparison).
-1 Overview of the TESSERA System
+We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see  (?) for details).
+To our knowledge, TESSERA is unprecedented in its ease of use, scale, and accuracy: no other foundation model provides analysis-ready outputs, is open, and provides global, annual coverage at 10 m resolution using only spectral-temporal features at pixel level (Supplementary Text table  LABEL:tab:sup_comparison).
+The work closest to ours is Google’s AlphaEarth foundation model  (?), which also provides global, annual, 10 m resolution embedding maps for 2017-2024.
+However, this work is neither open nor reproducible.
+Moreover, we demonstrate below that TESSERA embeddings match or outperform AlphaEarth embeddings in several downstream tasks.
+Overview of the TESSERA System
 TESSERA uses self-supervised learning to create per-pixel embeddings from a global dataset of unlabelled time-series of observations from Sentinel-1 Synthetic Aperture Radar (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams.
-It is trained by optimizing a modified Barlow Twins loss function [189].
+It is trained by optimizing a modified Barlow Twins loss function  (?).
 The original Barlow Twins approach seeks to create embeddings that are a) invariant to augmentations of its input, where these augmentations are typically rotations, reflections, and cropping of an image and b) whose dimensions are as decorrelated as possible.
 Inspired by the work by Lisaius et al.
-[95], we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions.
+(?), we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions.
 This is achieved by randomly sampling non-cloudy days from the Sentinel-2 annual observation time series so that model then learns to reconcile these partial views, interpolating missing data.
-The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Appendix A.7.2 for details, Equation 11).
+The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Supplementary Text Equation S11).
 We now discuss the system architecture.
@@
 Rows that are cloud-corrupted or otherwise invalid are indicated by a binary mask and ignored when creating augmentations.
-Model training and inferencing is done by random temporal sampling followed by flattening of these d-pixels (see Appendix A.4 and Fig.
-1 for details).
+Model training and inferencing is done by random temporal sampling followed by flattening of these d-pixels (see Fig.
+1 and Supplementary Text for details).
 In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
@@
 Under typical mid-latitude observation conditions, this compresses the input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
-Figure 1: Architecture and data processing pipeline of the TESSERA foundation model.
-a, Overview of the end-to-end workflow.
-Spatially aligned time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are first extracted for each pixel location, to form mode-specific ‘d-pixels’ (shown as vertical rectangle, with cloud-corrupted rows indicated in gray).
-To allow self-supervised training, two augmentations are created from each d-pixel.
-Each augmentation is a cloud-free random sample from the corresponding d-pixel.
-Augmentations are processed through a dual-branch architecture with weights shared between the branches, which is the key feature of Barlow Twin SSL.
-The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings.
-During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network.
-A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of these projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked (i.e, cloud-corrupted) days in the Sentinel-2 d-pixel.
-k b, Detailed structure of a single encoder branch.
-A d-pixel, represented as a matrix of temporal steps versus spectral bands (40 time-steps $\times$ 2 bands for Sentinel-1), is first embedded using a single-layer MLP (not shown).
-A Day-of-Year-based temporal encoding is added to this initial embedding.
-The resulting sequence is processed by an 4-block Transformer Encoder.
-Finally, a Gated Recurrent Unit (GRU) pooling layer processes the sequence of temporal features; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
 TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared).
@@
 Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context.
-An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [1].
-See Appendix A.7 and Figure 1 for architectural details.
+An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality  (?).
+See Figure 1 and Supplementary Text for architectural details.
 Modality-specific embeddings derived from dual encoders are fused using a multilayer perceptron (MLP) which is trained alongside the encoders.
@@
 This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate redundancy reduction during the self-supervised learning phase.
-A detailed breakdown of the model’s architectural parameters is provided in Appendix A.6 (Table 2).
-We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [1].
-1.1 Model Training
-The pretraining phase utilizes a large global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series.
+A detailed breakdown of the model’s architectural parameters is provided in the Supplementary table S2.
+We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance   (?).
+Model Training
+The pre-training phase utilizes a large global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series.
 To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed.
 This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features.
-The model was trained for one epoch in this dataset using distributed training techniques (see Appendix A.7.3 for further details on data preparation, the shuffling strategy, and optimization).
-The key hyperparameters used during pretraining are summarized in Appendix A.7.3 (Table 3).
-1.2 Model Inference and Global Embeddings Maps
-Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally.
+The model was trained for one epoch in this dataset using distributed training techniques (see  (?) for further details on data preparation, the shuffling strategy, and optimization).
+The key hyperparameters used during pre-training are summarized in Supplementary table S3.
+Model Inference and Global Embeddings Maps
+Following pre-training, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally.
 This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders.
@@
 These maps are designed as readily usable multichannel geospatial data layers.
-This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Appendix A.8).
-We are generating global embeddings for 2024 (approx.
+This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Supplementary Text ).
+We are currently generating global embeddings for 2024 (approx.
 250 TB) and will progressively extend coverage backward year by year until 2017.
@@
 2).
-N
-Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales.
-The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization.
-a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns.
-b, A multi-scale view of a complex agricultural landscape.
-b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields.
-c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
-2 Downstream Tasks
+Downstream Tasks
 The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks.
 These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation.
-(See Appendix D and Supplementary Fig.
-12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings.
+(See Supplementary fig.
+S5 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings.
 Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling.
 This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
-2.1 Crop Type Classification
-Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [177].
-The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [61].
-However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [61].
-In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [61, 113].
-Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [7, 80].
+Crop Type Classification
+Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector  (?).
+The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions  (?).
+However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes  (?).
+In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard  (?, ?).
+Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally  (?, ?).
 Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
-To evaluate performance, we used the INVEKOS Austrian crop dataset [3] from the 2021–2022 growing season.
-This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3].
-We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [151], and (3) Google Satellite Embedding (GSE) [57].
+To evaluate performance, we used the INVEKOS Austrian crop dataset  (?) from the 2021–2022 growing season.
+This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets)  (?).
+We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model  (?), and (3) AlphaEarth Google Satellite Embedding (GSE)  (?).
 For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison.
@@
 We also evaluated performance on a patch-based semantic segmentation task against several other foundation models.
-Details on data processing, model architectures, and training protocols are provided in Appendix E.
+Details on data processing, model architectures, and training protocols are provided in the Supplementary Text.
 TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labelled data are scarce (Fig.
@@
 These performance gains are due to the high quality of the learned embeddings.
-A UMAP analysis [102] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig.
+A UMAP analysis  (?) of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig.
 3d).
 This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
+Canopy Height Estimation
+Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+  (?, ?).
+It is correlated with ecosystem function  (?), forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness  (?, ?, ?).
+Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
+Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability.
+Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range  (?).
+GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes  (?), and does not produce unreliable estimates in mountainous terrain  (?).
+Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height  (?).
+Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors   (?, ?) - yet none of the three most widely used global CHM models currently implement sensor fusion  (?, ?, ?).
+Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500 features in Potapov et al.
+(?)), which may generalize poorly across ecosystems.
+These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation  (?), and inconsistencies in input quality.
+As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes  (?).
+Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
+These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling  (?, ?).
+We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo  (?) with openly available ground-truth data  (?).
+To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation.
+A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input.
+To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set.
+The accuracy metrics for the foundation models are reported as the average across the 12 runs.
+TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62 m.
+We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig.
+4).
+First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.
+(?), the ETH map by Lang et al.
+(?), and the Meta map by Tolan et al.
+(?).
+Although local regression models, such as ours, generally outperform global products in regional evaluations  (?), the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages  (?, ?, ?)—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
+We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
+(?), which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass.
+Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO  (?) and Google Satellite Embedding (GSE)  (?), using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
+Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
+In all comparisons, TESSERA outperformed competing methods.
+Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility.
+Instead, only the region and year of interest are needed to replicate the input data pipeline.
+By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
+Above-ground Forest Biomass Estimation
+Forests play a central role in the global carbon cycle  (?).
+This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery  (?).
+A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions.
+Thus, there has been a long-standing effort to link field measurements to global satellite observations  (?).
+Recently, with the creation of the BioMassters benchmark dataset  (?) (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice   (?).
+We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98.
+In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT  (?) and GSE  (?).
+First, a summary of the BioMassters dataset.
+This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations  (?, ?) informed by extensive field plots.
+Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel.
+The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig.
+5b) that would not appear in coarser resolution data.
+To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
+We compared a UNet-based regression model  (?) trained on TESSERA embeddings with SpectralGPT embeddings  (?) combined with a lightweight spatial-temporal encoder  (?) for temporal aggregation and a UPerNet  (?) regression head, because this combination has been found in the Pangaea benchmark  (?) to outperform other foundation models  (?, ?).
+Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase.
+Thus, we adopted the Pangaea implementation of their model  (?), which uses preprocessed satellite imagery from the BioMassters dataset.
+We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline.
+This model uses a UNet with a temporal attention encoder  (?).
+Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input.
+However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
+Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
+Our performance metrics are the mean RMSE, the coefficient of determination ($R^{2}$), and the mean bias (MB) in AGB across the test patches as a function of the fraction of labelled data used to train the regressor (Fig.
+5a).
+We make three significant findings.
+First, although custom models developed using supervised learning are typically difficult to outperform, even with a small fraction of training labels, our regressor nearly matches the performance of the BioMassters competition winner, which uses a supervised learning approach.
+Second, our regressor trained on 1% of the labelled data outperforms a UPerNet-based regressor trained on the entire training set using SpectralGPT embeddings.
+Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
+Burned Area Detection
+The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change  (?, ?, ?).
+Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends  (?), predicting wildfire occurrence  (?), assessing ecosystem recovery  (?), and establishing effective fire management strategies  (?).
+Remote sensing is widely used to detect burned areas and assess severity regionally and globally  (?).
+Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input  (?, ?).
+These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals  (?).
+Most existing methods use optical sensors  (?), which can detect fire-specific spectral signatures but are susceptible to cloud cover  (?), particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens  (?).
+SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover  (?, ?, ?).
+Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness  (?, ?).
+Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products  (?, ?, ?).
+TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig.
+6).
+We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6).
+Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset  (?).
+UMAP projections  (?) of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters.
+In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a).
+To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels.
+We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites.
+The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b).
+Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency.
+These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
+To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps.
+These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Figure 6C).
+The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
+Stocking Indices in Voluntary Carbon Markets
+The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere  (?).
+To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
+This is a costly and inefficient undertaking if it is done only by field sampling.
+To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra  (?)) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB.
+Upon acceptance, project developers can use the index to assess the performance of the project.
+Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
+To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.
+(?).
+Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see   (?) for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil.
+This rescaling ensures a fair comparison between the various products.
+A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Figure 7).
+As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
+TESSERA: a Global Open-Source Foundation Model
+In this work, we introduce TESSERA, an innovative remote sensing foundation model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
+TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing.
+We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models.
+Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
+Figure 1: Data processing pipeline.
+a, Overview.
+Time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are extracted for each pixel location to form mode-specific ‘d-pixels’ ( vertical rectangles; cloud-corrupted rows in gray).
+Two augmentations are created from each d-pixel as cloud-free random samples from the corresponding d-pixel.
+Augmentations are processed through a dual-branch architecture with weights shared between the branches.
+The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings.
+During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network.
+A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked days in the Sentinel-2 d-pixel.
+b, Detailed structure of a single encoder branch.
+Each d-pixel is embedded using a single-layer MLP (not shown) and a Day-of-Year-based temporal encoding is added.
+The resulting vector is processed by an 4-block Transformer Encoder and then a Gated Recurrent Unit (GRU) pooling layer; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
+Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales.
+The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization.
+a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns.
+b, A multi-scale view of a complex agricultural landscape.
+b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields.
+c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
 Figure 3: TESSERA embeddings achieve state-of-the-art performance in crop type classification.
@@
 TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
-2.2 Canopy height estimation
-Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [105, 50].
-It is correlated with ecosystem function [104], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [38, 26, 138].
-Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
-Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability.
-Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [23].
-GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [42], and does not produce unreliable estimates in mountainous terrain [53].
-Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [132].
-Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [160, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 120, 150].
-Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al.
-[120]), which may generalize poorly across ecosystems.
-These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality.
-As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [106].
-Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
-These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [107, 115].
-We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [75] with openly available ground-truth data [34].
-To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation.
-A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input.
-To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set.
-The accuracy metrics for the foundation models are reported as the average across the 12 runs.
-TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
-We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig.
-4).
-First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.
-[120], the ETH map by Lang et al.
-[84], and the Meta map by Tolan et al.
-[150].
-Although local regression models, such as ours, generally outperform global products in regional evaluations [128], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [120, 84, 150]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
-We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
-[85], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass.
-Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [151] and Google Satellite Embedding (GSE) [57], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
-Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
-In all comparisons, TESSERA outperformed competing methods.
-Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility.
-Instead, only the region and year of interest are needed to replicate the input data pipeline.
-By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
 Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation.
@@
 Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure.
-TESSERA achieves the highest coefficient of determination and the lowest RMSE (8.88±0.988.88$\pm$ 0.988.88 ± 0.98 m), indicating superior performance in structurally complex, tall tropical canopies.
+TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies.
 Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM.
@@
 A 1:1 reference line (blue) is shown in each panel for comparison.
-2.3 Above-ground forest biomass estimation
-Forests play a central role in the global carbon cycle [67].
-This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [46].
-A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions.
-Thus, there has been a long-standing effort to link field measurements to global satellite observations [194].
-Recently, with the creation of the BioMassters benchmark dataset [110] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [5].
-We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98.
-In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [68] and GSE [57].
-First, a summary of the BioMassters dataset.
-This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [126, 127] informed by extensive field plots.
-Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel.
-The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig.
-5b) that would not appear in coarser resolution data.
-To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
-We compared a UNet-based regression model [129] trained on TESSERA embeddings with SpectralGPT embeddings [68] combined with a lightweight spatial-temporal encoder [133] for temporal aggregation and a UPerNet [181] regression head, because this combination has been found in the Pangaea benchmark [100] to outperform other foundation models [110, 100].
-Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase.
-Thus, we adopted the Pangaea implementation of their model [100], which uses preprocessed satellite imagery from the BioMassters dataset.
-We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline.
-This model uses a UNet with a temporal attention encoder [110].
-Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input.
-However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
-Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
-Our performance metrics are the mean RMSE, the coefficient of determination ($R^{2}$), and the mean bias (MB) in AGB across the test patches as a function of the fraction of labelled data used to train the regressor (Fig.
-5a).
-We make three significant findings.
-First, although custom models developed using supervised learning are typically difficult to outperform, even with a small fraction of training labels, our regressor nearly matches the performance of the BioMassters competition winner, which uses a supervised learning approach.
-Second, our regressor trained on 1% of the labelled data outperforms a UPerNet-based regressor trained on the entire training set using SpectralGPT embeddings.
-Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
 Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings.
@@
 The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
-2.4 Burned Area Detection
-The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [6, 72, 116].
-Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [11], predicting wildfire occurrence [30], assessing ecosystem recovery [188], and establishing effective fire management strategies [64].
-Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82].
-Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 118].
-These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82].
-Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [65], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [66].
-SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 145, 156].
-Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [82, 31].
-Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [145, 156, 130].
-TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig.
-6).
 Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California.
-a, UMAP projection of 2020 TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events.
+a, UMAP projection of $20$ TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events.
 A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance.
@@
 As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
-We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6).
-Our analysis focused on two randomly selected 15$\times$15 km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [4].
-UMAP projections [102] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters.
-In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a).
-To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels.
-We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites.
-The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b).
-Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency.
-These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
-To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps.
-These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Figure 6C).
-The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
-2.5 Stocking indices in voluntary carbon markets
-The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [109].
-To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
-This is a costly and inefficient undertaking if it is done only by field sampling.
-To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [157]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB.
-Upon acceptance, project developers can use the index to assess the performance of the project.
-Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
-To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.
-[14].
-Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [1] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil.
-This rescaling ensures a fair comparison between the various products.
-A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Figure 7).
-As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
-Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market.
-Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details).
-Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 X + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA AGB converted using 0.407 X + 34.0; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X AGB prediction converted using 0.3 X + 54.3.
-Details on the conversion functions are provided in Supplemental section I.
-3 Conclusion
-In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
-TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing.
-We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models.
-Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
+Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market (data has been linearly rescaled).
+Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products.
+Note that the different products were originally provided in different units and converted to AGB (see   (?) for details).
 Acknowledgments
@@
 SJ was funded by a charitable donation from John Bernstein.
+JB received support from the UKRI CRCRM scheme (grant MR/Y024354).
 SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
+DAC and SK were funded by the NERC Centre for Landscape Regeneration ( NE/W00495X/1 )
 Author contributions:
@@
 Data and materials availability:
-Our framework integrates openly available data from the Copernicus Sentinel-2 (S2) [48] optical and Sentinel-1 (S1) [47] SAR missions, leveraging the strengths of each sensor.
+Our framework integrates openly available data from the Copernicus Sentinel-2 (S2)  (?) optical and Sentinel-1 (S1)  (?) SAR missions, leveraging the strengths of each sensor.
 We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22).
 To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations.
-For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [49].
-The Google Satellite Embedding V1 dataset [57] is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
-PRESTO embeddings can be created on demand in Google Earth Engine [58] using VertexAI [59].
-Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code [154].
-To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository [123], which is the official third-party software repository for Python.
-This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer [63] that can be passed directly into downstream tasks.
-The embeddings themselves are checksummed and indexed using Pooch [153] to check data integrity and handle retrieval.
-The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [3].
-The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset [4].
-The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [34].
-The AGB downstream task uses the open BioMassters dataset [110].
-The code for the winning model of the BioMassters competition is described in [110] and freely available on the corresponding GitHub page.
-The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark [100].
+For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox  (?).
+The Google Satellite Embedding V1 dataset  (?) is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
+PRESTO embeddings can be created on demand in Google Earth Engine  (?) using VertexAI  (?).
+Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code  (?).
+To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository  (?), which is the official third-party software repository for Python.
+This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer  (?) that can be passed directly into downstream tasks.
+The embeddings themselves are checksummed and indexed using Pooch  (?) to check data integrity and handle retrieval.
+The crop type identification downstream task uses the open INVEKOS Austrian crop dataset  (?).
+The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset  (?).
+The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok  (?).
+The AGB downstream task uses the open BioMassters dataset  (?).
+The code for the winning model of the BioMassters competition is described in  (?) and freely available on the corresponding GitHub page.
+The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark  (?).
 The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners; hence, data access will require a Data Transfer Agreement.
-Appendix A Model details
+Supplementary materials
+Materials and Methods
+Figures S1 to S7
+Tables S1- S5
+References (97-0)
+Supplementary Materials for
+TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis
+[…]
+Equal contribution. Corresponding author. Email: sk818@cam.ac.uk.
+This PDF file includes:
+Materials and Methods
+Figures S1 to S7
+Tables S1- S5
+Materials and Methods
+Model details
 Our input data consists of Sentinel-2 optical satellite imagery (Level-2A bottom-of-atmosphere) and Sentinel-1 Synthetic Aperture Radar (SAR, Radiometrically Terrain Corrected).
-A.1 Data Representation
-We consider remote sensing data [47, 48] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
-$R_{t}\in\mathbb{R}^{W\times H\times C}$ (1)
+Data Representation
+We consider remote sensing data  (?, ?) with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
+$R_{t}\in\mathbb{R}^{W\times H\times C}$ (S1)
 where $W$ is the width (longitude dimension), $H$ is the height (latitude dimension), and $C$ is the number of spectral channels.
 Each tile is accompanied by a corresponding binary mask $V_{t}$ of dimensions:
-$V_{t}\in\{0,1\}^{W\times H}$ (2)
-[…]
-A.2 Temporal Data Stacking
+$V_{t}\in\{0,1\}^{W\times H}$ (S2)
+where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
+Temporal Data Stacking
 We stack spatially aligned tiles over a time period that spans $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
-$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (3)
-$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (4)
-A.3 Time Series Extraction and d-pixel Definition
-[…]
-$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (5)
-[…]
-$P_{i,j}(c)=S(i,j,c)$ (6)
+$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (S3)
+$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (S4)
+Time Series Extraction and d-pixel Definition
+For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
+$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (S5)
+We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
+$P_{i,j}(c)=S(i,j,c)$ (S6)
 In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time.
 Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there are valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
-Table 1: Variable definitions for remote sensing data structure
-Variable Dimensions Description
-$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
-$V_{t}$ $W\times H$ Binary mask for tile at time $t$ (1=valid, 0=invalid/clouded)
-$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
-$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
-[…]
-[…]
-[…]
-$W,H$ Scalar Spatial dimensions (width, height)
-$C$ Scalar Number of spectral channels + polarizations
-$T$ Scalar Number of time steps
-A.4 The d-pixel Representation
+The d-pixel Representation
 We represent each 1- m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
@@
 Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored in the temporal sampling process.
-A.5 Dual-Encoder Architecture
+Dual-Encoder Architecture
 Given the distinct nature of Sentinel-1 SAR and Sentinel-2 MSI data, TESSERA employs two separate, parallel transformer-based encoder branches.
@@
 To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings.
-The core of each encoder consists of a stack of 4 standard Transformer blocks [155], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
+The core of each encoder consists of a stack of 4 standard Transformer blocks  (?), featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
 To derive a single vector summarizing the entire time series for each modality, an attention-pooling layer weighs the importance of different time steps before aggregation.
 The resulting modality-specific embeddings (one from the S1 encoder, one from the S2 encoder) are then fused using a multi-layer perceptron.
-A.6 Projector Network
+Projector Network
 The fused embedding from the dual-encoder stage is subsequently fed into a large projector network.
@@
 Each of the first five layers is a fully-connected linear layer followed by Batch Normalization and a ReLU non-linear activation function, making the network deep and highly non-linear.
-This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [189], to enable effective redundancy reduction during the self-supervised loss computation.
+This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work  (?), to enable effective redundancy reduction during the self-supervised loss computation.
 The final output of the projector is then used in the loss calculation.
@@
 The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
-Table 2: Architectural parameters of the TESSERA v1.0 model.
-Component Parameters Percentage of Total
-Encoder (Feature Extractor) 45,697,026 3.29%
-   Sentinel-1 Encoder 22,846,465 1.64%
-   Sentinel-2 Encoder 22,850,561 1.64%
-Fusion MLP 131,200 $<$0.01%
-Projector (for Training only) 1,344,700,421 96.71%
-Total Parameters 1,390,528,647 100.00%
-A.7 Self-Supervised Training
-A.7.1 Augmented View Generation
-The TESSERA model is trained using a modified Barlow Twins objective function [189].
+Self-Supervised Training
+The TESSERA model is trained using a modified Barlow Twins objective function  (?).
 For this objective, two distorted views, denoted as $Y_{A}$ and $Y_{B}$, are generated for each input d-pixel.
@@
 Thus, our augmentations are fundamentally about sampling from the available, inherently incomplete information streams, rather than artificially distorting a complete input.
-A.7.2 Loss Function
 The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
-The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as [189]:
-$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (7)
+The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as  (?):
+$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (S7)
 Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
@@
 The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
-We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [189].
+We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference  (?).
 To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al.
-[17].
+(?).
 For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$.
@@
 The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
-$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (8)
-$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (9)
-$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (10)
-where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model’s outputs. The total loss function optimized during training is a weighted sum:
-$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
-where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [17]) to be effective based on our experiments.
-A.7.3 Pre-training Details
+$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (S8)
+$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (S9)
+$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (S10)
+where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model output. The total loss function optimized during training is a weighted sum:
+$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (S11)
+where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by  (?)) to be effective based on our experiments.
 The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles.
@@
 The temporal context of each sampled observation was encoded by transforming its normalized Day-of-Year (DOY) into sine and cosine features, which were then concatenated with the corresponding spectral or backscatter measurements.
-Table 3: Pre-training hyperparameters for the TESSERA v1.0 model.
+As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each).
+We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled.
+AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$.
+The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training.
+The global batch size was 32768.
+Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig.
+S2.
+A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
+Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch.
+To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data.
+This pipeline is conceptually illustrated in Fig.
+S3a.
+The impact of this approach on training stability is empirically demonstrated in Fig.
+S3b.
+Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
+This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
+Operationally, the process begins with the aggregation of d-pixels from all MGRS tiles into a single, comprehensive pool.
+A global shuffling operation is performed on this pool, a critical step to break the spatial contiguity of data from individual tiles and ensure each training batch contains a diverse mix of geographical and environmental contexts.
+Following this global shuffle, the data augmentation required by the Barlow Twins framework is applied.
+As detailed in the Supplementary Text, this involves generating two distinct augmented views (e.g., $Y^{A}$ and $Y^{B}$) for each d-pixel.
+To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed a custom high-performance pipeline that handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation.
+The resulting pairs of augmented d-pixels are then serialized into a compact, pickle-like file format.
+These files are organized into manageable chunks and loaded by PyTorch DataLoader workers, which stream the data and assemble the final training batches.
+This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
+Global Embedding Map Generation
+A primary output of the TESSERA project is the generation of annual global embedding maps with 1- meter resolution for the years 2017-2024.
+To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used.
+For each 1- meter pixel on the globe and for each year:
+  1. 1.
+    The full Sentinel-1 and Sentinel-2 time series data at 1- meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
+  2. 2.
+    A fixed number of 40 timesteps is sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
+  3. 3.
+    These sampled time series are fed into their respective frozen TESSERA encoders.
+  4. 4.
+    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional embedding vector for that pixel for that year.
+This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 1- meter grid.
+Scaling with data and network size
+To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task.
+Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
+S4).
+Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
+We found that using 4 transformer layers, each with 4 attention heads, provided a strong baseline without incurring excessive computational overhead during inference.
+This makes the feature-extracting part of the model computationally efficient.
+In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning.
+As detailed in Section Projector Network, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers.
+Our parameter sweep focused on this computationally intensive hidden block.
+We varied the number of these hidden layers (referred to as ’depth’ in Fig.
+S4a) and their width (number of neurons per layer), with results shown in Fig.
+S4a.
+As discussed by Zbontar et al.
+(?), deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task.
+The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide).
+However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure.
+For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
+Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
+S4b.
+The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
+This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks.
+The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
+Ablation study
+To validate the contribution of each key component within the TESSERA framework, we performed an ablation study.
+We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
+Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score  (?).
+The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions.
+Given a batch of embeddings represented by the matrix $Z\in\mathbb{R}^{N\times D}$, where $N$ is the batch size and $D$ is the embedding dimension, we first compute its singular values $s_{1},s_{2},\ldots,s_{D}$.
+These are normalized to form a probability distribution $p_{i}=s_{i}/\sum_{j}s_{j}$.
+The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
+$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (S12)
+A score closer to 1 indicates that the embeddings are of higher effective rank, utilizing the full dimensionality of the feature space more effectively to create richer, less redundant representations.
+The results of the ablation study are summarized in Table S4.
+Our analysis reveals several key insights into the model’s architecture and training strategy.
+The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches.
+Removing mixup regularization also severely degrades results.
+When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
+With removal of Sentinel-1 data, task performance decreases moderately, demonstrating some gain from the fusion of optical and SAR data yields.
+However, performance is still relatively good even with only Sentinel-2 data.
+This highlights both TESSERA’s overall robustness and its continued efficacy in regions or historical periods where Sentinel-1 data are sparse or unavailable, which is not uncommon in practice.
+Finally, the effect of quantization is particularly noteworthy.
+We employ Quantization-Aware Training (QAT) to convert the embeddings from 32-bit floating-point (float32) to 8-bit integer (int8) format.
+Removing this quantization step (i.e.
+using full float32 precision) results in a marginal performance improvement.
+This demonstrates the success of our QAT strategy: we achieve a nearly four-fold reduction in the storage and bandwidth requirements for our final data products, with a negligible impact on embedding quality.
+This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
+Downstream Task Application Methodology
+A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data.
+Having pre-trained TESSERA on large unlabelled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications.
+The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
+The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. S5):
+  1. 1.
+    Download Embeddings for the Region of Interest: The Geotessera Python library  (?) allows users to download embeddings for a desired region and year in the form of a numpy array.
+  2. 2.
+    Prepare labelled Downstream Data: The labelled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
+  3. 3.
+    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA embeddings as input.
+    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (1-3 layers) ending in a softmax output layer.
+    • For pixel-wise regression (e.g., canopy height regression), the head is usually an MLP ending in a single linear output neuron.
+    • For tasks requiring spatial context from the embeddings (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA embeddings (e.g., $64\times 64\times 128$).
+The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions.
+  4. 4.
+    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA embeddings as input features and the corresponding labels.
+Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used.
+This training typically requires significantly less labelled data and computational power compared to training a deep model from scratch.
+  5. 5.
+    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA embeddings for the test samples and passing them through the trained head.
+Performance is evaluated using standard metrics relevant to the task.
+This workflow allows the use of TESSERA embeddings in a range of diverse applications, demonstrating its role as a foundational model for geospatial analysis.
+TESSERA users who would like to generate their own embeddings can load the frozen weights from the saved pre-training checkpoint into the TESSERA dual encoder architecture.
+For each input sample (e.g., pixel, object, or patch) in the region of interest, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year must be downloaded and processed using the d-pixel creation pipeline.
+These d-pixels are passed through the frozen encoders and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA embedding for that sample.
+To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
+Downstream Task: Austrian Crop Classification
+Dataset and Preprocessing
+We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season  (?).
+The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation.
+For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
+Pixel-wise Classification Baselines
+To provide a comprehensive performance comparison, we implemented three distinct models for pixel-wise classification:
+  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA embeddings were used as input to a simple MLP.
+The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
+  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model  (?) to generate its pixel embeddings.
+These embeddings were then fed into an MLP head identical to the one used for TESSERA.
+  • GSE + MLP: To include another key foundation model, we used the publicly available Google Satellite Embedding (GSE) V1 dataset  (?).
+These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
+  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data.
+For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year.
+The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector.
+The RF model was configured with 100 trees, while all other hyperparameters retained their default values as implemented in the scikit-learn library.
+For the experiments shown in the main text’s figure, we trained these models on randomly selected subsets of the data (from 1% to 30% for panel a; specified samples-per-class for panel b), using a fixed validation set for hyperparameter tuning and a held-out test set for final evaluation.
+Patch-wise Semantic Segmentation
+To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
+  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO  (?), and GSE  (?), a distinct workflow was used.
+We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$).
+These patches were then fed into a standard UNet architecture  (?), which learns to model the spatial relationships between the embeddings to produce a segmentation map.
+It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do.
+However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
+  • For other foundation models that are inherently patch-based (e.g., Prithvi  (?), Satlas  (?)), their encoders already process image patches.
+Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output.
+We did not freeze the encoders during the fine-tuning.
+Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. S6.
+Embedding Space Analysis
+The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP)  (?) algorithm to the embeddings of every pixel within the Austrian study area.
+UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology.
+It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure.
+UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
+In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
+To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
+  • The Silhouette score, $s(i)$, for a single data point $i$ measures how similar it is to its own cluster compared to other clusters. It is defined as:
+    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (S13)
+    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighbouring cluster.
+The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
+  • The Davies-Bouldin Index (DBI) evaluates clustering quality by computing the ratio of within-cluster scatter to between-cluster separation. For a set of $k$ clusters, it is defined as:
+    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (S14)
+    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$.
+Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
+Downstream Task: Canopy height estimation
+Study Areas and Ground Truth Datasets
+[…]
+Modelling methods
+To predict the canopy height from the different embeddings, we used a standard U-Net architecture  (?).
+It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations.
+For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube.
+The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values.
+The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches.
+At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation.
+In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training.
+Model training was done in Python using PyTorch.
+To ensure spatial generalization and avoid overfitting local patterns, the reference data were divided into training, validation, and test sets that are spatially disjoint.
+Each study region was divided into four spatial folds based on cardinal directions (bottom, top, left, right), each fold covering 50% of the area.
+For each fold, the remaining half of the region was used for training and validation.
+To account for inherent model randomness, we trained three independent models per fold, resulting in 12 trained models for each foundation model.
+This strategy captures variability due to both spatial heterogeneity and training randomness.
+The final performance of the model was evaluated on the test data held using the following metrics: coefficient of determination (R²), root mean square error (RMSE), and mean bias.
+Baselines
+TESSERA, PRESTO  (?), and GSE  (?) embeddings were compared by training identical models using each representation as input, with all other settings held constant.
+Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
+To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al.
+[2021] (?) (GLAD), Lang et al.
+[2023] (?) (ETH), and Tolan et al.
+[2024] (?) (Meta) as well as a regional map produced by Lang et al.
+[2021] (?).
+These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
+Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
+Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
+This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery.
+Cloud masking and compositing were performed in GEE, resulting in 12 channel input images, with one channel for each Sentinel-2 spectral band.
+Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
+Downstream Task: Burned area detection
+Study Areas and Ground Truth Datasets
+For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset  (?), referred to as Burn Area A and Burn Area B.
+Each Burn Area contains over two million 1 m pixels.
+Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure.
+In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
+Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land.
+In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
+To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set  (?).
+The dataset from the Monitoring Trends in Burn Severity (MTBS) program  (?) was used to identify burned areas and severity of the burn.
+Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
+Burn severity, defined as the measure of ecological damage caused by fire  (?), was derived from categorical and continuous severity metrics provided by MTBS.
+To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire.
+MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
+Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality  (?).
+Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations.
+The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine  (?).
+For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high).
+The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
+Dimensionality Reduction
+We used UMAP (Uniform Manifold Approximation and Projection)  (?) to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires.
+We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections.
+To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs.
+unburned), land cover class, or burn severity.
+These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
+To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
+UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space.
+UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
+Label Efficiency Analysis and Baselines
+We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available.
+The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set.
+We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%.
+The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance.
+Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
+We compared three types of input embeddings for the classification task.
+The first was TESSERA, using embeddings for 2020.
+The second was a task-specific composite baseline, constructed using prior knowledge of the fire event.
+This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector.
+The third type of input was Google Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset  (?).
+GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
+Downstream Task: Above-ground biomass estimation
+Dataset
+For the AGB regression task, we used the ground truth data from the BioMassters dataset  (?).
+This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution.
+Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre.
+Patch coordinates and collection years–provided by the authors of Ref.
+(?) upon request–were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE)  (?).
+UNet model architecture
+To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture  (?), using TESSERA or GSE embeddings   (?) as input.
+The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
+The network followed a standard UNet structure with two encoder and decoder stages.
+The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling.
+A 1024-channel bottleneck block was used prior to upsampling.
+The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double-convolution blocks.
+A final 1×1 convolution produced the single-channel regression output.
+The model had approximately 30 million trainable parameters.
+The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
+Following the evaluation protocol in Ref.
+(?), RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch.
+For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20).
+The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
+Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
+Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
+Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
+Baselines
+In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition  (?) and the SpectralGPT foundation model  (?) combined with a UPerNet  (?).
+For SpectralGPT, we used the implementation provided in Ref.
+(?), with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE  (?) for temporal aggregation, and a UPerNet as the regression head.
+This setup was found in  (?) to outperform other tested foundation models on the BioMassters dataset.
+To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
+Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss  (?) because it resulted in slightly better performance.
+Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
+For evaluating the supervised baseline, we downloaded the pre-trained weights of the BioMassters winning model and ran inference on the test set, applying a post-processing step to clip predictions to a non-negative range.
+Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model.
+According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
+Downstream Task: Stocking indices in Voluntary Carbon Markets
+Relative height modelling
+Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events.
+The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
+To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings.
+This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
+To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product  (?) from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil.
+From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables.
+A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
+The TerraClass Amazonia 2022 land cover map  (?) was used to distinguish between forest and non-forest.
+To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
+Stocking index baselines
+For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
+Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
+For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m)  (?), CTrees canopy tree height (5 m)  (?), ESA-CCI AGB (100 m)  (?), Meta CHM (1 m)   (?), and GTDX AGB (25 m)  (?).
+Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
+To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
+After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and $R^{2}$) directly comparable.
+The rescaling is as follows:
+  1. 1.
+    TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 $\times$ + 160.5
+  2. 2.
+[…]
+  3. 3.
+    CTrees canopy height converted using 1.806 $\times$ + 44.9
+  4. 4.
+    ESA AGB converted using 0.407 $\times$ + 34.0
+  5. 5.
+    Meta canopy height converted using 3.723 $\times$ + 46.6
+  6. 6.
+    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
+Interactive Habitat Mapping Tool
+Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure S7.
+The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles.
+To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
+The core of the tool is the interactive labelling workflow.
+A user can define custom habitat classes and provide training data by clicking on the map to place labelled points.
+These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier.
+The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view.
+This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise.
+We also provide a batch processing tool for larger regions of interest.
+The tools and their open source code are freely available in our github repository  (?).
+Figure S1: TESSERA is trained on globally-distributed training data.
+TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024.
+The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend.
+This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
+Figure S2: Training dynamics curves show that TESSERA rapidly reaches its asymptotes.
+a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss.
+b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay.
+c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction.
+d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant.
+e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
+Figure S3: Data shuffling greatly improves training stability.
+a, Schematic of the data shuffling and loading process.
+D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool.
+A custom Rust binary  (?) performs a global shuffle on this multi-terabyte dataset before applying augmentations.
+The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches.
+b, Comparison of training loss curves.
+The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias.
+The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
+Figure S4: Model performance improves with increasing projector size and additional computation.
+a, Hyperparameter sweep for the projector network’s architecture.
+Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis).
+The colour and size of the point indicate the validation F1 score and the number of parameters, respectively.
+Performance generally increases with both width and depth.
+The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints.
+The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency.
+b, Scaling laws for TESSERA.
+The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters).
+A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used.
+The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
+Figure S5: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications.
+Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
+Users can then fine-tune the entire model or only the head.
+Another approach (b) is where both the encoder and the projector (or decoder) are typically updated.
+In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input.
+This results in lightweight downstream models that consist only of a task-specific head.
+Figure S6: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
+Each row corresponds to a different patch size (32x32, 64x64, 128x128).
+The first column shows the ground truth segmentation.
+For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
+Figure S7: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
+We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
+Table S1: Variable definitions for remote sensing data structure
+Variable Dimensions Description
+$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
+$V_{t}$ $W\times H$ Binary mask for tile at time $t$ (1=valid, 0=invalid/clouded)
+$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
+$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
+$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
+$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
+$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
+$W,H$ Scalar Spatial dimensions (width, height)
+$C$ Scalar Number of spectral channels + polarizations
+$T$ Scalar Number of time steps
+Table S2: Architectural parameters of the TESSERA v1.0 model.
+Component Parameters Percentage of Total
+Encoder (Feature Extractor) 45,697,026 3.29%
+   Sentinel-1 Encoder 22,846,465 1.64%
+   Sentinel-2 Encoder 22,850,561 1.64%
+Fusion MLP 131,200 <<0.01%
+Projector (for Training only) 1,344,700,421 96.71%
+Total Parameters 1,390,528,647 100.00%
+Table S3: Pre-training hyperparameters for the TESSERA v1.0 model.
 Description Symbol Value
@@
 Quantization Bits (QAT) - 8
-Figure 8: TESSERA is trained on globally-distributed training data.
-TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024.
-The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend.
-This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
-As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each).
-We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled.
-AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$.
-The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training.
-The global batch size was 32768.
-Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary Fig.
-9.
-Figure 9: Training dynamics curves show that TESSERA rapidly reaches its asymptotes.
-a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss.
-b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay.
-c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction.
-d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant.
-e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
-A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
-Given that d-pixels within an individual MGRS tile exhibit high spatial autocorrelation, a naive sequential or locally-shuffled data loading process would expose the model to strong geographic biases in each batch.
-To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data.
-This pipeline is conceptually illustrated in Fig.
-10a.
-The impact of this approach on training stability is empirically demonstrated in Fig.
-10b.
-Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
-This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
-Figure 10: Data shuffling greatly improves training stability.
-a, Schematic of the data shuffling and loading process.
-D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool.
-A custom Rust binary [154] performs a global shuffle on this multi-terabyte dataset before applying augmentations.
-The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches.
-b, Comparison of training loss curves.
-The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias.
-The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
-Operationally, the process begins with the aggregation of d-pixels from all MGRS tiles into a single, comprehensive pool.
-A global shuffling operation is performed on this pool, a critical step to break the spatial contiguity of data from individual tiles and ensure each training batch contains a diverse mix of geographical and environmental contexts.
-Following this global shuffle, the data augmentation required by the Barlow Twins framework is applied.
-As detailed in Appendix 5.1, this involves generating two distinct augmented views (e.g., $Y^{A}$ and $Y^{B}$) for each d-pixel.
-To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed a custom high-performance pipeline that handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation.
-The resulting pairs of augmented d-pixels are then serialized into a compact, pickle-like file format.
-These files are organized into manageable chunks and loaded by PyTorch DataLoader workers, which stream the data and assemble the final training batches.
-This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
-A.8 Global Embedding Map Generation
-A primary output of the TESSERA project is the generation of annual global embedding maps with 1- meter resolution for the years 2017-2024.
-To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used.
-For each 1- meter pixel on the globe and for each year:
-  1. 1.
-    The full Sentinel-1 and Sentinel-2 time series data at 1- meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
-  2. 2.
-    A fixed number of 40 timesteps is sampled from the valid observations within the year for both Sentinel-1 and Sentinel-2 data, along with their DOY positional encodings.
-  3. 3.
-    These sampled time series are fed into their respective frozen TESSERA encoders.
-  4. 4.
-    The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional embedding vector for that pixel for that year.
-This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 1- meter grid.
-Appendix B Scaling with data and network size
-To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task.
-Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
-11).
-Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
-We found that using 4 transformer layers, each with 4 attention heads, provided a strong baseline without incurring excessive computational overhead during inference.
-This makes the feature-extracting part of the model computationally efficient.
-In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning.
-As detailed in Section A.6, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers.
-Our parameter sweep focused on this computationally intensive hidden block.
-We varied the number of these hidden layers (referred to as ’depth’ in Fig.
-11a) and their width (number of neurons per layer), with results shown in Fig.
-11a.
-As discussed by Zbontar et al.
-[189], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task.
-The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide).
-However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure.
-For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
-Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
-11b.
-The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
-This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks.
-The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
-Figure 11: Model performance improves with increasing projector size and additional computation.
-a, Hyperparameter sweep for the projector network’s architecture.
-Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis).
-The colour and size of the point indicate the validation F1 score and the number of parameters, respectively.
-Performance generally increases with both width and depth.
-The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints.
-The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency.
-b, Scaling laws for TESSERA.
-The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters).
-A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used.
-The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
-Appendix C Ablation study
-To validate the contribution of each key component within the TESSERA framework, we performed an ablation study.
-We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
-Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score [55].
-The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions.
-Given a batch of embeddings represented by the matrix $Z\in\mathbb{R}^{N\times D}$, where $N$ is the batch size and $D$ is the embedding dimension, we first compute its singular values $s_{1},s_{2},\ldots,s_{D}$.
-These are normalized to form a probability distribution $p_{i}=s_{i}/\sum_{j}s_{j}$.
-The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
-$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (12)
-A score closer to 1 indicates that the embeddings are of higher effective rank, utilizing the full dimensionality of the feature space more effectively to create richer, less redundant representations.
-The results of the ablation study are summarized in Table 4.
-Our analysis reveals several key insights into the model’s architecture and training strategy.
-The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches.
-Removing mixup regularization also severely degrades results.
-When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
-With removal of Sentinel-1 data, task performance decreases moderately, demonstrating some gain from the fusion of optical and SAR data yields.
-However, performance is still relatively good even with only Sentinel-2 data.
-This highlights both TESSERA’s overall robustness and its continued efficacy in regions or historical periods where Sentinel-1 data are sparse or unavailable, which is not uncommon in practice.
-Finally, the effect of quantization is particularly noteworthy.
-We employ Quantization-Aware Training (QAT) to convert the embeddings from 32-bit floating-point (float32) to 8-bit integer (int8) format.
-Removing this quantization step (i.e.
-using full float32 precision) results in a marginal performance improvement.
-This demonstrates the success of our QAT strategy: we achieve a nearly four-fold reduction in the storage and bandwidth requirements for our final data products, with a negligible impact on embedding quality.
-This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
-Table 4: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance.
+Table S4: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance.
 We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset.
@@
 w/o Embedding Quantization 0.779 0.972
-Appendix D Downstream Task Application Methodology
-A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data.
-Having pre-trained TESSERA on large unlabelled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications.
-The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
-The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary Fig. 12):
-Figure 12: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications.
-Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
-Users can then fine-tune the entire model or only the head.
-Another approach (b) is where both the encoder and the projector (or decoder) are typically updated.
-In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input.
-This results in lightweight downstream models that consist only of a task-specific head.
-  1. 1.
-    Download Embeddings for the Region of Interest: The Geotessera Python library [154] allows users to download embeddings for a desired region and year in the form of a numpy array.
-  2. 2.
-    Prepare labelled Downstream Data: The labelled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
-  3. 3.
-    Design Task-Specific Head: A lightweight, task-specific neural network module (the ”head”) is designed. This head takes the extracted TESSERA embeddings as input.
-    • For pixel-wise classification (e.g., crop classification), the head is typically a shallow MLP (1-3 layers) ending in a softmax output layer.
-    • For pixel-wise regression (e.g., canopy height regression), the head is usually an MLP ending in a single linear output neuron.
-    • For tasks requiring spatial context from the embeddings (e.g., canopy height mapping over an area, semantic segmentation), the input to the head can be a patch of TESSERA embeddings (e.g., 64×64×12864$\times$ 64$\times$ 12864 × 64 × 128).
-The head might then be a convolutional architecture, such as a UNet, that processes these spatial feature maps to produce dense predictions.
-  4. 4.
-    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA embeddings as input features and the corresponding labels.
-Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used.
-This training typically requires significantly less labelled data and computational power compared to training a deep model from scratch.
-  5. 5.
-    Evaluation: Once the head is trained, inference is performed on a test set by extracting TESSERA embeddings for the test samples and passing them through the trained head.
-Performance is evaluated using standard metrics relevant to the task.
-This workflow allows the use of TESSERA embeddings in a range of diverse applications, demonstrating its role as a foundational model for geospatial analysis.
-TESSERA users who would like to generate their own embeddings can load the frozen weights from the saved pre-training checkpoint into the TESSERA dual encoder architecture.
-For each input sample (e.g., pixel, object, or patch) in the region of interest, its corresponding Sentinel-1 and Sentinel-2 time series data for the relevant year must be downloaded and processed using the d-pixel creation pipeline.
-These d-pixels are passed through the frozen encoders and their outputs are fused by the MLP to generate the final 128-dimensional TESSERA embedding for that sample.
-To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
-Appendix E Downstream Task: Austrian Crop Classification
-E.1 Dataset and Preprocessing
-We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [3].
-The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation.
-For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
-E.2 Pixel-wise Classification Baselines
-To provide a comprehensive performance comparison, we implemented three distinct models for pixel-wise classification:
-  • TESSERA + MLP: The primary model, where the frozen 128-dimensional TESSERA embeddings were used as input to a simple MLP.
-The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
-  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model [151] to generate its pixel embeddings.
-These embeddings were then fed into an MLP head identical to the one used for TESSERA.
-  • GSE + MLP: To include another key foundation model, we used the publicly available Google Satellite Embedding (GSE) V1 dataset [57].
-These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
-  • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data.
-For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year.
-The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector.
-The RF model was configured with 100 trees, while all other hyperparameters retained their default values as implemented in the scikit-learn library.
-For the experiments shown in the main text’s figure, we trained these models on randomly selected subsets of the data (from 1% to 30% for panel a; specified samples-per-class for panel b), using a fixed validation set for hyperparameter tuning and a held-out test set for final evaluation.
-E.3 Patch-wise Semantic Segmentation
-To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
-  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [151], and GSE [57], a distinct workflow was used.
-We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$).
-These patches were then fed into a standard UNet architecture [129], which learns to model the spatial relationships between the embeddings to produce a segmentation map.
-It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do.
-However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
-  • For other foundation models that are inherently patch-based (e.g., Prithvi [143], Satlas [19]), their encoders already process image patches.
-Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output.
-We did not freeze the encoders during the fine-tuning.
-Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 13.
-E.4 Embedding Space Analysis
-The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [102] algorithm to the embeddings of every pixel within the Austrian study area.
-UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology.
-It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure.
-UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
-In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
-To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
-  • The Silhouette score, $s(i)$, for a single data point $i$ measures how similar it is to its own cluster compared to other clusters. It is defined as:
-    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (13)
-    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighbouring cluster.
-The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
-  • The Davies-Bouldin Index (DBI) evaluates clustering quality by computing the ratio of within-cluster scatter to between-cluster separation. For a set of $k$ clusters, it is defined as:
-    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (14)
-    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$.
-Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
-Figure 13: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
-Each row corresponds to a different patch size (32x32, 64x64, 128x128).
-The first column shows the ground truth segmentation.
-For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
-Appendix F Downstream Task: Canopy height estimation
-F.1 Study Areas and Ground Truth Datasets
-Canopy height prediction was evaluated in a 5$\times$6 km area in the Danum Valley, Borneo.
-This region is covered in tropical rainforest dominated by species in the Dipterocarpaceae family, which hosts some of the tallest trees in the tropics [137, 119].
-Ground truth data was derived from airborne laser scanning (ALS) performed in February 2020 [34].
-ALS data provides detailed 3D point clouds of vegetation and terrain structure, which were processed using the  lspikefree algorithm [52] to estimate top-of-canopy height at 1 m resolution.
-These canopy height rasters were then downsampled using mean pooling to 10 m to match the resolution of the TESSERA representations.
-F.2 Modelling methods
-To predict the canopy height from the different embeddings, we used a standard U-Net architecture [129].
-It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations.
-For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube.
-The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values.
-The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches.
-At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation.
-In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training.
-Model training was done in Python using PyTorch.
-To ensure spatial generalization and avoid overfitting local patterns, the reference data were divided into training, validation, and test sets that are spatially disjoint.
-Each study region was divided into four spatial folds based on cardinal directions (bottom, top, left, right), each fold covering 50% of the area.
-For each fold, the remaining half of the region was used for training and validation.
-To account for inherent model randomness, we trained three independent models per fold, resulting in 12 trained models for each foundation model.
-This strategy captures variability due to both spatial heterogeneity and training randomness.
-The final performance of the model was evaluated on the test data held using the following metrics: coefficient of determination (R²), root mean square error (RMSE), and mean bias.
-F.3 Baselines
-TESSERA, PRESTO [151], and GSE [57] embeddings were compared by training identical models using each representation as input, with all other settings held constant.
-Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
-To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al.
-[2021][120] (GLAD), Lang et al.
-[2023][84] (ETH), and Tolan et al.
-[2024][150] (Meta) as well as a regional map produced by Lang et al.
-[2021][85].
-These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
-Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
-Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
-This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery.
-Cloud masking and compositing were performed in GEE, resulting in 12 channel input images, with one channel for each Sentinel-2 spectral band.
-Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
-Appendix G Downstream Task: Burned area detection
-G.1 Study Areas and Ground Truth Datasets
-For fire disturbance detection, two 15$\times$15 km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [77], referred to as Burn Area A and Burn Area B.
-Each Burn Area contains over two million 1 m pixels.
-Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure.
-In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
-Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land.
-In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
-To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [2].
-The dataset from the Monitoring Trends in Burn Severity (MTBS) program [4] was used to identify burned areas and severity of the burn.
-Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
-Burn severity, defined as the measure of ecological damage caused by fire [56], was derived from categorical and continuous severity metrics provided by MTBS.
-To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire.
-MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
-Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [56].
-Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations.
-The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [58].
-For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high).
-The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
-G.2 Dimensionality Reduction
-We used UMAP (Uniform Manifold Approximation and Projection) [102] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires.
-We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections.
-To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs.
-unburned), land cover class, or burn severity.
-These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
-To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
-UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space.
-UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
-G.3 Label Efficiency Analysis and Baselines
-We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available.
-The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set.
-We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%.
-The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance.
-Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
-We compared three types of input embeddings for the classification task.
-The first was TESSERA, using embeddings for 2020.
-The second was a task-specific composite baseline, constructed using prior knowledge of the fire event.
-This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector.
-The third type of input was Google Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [57].
-GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
-Appendix H Downstream Task: Above-ground biomass estimation
-H.1 Dataset
-For the AGB regression task, we used the ground truth data from the BioMassters dataset [110].
-This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a 2,560$\times$2,560  m area at 10 meter pixel resolution.
-Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre.
-Patch coordinates and collection years–provided by the authors of Ref.
-[110] upon request–were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE) [57].
-H.2 UNet model architecture
-To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [129], using TESSERA or GSE embeddings  [57] as input.
-The model received 256$\times$256$\times$128 and 256$\times$256$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
-The network followed a standard UNet structure with two encoder and decoder stages.
-The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling.
-A 1024-channel bottleneck block was used prior to upsampling.
-The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double-convolution blocks.
-A final 1×1 convolution produced the single-channel regression output.
-The model had approximately 30 million trainable parameters.
-The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
-Following the evaluation protocol in Ref.
-[110], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch.
-For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20).
-The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
-Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
-Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
-Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
-H.3 Baselines
-In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [110] and the SpectralGPT foundation model [68] combined with a UPerNet [181].
-For SpectralGPT, we used the implementation provided in Ref.
-[100], with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE [133] for temporal aggregation, and a UPerNet as the regression head.
-This setup was found in [100] to outperform other tested foundation models on the BioMassters dataset.
-To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
-Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [100] because it resulted in slightly better performance.
-Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
-For evaluating the supervised baseline, we downloaded the pre-trained weights of the BioMassters winning model and ran inference on the test set, applying a post-processing step to clip predictions to a non-negative range.
-Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model.
-According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
-Appendix I Downstream Task: Stocking indices in Voluntary Carbon Markets
-I.1 Relative height modelling
-Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events.
-The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
-To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings.
-This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
-To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [42] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil.
-From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables.
-A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
-The TerraClass Amazonia 2022 land cover map [148] was used to distinguish between forest and non-forest.
-To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
-I.2 Stocking index baselines
-For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
-Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
-For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [84], CTrees canopy tree height (5 m) [158], ESA-CCI AGB (100 m) [134], Meta CHM (1 m)  [150], and GTDX AGB (25 m) [43].
-Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
-To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
-After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and R2) directly comparable.
-Appendix J Interactive Habitat Mapping Tool
-Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure 14.
-The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles.
-To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
-The core of the tool is the interactive labelling workflow.
-A user can define custom habitat classes and provide training data by clicking on the map to place labelled points.
-These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier.
-The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view.
-This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise.
-We also provide a batch processing tool for larger regions of interest.
-The tools and their open source code are freely available in our github repository [154].
-Figure 14: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
-We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
-Appendix K Supplementary Table
-The table below compares Tessera along several dimensions with a comprehensive list of remote sensing foundation models [74].
-Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy.
+Table S5: TESSERA provides unprecedented ease of use, scale, and accuracy.
 Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m).
@@
 TESSERA (2025) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-GeoKR (2021) [87] $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-CVPRW (2021) [141] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
-GASSL (2021) [15] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
-SeCo (2021) [101] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-DINO-MM (2022) [169] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
-SatMAE (2022) [33] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
-RS-BYOL (2022) [76] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
-GeCo (2022) [88] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
-RingMo (2022) [142] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-RVSA (2022) [165] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
-RSP (2022) [162] $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
-MATTER (2022) [8] $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-CSPT (2022) [192] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-CVPRW (2022) [135] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-BFM (2023) [27] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-TOV (2023) [147] ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
-CMID (2023) [108] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-RingMo-Sense (2023) [187] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
-lal-SimCLR (2023) [121] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-CACo (2023) [99] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SatLas (2023) [19] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-GFM (2023) [103] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Scale-MAE (2023) [125] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-DINO-MC (2023) [176] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-CROMA (2023) [54] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Cross-Scale MAE (2023) [146] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-DeCUR (2023) [168] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
-Presto (2023) [151] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-CtxMIM (2023) [191] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-FG-MAE (2023) [172] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Prithvi (2023) [136] $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-RingMo-Lite (2023) [174] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-IGARSS (2023) [51] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-EarthPTi (2023) [139] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-USat (2023) [73] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-AIEarth (2023) [184] p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Clay (2023) [32] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Hydro (2023) [36] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-U-Barn (2023) [45] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
-GeRSP (2023) [70] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SwiMDiff (2023) [149] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-OFA-Net (2023) [183] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-SML-FR (2023) [41] p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-Spectral-GP (2024) [68] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-S2MAE (2024) [90] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SatMAE++ (2024) [112] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-msGFM (2024) [62] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SkySense (2024) [60] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-MTP (2024) [164] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-DOFA (2024) [182] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-MMEarth (2024) [111] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-LeMeViT (2024) [79] $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
-SoftCon (2024) [170] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-RS-DFM (2024) [175] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-A2MAE (2024) [190] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-HyperSIGM (2024) [161] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SelectiveMAE (2024) [166] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-OmniSat (2024) [12] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
-MM-VSE (2024) [124] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
-MA3E (2024) [94] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SpectralEarth (2024) [93] $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SenPa-MAE (2024) [122] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-RingMo-Ae (2024) [39] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SAR-JEPA (2024) [89] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
-PIS (2024) [10] $\times$ $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
-OReole-FM (2024) [40] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-PIEVIT (2024) [98] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SatVisionTOA (2024) [140] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-RS-vHeat (2024) [69] $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
-Prithvi-EO-2.0 (2024) [143] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-AnySat (2024) [13] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-WildSAT (2024) [37] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SeaMo (2024) [92] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-FoMo (2025) [24] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-SatMamba (2025) [44] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Galileo (2025) [152] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
-SatDiFuser (2025)[78] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-RoMA (2025) [167] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Panopticon (2025) [159] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-FedSense (2025) [144] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Copernicus-FM (2025) [173] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-HyperFree (2025) [86] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-DynamicVis (2025) [28] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-FlexiMo (2025) [91] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Google Satellite (2025)[57] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-RingMoE (2025) [20] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+GeoKR (2021)  (?) $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+CVPRW (2021)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
+GASSL (2021)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
+SeCo (2021)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+DINO-MM (2022)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
+SatMAE (2022)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
+RS-BYOL (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
+GeCo (2022)  (?) $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
+RingMo (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+RVSA (2022)  (?) $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
+RSP (2022)  (?) $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
+MATTER (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+CSPT (2022)  (?) $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+CVPRW (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+BFM (2023)  (?) $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+TOV (2023)  (?) ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
+CMID (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+RingMo-Sense (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
+lal-SimCLR (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+CACo (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SatLas (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+GFM (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Scale-MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+DINO-MC (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+CROMA (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Cross-Scale MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+DeCUR (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
+Presto (2023)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+CtxMIM (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+FG-MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Prithvi (2023)  (?) $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+RingMo-Lite (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+IGARSS (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+EarthPTi (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+USat (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+AIEarth (2023)  (?) p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Clay (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Hydro (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+U-Barn (2023)  (?) $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
+GeRSP (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SwiMDiff (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+OFA-Net (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+SML-FR (2023)  (?) p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Spectral-GP (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+S2MAE (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SatMAE++ (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+msGFM (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SkySense (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+MTP (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+DOFA (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+MMEarth (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+LeMeViT (2024)  (?) $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
+SoftCon (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+RS-DFM (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+A2MAE (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+HyperSIGM (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SelectiveMAE (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+OmniSat (2024)  (?) $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
+MM-VSE (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
+MA3E (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SpectralEarth (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SenPa-MAE (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+RingMo-Ae (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SAR-JEPA (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
+PIS (2024)  (?) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
+OReole-FM (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+PIEVIT (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SatVisionTOA (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+RS-vHeat (2024)  (?) $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
+Prithvi-EO-2.0 (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+AnySat (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+WildSAT (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SeaMo (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+FoMo (2025)  (?) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+SatMamba (2025)  (?) $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Galileo (2025)  (?) p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
+SatDiFuser (2025) (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+RoMA (2025)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Panopticon (2025)  (?) $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+FedSense (2025)  (?) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Copernicus-FM (2025)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+HyperFree (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+DynamicVis (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+FlexiMo (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Google Satellite (2025) (?) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+RingMoE (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$

### v5

@@
-TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis
+TESSERA: Precomputed FAIR Global Pixel Embeddings for Earth Representation and Analysis
 Abstract
-Satellite remote sensing enables a wide range of downstream applications, including habitat mapping, carbon accounting, and strategies for conservation and sustainable land use.
-However, satellite time series are voluminous and often corrupted, making them challenging to use.
-We present TESSERA, an open, global, land-oriented remote sensing foundation model that uses self-supervised learning to generate ‘ready-to-use’ embeddings at 10 m scale from pixel-level satellite time-series data.
-TESSERA uses two encoders to combine optical data with synthetic aperture radar backscatter coefficients at 10 m resolution to create embeddings that are fused with a multilayer perceptron to create annual global embedding maps.
-We compare our work with state-of-the-art task-specific models and other foundation models in five diverse downstream tasks and find that TESSERA closely matches or outperforms these baselines.
-We believe that TESSERA’s ease of use, state-of-the-art performance, openness, and computation- and labelled data-efficiency will prove transformative in a wide range of ecological applications.
-Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation  (?, ?, ?).
-Petabytes of EO data are available from sensors in various modalities, and declining launch costs promise continued growth in the future  (?).
-However, practitioners face a fundamental bottleneck: the scarcity of large, accurately labelled datasets required for supervised training  (?).
-The direct use of EO data also presents challenges as data quality is affected by sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling, preventing full and easy use of this valuable resource  (?, ?, ?).
-EO data users usually mitigate these issues by annual or seasonal compositing  (?), then using supervised machine learning to map from composited images to task-specific classification, such as crop identification  (?).
-However, compositing removes the critical temporal signal arising from phenology (i.e.
-timing of biological events such as flowering) and training a task-specific ML model is computationally onerous.
-Recent EO approaches tackle the latter problem using foundation models: general-purpose models pre-trained on a broad set of EO data  (?).
-These include both supervised pre-training on large labelled datasets for spatio-temporal analysis  (?, ?, ?, ?), and self-supervised learning (SSL) strategies that learn from the inherent structure of the data itself.
-Common SSL techniques include generative methods, such as masked autoencoders  (?, ?, ?, ?), diffusion-based models   (?, ?, ?), and contrastive methods  (?, ?, ?, ?).
-General purpose foundation models must be adapted (or ‘fine-tuned’) for specific downstream tasks such as habitat mapping and semantic segmentation  (?, ?, ?, ?, ?).
-Although several existing SSL approaches create rich multidimensional ‘embeddings’ from unlabelled data through auxiliary tasks that reveal latent patterns  (?, ?), they still do not address the loss of the temporal signal during compositing.
-In recent work, Lisaius et al.
-(?) have demonstrated that the Barlow Twins loss  (?), grounded in redundancy reduction principles  (?), can be adapted to preserve the spectral-temporal signal in latent embeddings when processing dense EO time series.
-Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale to enable state-of-the-art performance across a diverse array of complex tasks.
-Our experiments demonstrate that these TESSERA-derived embeddings outperform state-of-the-art methods in tasks as diverse as estimating canopy height in rainforests in Borneo, detecting crop types in Austria, mapping above-ground biomass in Finland, and identifying fire scars in California.
-Importantly, our outputs are ‘ready to use’: users need neither to fine-tune the model, nor download and process voluminous satellite data, nor perform inferencing other than what is needed to perform a downstream task.
-This makes TESSERA frugal in terms of both data handling and computation.
-Users can also keep their data private instead of having to upload it to a centralized repository.
-This removes a crucial barrier to use.
-We demonstrate that TESSERA is label-efficient: Only a relatively small number of labels for task-specific training is needed to achieve state-of-the-art performance.
-Moreover, our model is open source, enabling extensibility and scientific reproducibility.
-We have made a gap-free, global, annual, 10 m resolution embedding map for 2024 openly available through the Geotessera Python library  (?) and aim to eventually provide global maps for the entire period 2017-2024.
-These information-rich precomputed maps greatly simplify a wide range of downstream EO applications.
-We have also released interactive and batch tools for habitat mapping using TESSERA embeddings (see  (?) for details).
-To our knowledge, TESSERA is unprecedented in its ease of use, scale, and accuracy: no other foundation model provides analysis-ready outputs, is open, and provides global, annual coverage at 10 m resolution using only spectral-temporal features at pixel level (Supplementary Text table  LABEL:tab:sup_comparison).
-The work closest to ours is Google’s AlphaEarth foundation model  (?), which also provides global, annual, 10 m resolution embedding maps for 2017-2024.
-However, this work is neither open nor reproducible.
-Moreover, we demonstrate below that TESSERA embeddings match or outperform AlphaEarth embeddings in several downstream tasks.
-Overview of the TESSERA System
+Petabytes of satellite Earth Observation (EO) data are freely available and can address critical global challenges.
+However, EO data quality is poor due to clouds and variable lighting conditions.
+To address this, practitioners typically use compositing, but this critically removes the temporal phenological signal.
+Moreover, supervised machine learning to map composited pixels to task-specific classes requires accurately labelled data that are rarely available.
+We present TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings requiring only a few labels for task-specific training to achieve state-of-the-art performance across diverse complex tasks.
+TESSERA uses two encoders that combine optical data with synthetic aperture radar backscatter coefficients at 10m resolution, creating embeddings fused with a multilayer perceptron to generate annual global embedding maps.
+TESSERA closely matches or outperforms state-of-the-art task-specific models and other foundation models across five diverse downstream tasks.
+It is unprecedented in ease of use, scale, and accuracy: no other open foundation model provides precomputed outputs with global, annual coverage at 10m resolution.
+keywords:
+Satellite Remote Sensing, Earth Observation, Foundation Model
+1 Introduction
+Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [1, 2, 3].
+Although petabytes of EO data are available from sensors in various modalities [4], practitioners face fundamental bottlenecks: the scarcity of large, accurately labelled datasets required for supervised training [5], and poor data quality due to sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling [6, 7, 8].
+To mitigate these issues, users typically employ annual or seasonal compositing [9], then apply supervised machine learning to map from composited images to task-specific classification [10].
+However, compositing removes the critical temporal signal arising from phenology, and training task-specific ML models is computationally onerous.
+Recent EO approaches address the labelling problem using foundation models: general-purpose models pre-trained on broad EO datasets [11].
+These include supervised pre-training on large labelled datasets [12, 13, 14, 15], and self-supervised learning (SSL) strategies that learn from data structure itself.
+Common SSL techniques include masked autoencoders [16, 17, 18, 19], diffusion-based models [20, 21, 22], and contrastive methods [23, 24, 25, 26].
+These foundation models require fine-tuning for specific downstream tasks such as habitat mapping and semantic segmentation [27, 28, 29, 30, 31].
+Although existing SSL approaches create rich multidimensional embeddings from unlabelled data [32, 33], they still do not address temporal signal loss during compositing.
+Lisaius et al.
+[5] demonstrated that the Barlow Twins loss [34], grounded in redundancy reduction principles [35], can preserve spectral-temporal signals in latent embeddings when processing dense EO time series.
+Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale.
+TESSERA outperforms state-of-the-art methods across diverse tasks including estimating canopy height in Borneo rainforests, detecting crop types in Austria, mapping above-ground biomass in Finland, identifying fire scars in California and evaluating agroforestry stocking indices in Brazil.
+Critically, TESSERA provides precomputed outputs that are fully compliant with FAIR (Findable, Accessible, Interoperable, and Reusable) data principles [36].
+The embeddings are findable through a per-tile version-controlled registry, openly accessible without requiring inference, interoperable via standard numpy arrays and GeoTIFF metadata, and reusable under an open license with fully reproducible model training from open Sentinel-1/2 data.
+Users need neither fine-tune the model, download voluminous satellite data, nor perform inferencing beyond downstream tasks.
+This computational frugality, data privacy preservation, and FAIR compliance removes crucial barriers to use.
+TESSERA is label-efficient, requiring only small numbers of labels for task-specific training to achieve state-of-the-art performance, and is open source for extensibility and reproducibility.
+We provide a gap-free, global, annual, 10 m resolution embedding map for 2024 through the Geotessera Python library [37], with plans for global maps covering 2017-2024.
+We have also released interactive and batch tools for habitat mapping [38].
+TESSERA is unprecedented in combining ease of use, scale, and accuracy: no other foundation model provides precomputed outputs while being open and offering global, annual coverage at 10 m resolution using pixel-level spectral-temporal features (Supplementary Text Table 5).
+The closest work is Google’s AlphaEarth Satellite Embedding foundation model (GSE) [39], which provides similar global coverage but is neither open nor reproducible.
+We demonstrate that TESSERA embeddings match or outperform AlphaEarth embeddings in several downstream tasks.
+2 Methods
+Figure 1: Data processing pipeline.
+a, Overview.
+Time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are extracted for each pixel location to form mode-specific ‘d-pixels’ ( vertical rectangles; cloud-corrupted rows in gray).
+Two augmentations are created from each d-pixel as cloud-free random samples from the corresponding d-pixel.
+Augmentations are processed through a dual-branch architecture with weights shared between the branches.
+The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings.
+During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network.
+A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked days in the Sentinel-2 d-pixel.
+b, Detailed structure of a single encoder branch.
+Each d-pixel is embedded using a single-layer MLP (not shown) and a Day-of-Year-based temporal encoding is added.
+The resulting vector is processed by an 4-block Transformer Encoder and then a Gated Recurrent Unit (GRU) pooling layer; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
 TESSERA uses self-supervised learning to create per-pixel embeddings from a global dataset of unlabelled time-series of observations from Sentinel-1 Synthetic Aperture Radar (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams.
-It is trained by optimizing a modified Barlow Twins loss function  (?).
+It is trained by optimizing a modified Barlow Twins loss function [34].
 The original Barlow Twins approach seeks to create embeddings that are a) invariant to augmentations of its input, where these augmentations are typically rotations, reflections, and cropping of an image and b) whose dimensions are as decorrelated as possible.
 Inspired by the work by Lisaius et al.
-(?), we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions.
+[5], we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions.
 This is achieved by randomly sampling non-cloudy days from the Sentinel-2 annual observation time series so that model then learns to reconcile these partial views, interpolating missing data.
-The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Supplementary Text Equation S11).
+The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Supplementary Text Equation 11).
 We now discuss the system architecture.
@@
 Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context.
-An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality  (?).
+An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [38].
 See Figure 1 and Supplementary Text for architectural details.
@@
 This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate redundancy reduction during the self-supervised learning phase.
-A detailed breakdown of the model’s architectural parameters is provided in the Supplementary table S2.
-We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance   (?).
-Model Training
-The pre-training phase utilizes a large global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series.
+A detailed breakdown of the model’s architectural parameters is provided in the Supplementary table 2.
+We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [38].
+2.1 Model Training
+The pre-training phase utilizes a large multi-year global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series from the time period 2017-2024.
 To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed.
 This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features.
-The model was trained for one epoch in this dataset using distributed training techniques (see  (?) for further details on data preparation, the shuffling strategy, and optimization).
-The key hyperparameters used during pre-training are summarized in Supplementary table S3.
-Model Inference and Global Embeddings Maps
+The model was trained for one epoch in this dataset using distributed training techniques (see [38] for further details on data preparation, the shuffling strategy, and optimization).
+The key hyperparameters used during pre-training are summarized in Supplementary table 4.
+Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales.
+The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization.
+a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns.
+b, A multi-scale view of a complex agricultural landscape.
+b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields.
+c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
+2.2 Model Inference and Global Embeddings Maps
 Following pre-training, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally.
@@
 2).
-Downstream Tasks
+3 Downstream Tasks
 The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks.
@@
 (See Supplementary fig.
-S5 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings.
+12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings.
 Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling.
 This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
-Crop Type Classification
-Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector  (?).
-The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions  (?).
-However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes  (?).
-In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard  (?, ?).
-Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally  (?, ?).
+3.1 Crop Type Classification
+Figure 3: TESSERA embeddings achieve state-of-the-art performance in crop type classification.
+The evaluation was conducted on the 2022 Austrian crop dataset.
+a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio.
+TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google AlphaEarth Satellite Embedding (GSE).
+b, Patch-based semantic segmentation performance comparison.
+TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios.
+c, Weighted F1 score in few-shot learning scenarios.
+TESSERA maintains a significant performance advantage even with very few training samples per class.
+Error bars represent the standard deviation over multiple runs.
+d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes.
+TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
+Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [40].
+The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [7].
+However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [7].
+In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [7, 41].
+Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [42, 43].
 Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
-To evaluate performance, we used the INVEKOS Austrian crop dataset  (?) from the 2021–2022 growing season.
-This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets)  (?).
-We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model  (?), and (3) AlphaEarth Google Satellite Embedding (GSE)  (?).
+To evaluate performance, we used the INVEKOS Austrian crop dataset [44] from the 2021–2022 growing season.
+This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [44].
+We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [19], and (3) Google AlphaEarth Satellite Embedding (GSE) [39].
 For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison.
@@
 TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig.
+3c).
+Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig.
 3b).
-Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig.
-3c).
 These performance gains are due to the high quality of the learned embeddings.
-A UMAP analysis  (?) of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig.
+A UMAP analysis [45] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig.
 3d).
 This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
-Canopy Height Estimation
-Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+  (?, ?).
-It is correlated with ecosystem function  (?), forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness  (?, ?, ?).
+3.2 Canopy Height Estimation
+Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation.
+Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo.
+Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google AlphaEarth Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure.
+TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies.
+Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM.
+All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations.
+These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration.
+A 1:1 reference line (blue) is shown in each panel for comparison.
+Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [46, 47].
+It is correlated with ecosystem function [48], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [49, 50, 51].
 Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
 Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability.
-Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range  (?).
-GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes  (?), and does not produce unreliable estimates in mountainous terrain  (?).
-Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height  (?).
-Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors   (?, ?) - yet none of the three most widely used global CHM models currently implement sensor fusion  (?, ?, ?).
+Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [52].
+GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [53], and does not produce unreliable estimates in mountainous terrain [54].
+Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [55].
+Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [56, 57] - yet none of the three most widely used global CHM models currently implement sensor fusion [58, 59, 60].
 Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500 features in Potapov et al.
-(?)), which may generalize poorly across ecosystems.
-These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation  (?), and inconsistencies in input quality.
-As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes  (?).
+[59]), which may generalize poorly across ecosystems.
+These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [61], and inconsistencies in input quality.
+As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [62].
 Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
-These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling  (?, ?).
-We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo  (?) with openly available ground-truth data  (?).
+These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [63, 64].
+We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [65] with openly available ground-truth data [66].
 To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation.
@@
 The accuracy metrics for the foundation models are reported as the average across the 12 runs.
-TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62 m.
+TESSERA achieved a coefficient of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62 m.
 We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig.
@@
 First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.
-(?), the ETH map by Lang et al.
-(?), and the Meta map by Tolan et al.
-(?).
-Although local regression models, such as ours, generally outperform global products in regional evaluations  (?), the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages  (?, ?, ?)—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
+[59], the ETH map by Lang et al.
+[58], and the Meta map by Tolan et al.
+[60].
+Although local regression models, such as ours, generally outperform global products in regional evaluations [67], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [59, 58, 60]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
 We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
-(?), which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass.
-Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO  (?) and Google Satellite Embedding (GSE)  (?), using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
+[68], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass.
+Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [19] and Google AlphaEarth Satellite Embedding (GSE) [39], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
 Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
@@
 By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
-Above-ground Forest Biomass Estimation
-Forests play a central role in the global carbon cycle  (?).
-This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery  (?).
+3.3 Above-ground Forest Biomass Estimation
+Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings.
+a, TESSERA achieves lower RMSE values than SpectralGPT+UPerNet and results similar to the state-of-the-art supervised baseline and Google Satellite Embedding foundation model (top row).
+The coefficient of determination ($R^{2}$) and mean bias (MB) are shown in the middle and bottom row as functions of label availability.
+b, Mean predicted AGB values show increasing bias at higher ground-truth AGB values, which, however, are sparse in the dataset (the y-axis in the upper panel is on a logarithmic scale).
+Mean predictions are calculated for ground-truth AGB values grouped into 5 t/ha intervals.
+c, Predicted AGB map for an example patch using 4% of labels in training.
+d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (left) and the supervised baseline (right).
+The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
+Forests play a central role in the global carbon cycle [69].
+This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [70].
 A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions.
-Thus, there has been a long-standing effort to link field measurements to global satellite observations  (?).
-Recently, with the creation of the BioMassters benchmark dataset  (?) (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice   (?).
+Thus, there has been a long-standing effort to link field measurements to global satellite observations [71].
+Recently, with the creation of the BioMassters benchmark dataset [72] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [73].
 We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98.
-In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT  (?) and GSE  (?).
+In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [18] and GSE [39].
 First, a summary of the BioMassters dataset.
-This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations  (?, ?) informed by extensive field plots.
+This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [74, 75] informed by extensive field plots.
 Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel.
@@
 To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
-We compared a UNet-based regression model  (?) trained on TESSERA embeddings with SpectralGPT embeddings  (?) combined with a lightweight spatial-temporal encoder  (?) for temporal aggregation and a UPerNet  (?) regression head, because this combination has been found in the Pangaea benchmark  (?) to outperform other foundation models  (?, ?).
+We compared a UNet-based regression model [76] trained on TESSERA embeddings with SpectralGPT embeddings [18] combined with a lightweight spatial-temporal encoder [77] for temporal aggregation and a UPerNet [78] regression head, because this combination has been found in the Pangaea benchmark [79] to outperform other foundation models [72, 79].
 Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase.
-Thus, we adopted the Pangaea implementation of their model  (?), which uses preprocessed satellite imagery from the BioMassters dataset.
+Thus, we adopted the Pangaea implementation of their model [79], which uses preprocessed satellite imagery from the BioMassters dataset.
 We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline.
-This model uses a UNet with a temporal attention encoder  (?).
+This model uses a UNet with a temporal attention encoder [72].
 Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input.
@@
 Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
-Burned Area Detection
-The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change  (?, ?, ?).
-Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends  (?), predicting wildfire occurrence  (?), assessing ecosystem recovery  (?), and establishing effective fire management strategies  (?).
-Remote sensing is widely used to detect burned areas and assess severity regionally and globally  (?).
-Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input  (?, ?).
-These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals  (?).
-Most existing methods use optical sensors  (?), which can detect fire-specific spectral signatures but are susceptible to cloud cover  (?), particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens  (?).
-SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover  (?, ?, ?).
-Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness  (?, ?).
-Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products  (?, ?, ?).
-TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig.
-6).
-We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6).
-Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset  (?).
-UMAP projections  (?) of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters.
-In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a).
-To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels.
-We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites.
-The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b).
-Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency.
-These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
-To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps.
-These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Figure 6C).
-The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
-Stocking Indices in Voluntary Carbon Markets
-The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere  (?).
-To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
-This is a costly and inefficient undertaking if it is done only by field sampling.
-To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra  (?)) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB.
-Upon acceptance, project developers can use the index to assess the performance of the project.
-Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
-To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.
-(?).
-Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see   (?) for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil.
-This rescaling ensures a fair comparison between the various products.
-A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Figure 7).
-As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
-TESSERA: a Global Open-Source Foundation Model
-In this work, we introduce TESSERA, an innovative remote sensing foundation model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
-TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing.
-We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models.
-Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
-Figure 1: Data processing pipeline.
-a, Overview.
-Time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are extracted for each pixel location to form mode-specific ‘d-pixels’ ( vertical rectangles; cloud-corrupted rows in gray).
-Two augmentations are created from each d-pixel as cloud-free random samples from the corresponding d-pixel.
-Augmentations are processed through a dual-branch architecture with weights shared between the branches.
-The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings.
-During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network.
-A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked days in the Sentinel-2 d-pixel.
-b, Detailed structure of a single encoder branch.
-Each d-pixel is embedded using a single-layer MLP (not shown) and a Day-of-Year-based temporal encoding is added.
-The resulting vector is processed by an 4-block Transformer Encoder and then a Gated Recurrent Unit (GRU) pooling layer; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
-Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales.
-The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization.
-a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns.
-b, A multi-scale view of a complex agricultural landscape.
-b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields.
-c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
-Figure 3: TESSERA embeddings achieve state-of-the-art performance in crop type classification.
-The evaluation was conducted on the 2022 Austrian crop dataset.
-a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio.
-TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google Satellite Embedding (GSE).
-b, Patch-based semantic segmentation performance comparison.
-TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios.
-c, Weighted F1 score in few-shot learning scenarios.
-TESSERA maintains a significant performance advantage even with very few training samples per class.
-Error bars represent the standard deviation over multiple runs.
-d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes.
-TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
-Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation.
-Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo.
-Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure.
-TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies.
-Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM.
-All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations.
-These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration.
-A 1:1 reference line (blue) is shown in each panel for comparison.
-Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings.
-a, TESSERA achieves lower RMSE values than SpectralGPT+UPerNet and results similar to the state-of-the-art supervised baseline and Google Satellite Embedding foundation model (top row).
-The coefficient of determination ($R^{2}$) and mean bias (MB) are shown in the middle and bottom row as functions of label availability.
-b, Mean predicted AGB values show increasing bias at higher ground-truth AGB values, which, however, are sparse in the dataset (the y-axis in the upper panel is on a logarithmic scale).
-Mean predictions are calculated for ground-truth AGB values grouped into 5 t/ha intervals.
-c, Predicted AGB map for an example patch using 4% of labels in training.
-d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (left) and the supervised baseline (right).
-The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
+3.4 Burned Area Detection
 Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California.
@@
 unburned classification performance using random forest classifiers trained on different input representations.
-TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google Satellite Embeddings (GSE).
+TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google AlphaEarth Satellite Embeddings (GSE).
 c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas.
 As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
+The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [80, 81, 82].
+Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [83], predicting wildfire occurrence [84], assessing ecosystem recovery [85], and establishing effective fire management strategies [86].
+Remote sensing is widely used to detect burned areas and assess severity regionally and globally [87].
+Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [87, 88].
+These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [87].
+Most existing methods use optical sensors [87], which can detect fire-specific spectral signatures but are susceptible to cloud cover [89], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [90].
+SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [87, 91, 92].
+Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [87, 93].
+Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [91, 92, 94].
+TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig.
+6).
+We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Fig.
+6).
+Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [95].
+UMAP projections [45] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters.
+In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Fig.
+6a).
+To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels.
+We compared three input representations: (1) TESSERA embeddings, (2) Google AlphaEarth Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post- fire Sentinel-2 composites.
+The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Fig.
+6b).
+Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency.
+These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
+To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps.
+These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Fig.
+6c).
+The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
+3.5 Stocking Indices in Voluntary Carbon Markets
 Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market (data has been linearly rescaled).
 Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products.
-Note that the different products were originally provided in different units and converted to AGB (see   (?) for details).
-Acknowledgments
-We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
+Note that the different products were originally provided in different units and converted to AGB (see  [38] for details).
+The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [96].
+To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
+This is a costly and inefficient undertaking if it is done only by field sampling.
+To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [97]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB.
+Upon acceptance, project developers can use the index to assess the performance of the project.
+Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
+To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.
+[98].
+Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [38] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil.
+This rescaling ensures a fair comparison between the various products.
+A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Fig.
+7).
+As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
+4 Conclusion
+In this work, we introduce TESSERA, an innovative remote sensing foundation model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
+TESSERA embeddings establish new performance standards, while our open-source methodology, FAIR-compliant data distribution through version-controlled registries, standard geospatial formats, and the GeoTessera library permit reproducible science and enable use of our work even by non-specialists who lack access to high-performance computing.
+We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models.
+Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
+5 Acknowledgments
+We gratefully acknowledge support from Mantle Labs and help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
 Funding:
@@
 SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
-DAC and SK were funded by the NERC Centre for Landscape Regeneration ( NE/W00495X/1 )
+DAC and SK were funded by the NERC Centre for Landscape Regeneration (NE/W00495X/1).
+TJ was funded by NERC grant NE/S010750/1.
+MCL was funded by UKRI and Mantle Labs.
 Author contributions:
@@
 System implementation: ZF, RY, MCL, AM
-Data curation: ZF, RY, JK, SJ, AM, DAC, SK
+Data collection: TJ
+Data analysis: ZF, RY, JK, SJ, AM, TJ, DAC, SK
 Downstream tasks: MCL, CA, SS, DAC, CA, MI, ZF, JK
@@
 Data and materials availability:
-Our framework integrates openly available data from the Copernicus Sentinel-2 (S2)  (?) optical and Sentinel-1 (S1)  (?) SAR missions, leveraging the strengths of each sensor.
+Our framework integrates openly available data from the Copernicus Sentinel-2 (S2) [99] optical and Sentinel-1 (S1) [100] SAR missions, leveraging the strengths of each sensor.
 We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22).
 To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations.
-For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox  (?).
-The Google Satellite Embedding V1 dataset  (?) is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
-PRESTO embeddings can be created on demand in Google Earth Engine  (?) using VertexAI  (?).
-Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code  (?).
-To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository  (?), which is the official third-party software repository for Python.
-This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer  (?) that can be passed directly into downstream tasks.
-The embeddings themselves are checksummed and indexed using Pooch  (?) to check data integrity and handle retrieval.
-The crop type identification downstream task uses the open INVEKOS Austrian crop dataset  (?).
-The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset  (?).
-The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok  (?).
-The AGB downstream task uses the open BioMassters dataset  (?).
-The code for the winning model of the BioMassters competition is described in  (?) and freely available on the corresponding GitHub page.
-The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark  (?).
+For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [101].
+The Google Satellite Embedding V1 dataset [39] is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
+PRESTO embeddings can be created on demand in Google Earth Engine [102] using VertexAI [103].
+Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code [37].
+To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository [104], which is the official third-party software repository for Python.
+This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer [105] that can be passed directly into downstream tasks.
+The embeddings are indexed using Pooch [106] with per-tile checksums stored in an open Git repository, allowing users to pin specific versions for reproducibility or automatically refresh to the latest embeddings as needed.
+Additionally, a Parquet index is published daily for rapid spatial queries across the entire dataset.
+The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [44].
+The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset [95].
+The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [66].
+The AGB downstream task uses the open BioMassters dataset [72].
+The code for the winning model of the BioMassters competition is described in [72] and freely available on the corresponding GitHub page.
+The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark [79].
 The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners; hence, data access will require a Data Transfer Agreement.
-Supplementary materials
-Materials and Methods
-Figures S1 to S7
-Tables S1- S5
-References (97-0)
-Supplementary Materials for
-TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis
-[…]
-Equal contribution. Corresponding author. Email: sk818@cam.ac.uk.
-This PDF file includes:
-Materials and Methods
-Figures S1 to S7
-Tables S1- S5
-Materials and Methods
-Model details
+Appendix A Materials and Methods
+Table 1: Variable definitions for remote sensing data structure
+Variable Dimensions Description
+$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
+$V_{t}$ $W\times H$ Binary mask for tile at time $t$ (1=valid, 0=invalid/clouded)
+$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
+$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
+$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
+$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
+$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
+$W,H$ Scalar Spatial dimensions (width, height)
+$C$ Scalar Number of spectral channels + polarizations
+$T$ Scalar Number of time steps
+A.1 Model details
 Our input data consists of Sentinel-2 optical satellite imagery (Level-2A bottom-of-atmosphere) and Sentinel-1 Synthetic Aperture Radar (SAR, Radiometrically Terrain Corrected).
 Data Representation
-We consider remote sensing data  (?, ?) with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
-$R_{t}\in\mathbb{R}^{W\times H\times C}$ (S1)
+We consider remote sensing data [100, 99] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
+$R_{t}\in\mathbb{R}^{W\times H\times C}$ (1)
 where $W$ is the width (longitude dimension), $H$ is the height (latitude dimension), and $C$ is the number of spectral channels.
 Each tile is accompanied by a corresponding binary mask $V_{t}$ of dimensions:
-$V_{t}\in\{0,1\}^{W\times H}$ (S2)
+$V_{t}\in\{0,1\}^{W\times H}$ (2)
 where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
@@
 We stack spatially aligned tiles over a time period that spans $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
-$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (S3)
-$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (S4)
+$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (3)
+$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (4)
 Time Series Extraction and d-pixel Definition
 For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
-$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (S5)
+$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (5)
 We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
-$P_{i,j}(c)=S(i,j,c)$ (S6)
+$P_{i,j}(c)=S(i,j,c)$ (6)
 In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time.
@@
 The d-pixel Representation
-We represent each 1- m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
+We represent each 10 m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
 1 a).
@@
 To preserve sequence order and incorporate temporal context, learnable positional encodings based on the Day-of-Year (DOY) of each observation are added to these embeddings.
-The core of each encoder consists of a stack of 4 standard Transformer blocks  (?), featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
+The core of each encoder consists of a stack of 4 standard Transformer blocks [107], featuring multi-head self-attention and feed-forward layers to learn temporal patterns within the data streams.
 To derive a single vector summarizing the entire time series for each modality, an attention-pooling layer weighs the importance of different time steps before aggregation.
@@
 Each of the first five layers is a fully-connected linear layer followed by Batch Normalization and a ReLU non-linear activation function, making the network deep and highly non-linear.
-This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work  (?), to enable effective redundancy reduction during the self-supervised loss computation.
+This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [34], to enable effective redundancy reduction during the self-supervised loss computation.
 The final output of the projector is then used in the loss calculation.
@@
 The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
+Table 2: Architectural parameters of the TESSERA v1.0 model.
+Component Parameters Percentage of Total
+Encoder (Feature Extractor) 45,697,026 3.29%
+   Sentinel-1 Encoder 22,846,465 1.64%
+   Sentinel-2 Encoder 22,850,561 1.64%
+Fusion MLP 131,200 <<0.01%
+Projector (for Training only) 1,344,700,421 96.71%
+Total Parameters 1,390,528,647 100.00%
 Self-Supervised Training
-The TESSERA model is trained using a modified Barlow Twins objective function  (?).
+The TESSERA model is trained using a modified Barlow Twins objective function [34].
 For this objective, two distorted views, denoted as $Y_{A}$ and $Y_{B}$, are generated for each input d-pixel.
@@
 The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
-The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as  (?):
-$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (S7)
+The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as [34]:
+$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (7)
 Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
@@
 The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
-We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference  (?).
+We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [34].
 To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al.
-(?).
+[108].
 For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$.
@@
 The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
-$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (S8)
-$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (S9)
-$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (S10)
+$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (8)
+$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (9)
+$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (10)
 where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model output. The total loss function optimized during training is a weighted sum:
-$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (S11)
-where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by  (?)) to be effective based on our experiments.
-The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles.
+$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
+where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [108]) to be effective based on our experiments.
+The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles chosen from the years 2017-2024.
 d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400.
@@
 The temporal context of each sampled observation was encoded by transforming its normalized Day-of-Year (DOY) into sine and cosine features, which were then concatenated with the corresponding spectral or backscatter measurements.
+Figure 8: TESSERA is trained on globally-distributed training data.
+TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024.
+The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend.
+This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
 As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each).
@@
 The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training.
-The global batch size was 32768.
+The global batch size was 32,768.
 Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig.
-S2.
+9.
 A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
@@
 This pipeline is conceptually illustrated in Fig.
-S3a.
+10a.
 The impact of this approach on training stability is empirically demonstrated in Fig.
-S3b.
+10b.
 Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
 This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
+Figure 9: Training dynamics curves show that TESSERA rapidly reaches its asymptotes.
+a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss.
+b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay.
+c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction.
+d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant.
+e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
 Operationally, the process begins with the aggregation of d-pixels from all MGRS tiles into a single, comprehensive pool.
@@
 This end-to-end pipeline ensures that each batch presented to the model is a well-shuffled, globally diverse representation of Earth’s surface characteristics, which is fundamental for training a robust pixel-wise foundation model like TESSERA.
+Figure 10: Data shuffling greatly improves training stability.
+a, Schematic of the data shuffling and loading process.
+D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool.
+A custom Rust binary [37] performs a global shuffle on this multi-terabyte dataset before applying augmentations.
+The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches.
+b, Comparison of training loss curves.
+The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias.
+The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
 Global Embedding Map Generation
-A primary output of the TESSERA project is the generation of annual global embedding maps with 1- meter resolution for the years 2017-2024.
+A primary output of the TESSERA project is the generation of annual global embedding maps with 10 meter resolution for the years 2017-2024.
 To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used.
-For each 1- meter pixel on the globe and for each year:
+For each 10 meter pixel on the globe and for each year:
   1. 1.
-    The full Sentinel-1 and Sentinel-2 time series data at 1- meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
+    The full Sentinel-1 and Sentinel-2 time series data at 10 meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
   2. 2.
@@
     The outputs from the S1 and S2 encoders are fused by the MLP, producing a 128-dimensional embedding vector for that pixel for that year.
-This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 1- meter grid.
-Scaling with data and network size
+This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10 meter grid.
+A.2 Scaling with data and network size
 To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task.
 Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
-S4).
+11).
 Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
@@
 In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning.
-As detailed in Section Projector Network, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers.
+As detailed in Section A.1, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers.
 Our parameter sweep focused on this computationally intensive hidden block.
 We varied the number of these hidden layers (referred to as ’depth’ in Fig.
-S4a) and their width (number of neurons per layer), with results shown in Fig.
-S4a.
+11a) and their width (number of neurons per layer), with results shown in Fig.
+11a.
 As discussed by Zbontar et al.
-(?), deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task.
+[34], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task.
 The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide).
-However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure.
+However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 16,384 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure.
 For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
+Figure 11: Model performance improves with increasing projector size and additional computation.
+a, Hyperparameter sweep for the projector network’s architecture.
+Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis).
+The colour and size of the point indicate the validation F1 score and the number of parameters, respectively.
+Performance generally increases with both width and depth.
+The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints.
+The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency.
+b, Scaling laws for TESSERA.
+The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters).
+A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used.
+The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
 Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
-S4b.
+11b.
 The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
@@
 The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
+Figure 12: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications.
+Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
+Users can then fine-tune the entire model or only the head.
+Another approach (b) is where both the encoder and the projector (or decoder) are typically updated.
+In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input.
+This results in lightweight downstream models that consist only of a task-specific head.
 Ablation study
+Table 3: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance.
+We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset.
+The baseline represents the full model.
+Each subsequent row shows the result of removing the specified component.
+Performance is measured by validation F1 score and the RankMe score.
+Arrows ($\uparrow$) indicate that higher values are better.
+The results demonstrate that all components contribute positively to the model’s overall performance.
+Configuration Validation F1 Score ($\uparrow$) RankMe Score ($\uparrow$)
+Baseline (Full Model) 0.773 0.963
+w/o Global Data Shuffling 0.681 0.847
+w/o Mixup Regularization 0.662 0.857
+w/o Sentinel-1 Data 0.742 0.931
+w/o Shuffling & Mixup 0.626 0.867
+w/o Embedding Quantization 0.779 0.972
 To validate the contribution of each key component within the TESSERA framework, we performed an ablation study.
 We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
-Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score  (?).
+Performance was evaluated using two metrics: the standard validation overall F1 score, and a measure of the effective rank of the learned embeddings, which we refer to as the RankMe score [109].
 The RankMe score quantifies the richness of the embeddings by measuring how uniformly the information is distributed across the feature dimensions.
@@
 The RankMe score is the Shannon entropy of this distribution, normalized by the maximum possible entropy:
-$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (S12)
+$\mathcal{R}_{\text{RankMe}}=\frac{-\sum_{i=1}^{D}p_{i}\log(p_{i})}{\log(D)}$ (12)
 A score closer to 1 indicates that the embeddings are of higher effective rank, utilizing the full dimensionality of the feature space more effectively to create richer, less redundant representations.
-The results of the ablation study are summarized in Table S4.
+The results of the ablation study are summarized in Table 3.
 Our analysis reveals several key insights into the model’s architecture and training strategy.
@@
 This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
-Downstream Task Application Methodology
+A.3 Downstream Task Application Methodology
 A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data.
@@
 The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
-The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. S5):
+The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. 12):
   1. 1.
-    Download Embeddings for the Region of Interest: The Geotessera Python library  (?) allows users to download embeddings for a desired region and year in the form of a numpy array.
+    Download Embeddings for the Region of Interest: The Geotessera Python library [37] allows users to download embeddings for a desired region and year in the form of a numpy array.
   2. 2.
@@
 To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
-Downstream Task: Austrian Crop Classification
+A.4 Downstream Task: Austrian Crop Classification
 Dataset and Preprocessing
-We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season  (?).
+We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [44].
 The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation.
@@
 The MLP consisted of two hidden layers with 256 and 128 neurons, respectively, using ReLU activation functions, followed by a softmax output layer for the 17 classes.
-  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model  (?) to generate its pixel embeddings.
+  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model [19] to generate its pixel embeddings.
 These embeddings were then fed into an MLP head identical to the one used for TESSERA.
-  • GSE + MLP: To include another key foundation model, we used the publicly available Google Satellite Embedding (GSE) V1 dataset  (?).
+  • GSE + MLP: To include another key foundation model, we used the publicly available Google AphaEarth Satellite Embedding (GSE) V1 dataset [39].
 These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
@@
 For the experiments shown in the main text’s figure, we trained these models on randomly selected subsets of the data (from 1% to 30% for panel a; specified samples-per-class for panel b), using a fixed validation set for hyperparameter tuning and a held-out test set for final evaluation.
-Patch-wise Semantic Segmentation
-To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
-  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO  (?), and GSE  (?), a distinct workflow was used.
-We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$).
-These patches were then fed into a standard UNet architecture  (?), which learns to model the spatial relationships between the embeddings to produce a segmentation map.
-It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do.
-However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
-  • For other foundation models that are inherently patch-based (e.g., Prithvi  (?), Satlas  (?)), their encoders already process image patches.
-Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output.
-We did not freeze the encoders during the fine-tuning.
-Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. S6.
-Embedding Space Analysis
-The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP)  (?) algorithm to the embeddings of every pixel within the Austrian study area.
-UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology.
-It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure.
-UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
-In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
-To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
-  • The Silhouette score, $s(i)$, for a single data point $i$ measures how similar it is to its own cluster compared to other clusters. It is defined as:
-    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (S13)
-    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighbouring cluster.
-The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
-  • The Davies-Bouldin Index (DBI) evaluates clustering quality by computing the ratio of within-cluster scatter to between-cluster separation. For a set of $k$ clusters, it is defined as:
-    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (S14)
-    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$.
-Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
-Downstream Task: Canopy height estimation
-Study Areas and Ground Truth Datasets
-[…]
-Modelling methods
-To predict the canopy height from the different embeddings, we used a standard U-Net architecture  (?).
-It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations.
-For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube.
-The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values.
-The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches.
-At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation.
-In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training.
-Model training was done in Python using PyTorch.
-To ensure spatial generalization and avoid overfitting local patterns, the reference data were divided into training, validation, and test sets that are spatially disjoint.
-Each study region was divided into four spatial folds based on cardinal directions (bottom, top, left, right), each fold covering 50% of the area.
-For each fold, the remaining half of the region was used for training and validation.
-To account for inherent model randomness, we trained three independent models per fold, resulting in 12 trained models for each foundation model.
-This strategy captures variability due to both spatial heterogeneity and training randomness.
-The final performance of the model was evaluated on the test data held using the following metrics: coefficient of determination (R²), root mean square error (RMSE), and mean bias.
-Baselines
-TESSERA, PRESTO  (?), and GSE  (?) embeddings were compared by training identical models using each representation as input, with all other settings held constant.
-Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
-To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al.
-[2021] (?) (GLAD), Lang et al.
-[2023] (?) (ETH), and Tolan et al.
-[2024] (?) (Meta) as well as a regional map produced by Lang et al.
-[2021] (?).
-These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
-Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
-Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
-This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery.
-Cloud masking and compositing were performed in GEE, resulting in 12 channel input images, with one channel for each Sentinel-2 spectral band.
-Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
-Downstream Task: Burned area detection
-Study Areas and Ground Truth Datasets
-For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset  (?), referred to as Burn Area A and Burn Area B.
-Each Burn Area contains over two million 1 m pixels.
-Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure.
-In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
-Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land.
-In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
-To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set  (?).
-The dataset from the Monitoring Trends in Burn Severity (MTBS) program  (?) was used to identify burned areas and severity of the burn.
-Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
-Burn severity, defined as the measure of ecological damage caused by fire  (?), was derived from categorical and continuous severity metrics provided by MTBS.
-To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire.
-MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
-Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality  (?).
-Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations.
-The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine  (?).
-For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high).
-The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
-Dimensionality Reduction
-We used UMAP (Uniform Manifold Approximation and Projection)  (?) to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires.
-We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections.
-To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs.
-unburned), land cover class, or burn severity.
-These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
-To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
-UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space.
-UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
-Label Efficiency Analysis and Baselines
-We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available.
-The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set.
-We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%.
-The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance.
-Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
-We compared three types of input embeddings for the classification task.
-The first was TESSERA, using embeddings for 2020.
-The second was a task-specific composite baseline, constructed using prior knowledge of the fire event.
-This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector.
-The third type of input was Google Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset  (?).
-GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
-Downstream Task: Above-ground biomass estimation
-Dataset
-For the AGB regression task, we used the ground truth data from the BioMassters dataset  (?).
-This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution.
-Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre.
-Patch coordinates and collection years–provided by the authors of Ref.
-(?) upon request–were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE)  (?).
-UNet model architecture
-To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture  (?), using TESSERA or GSE embeddings   (?) as input.
-The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
-The network followed a standard UNet structure with two encoder and decoder stages.
-The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling.
-A 1024-channel bottleneck block was used prior to upsampling.
-The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double-convolution blocks.
-A final 1×1 convolution produced the single-channel regression output.
-The model had approximately 30 million trainable parameters.
-The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
-Following the evaluation protocol in Ref.
-(?), RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch.
-For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20).
-The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
-Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
-Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
-Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
-Baselines
-In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition  (?) and the SpectralGPT foundation model  (?) combined with a UPerNet  (?).
-For SpectralGPT, we used the implementation provided in Ref.
-(?), with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE  (?) for temporal aggregation, and a UPerNet as the regression head.
-This setup was found in  (?) to outperform other tested foundation models on the BioMassters dataset.
-To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
-Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss  (?) because it resulted in slightly better performance.
-Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
-For evaluating the supervised baseline, we downloaded the pre-trained weights of the BioMassters winning model and ran inference on the test set, applying a post-processing step to clip predictions to a non-negative range.
-Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model.
-According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
-Downstream Task: Stocking indices in Voluntary Carbon Markets
-Relative height modelling
-Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events.
-The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
-To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings.
-This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
-To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product  (?) from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil.
-From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables.
-A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
-The TerraClass Amazonia 2022 land cover map  (?) was used to distinguish between forest and non-forest.
-To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
-Stocking index baselines
-For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
-Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
-For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m)  (?), CTrees canopy tree height (5 m)  (?), ESA-CCI AGB (100 m)  (?), Meta CHM (1 m)   (?), and GTDX AGB (25 m)  (?).
-Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
-To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
-After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and $R^{2}$) directly comparable.
-The rescaling is as follows:
-  1. 1.
-    TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 $\times$ + 160.5
-  2. 2.
-[…]
-  3. 3.
-    CTrees canopy height converted using 1.806 $\times$ + 44.9
-  4. 4.
-    ESA AGB converted using 0.407 $\times$ + 34.0
-  5. 5.
-    Meta canopy height converted using 3.723 $\times$ + 46.6
-  6. 6.
-    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
-Interactive Habitat Mapping Tool
-Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure S7.
-The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles.
-To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
-The core of the tool is the interactive labelling workflow.
-A user can define custom habitat classes and provide training data by clicking on the map to place labelled points.
-These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier.
-The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view.
-This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise.
-We also provide a batch processing tool for larger regions of interest.
-The tools and their open source code are freely available in our github repository  (?).
-Figure S1: TESSERA is trained on globally-distributed training data.
-TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024.
-The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend.
-This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
-Figure S2: Training dynamics curves show that TESSERA rapidly reaches its asymptotes.
-a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss.
-b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay.
-c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction.
-d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant.
-e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
-Figure S3: Data shuffling greatly improves training stability.
-a, Schematic of the data shuffling and loading process.
-D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool.
-A custom Rust binary  (?) performs a global shuffle on this multi-terabyte dataset before applying augmentations.
-The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches.
-b, Comparison of training loss curves.
-The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias.
-The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
-Figure S4: Model performance improves with increasing projector size and additional computation.
-a, Hyperparameter sweep for the projector network’s architecture.
-Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis).
-The colour and size of the point indicate the validation F1 score and the number of parameters, respectively.
-Performance generally increases with both width and depth.
-The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints.
-The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency.
-b, Scaling laws for TESSERA.
-The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters).
-A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used.
-The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
-Figure S5: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications.
-Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
-Users can then fine-tune the entire model or only the head.
-Another approach (b) is where both the encoder and the projector (or decoder) are typically updated.
-In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input.
-This results in lightweight downstream models that consist only of a task-specific head.
-Figure S6: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
-Each row corresponds to a different patch size (32x32, 64x64, 128x128).
-The first column shows the ground truth segmentation.
-For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
-Figure S7: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
-We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
-Table S1: Variable definitions for remote sensing data structure
-Variable Dimensions Description
-$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
-$V_{t}$ $W\times H$ Binary mask for tile at time $t$ (1=valid, 0=invalid/clouded)
-$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
-$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
-$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
-$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
-$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
-$W,H$ Scalar Spatial dimensions (width, height)
-$C$ Scalar Number of spectral channels + polarizations
-$T$ Scalar Number of time steps
-Table S2: Architectural parameters of the TESSERA v1.0 model.
-Component Parameters Percentage of Total
-Encoder (Feature Extractor) 45,697,026 3.29%
-   Sentinel-1 Encoder 22,846,465 1.64%
-   Sentinel-2 Encoder 22,850,561 1.64%
-Fusion MLP 131,200 <<0.01%
-Projector (for Training only) 1,344,700,421 96.71%
-Total Parameters 1,390,528,647 100.00%
-Table S3: Pre-training hyperparameters for the TESSERA v1.0 model.
+Table 4: Pre-training hyperparameters for the TESSERA v1.0 model.
 Description Symbol Value
@@
 Quantization Bits (QAT) - 8
-Table S4: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance.
-We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset.
-The baseline represents the full model.
-Each subsequent row shows the result of removing the specified component.
-Performance is measured by validation F1 score and the RankMe score.
-Arrows ($\uparrow$) indicate that higher values are better.
-The results demonstrate that all components contribute positively to the model’s overall performance.
-Configuration Validation F1 Score ($\uparrow$) RankMe Score ($\uparrow$)
-Baseline (Full Model) 0.773 0.963
-w/o Global Data Shuffling 0.681 0.847
-w/o Mixup Regularization 0.662 0.857
-w/o Sentinel-1 Data 0.742 0.931
-w/o Shuffling & Mixup 0.626 0.867
-w/o Embedding Quantization 0.779 0.972
-Table S5: TESSERA provides unprecedented ease of use, scale, and accuracy.
+Patch-wise Semantic Segmentation
+To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
+  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [19], and GSE [39], a distinct workflow was used.
+We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$).
+These patches were then fed into a standard UNet architecture [76], which learns to model the spatial relationships between the embeddings to produce a segmentation map.
+It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do.
+However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
+  • For other foundation models that are inherently patch-based (e.g., Prithvi [110], Satlas [15]), their encoders already process image patches.
+Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output.
+We did not freeze the encoders during the fine-tuning.
+Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 13.
+Figure 13: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
+Each row corresponds to a different patch size (32x32, 64x64, 128x128).
+The first column shows the ground truth segmentation.
+For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
+Embedding Space Analysis
+The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [45] algorithm to the embeddings of every pixel within the Austrian study area.
+UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology.
+It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure.
+UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
+In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
+To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
+  • The Silhouette score, $s(i)$, for a single data point $i$ measures how similar it is to its own cluster compared to other clusters. It is defined as:
+    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (13)
+    where $a(i)$ is the mean distance between $i$ and all other points in the same cluster, and $b(i)$ is the mean distance from $i$ to all points in the nearest neighbouring cluster.
+The score ranges from -1 to 1, where a high value indicates that the object is well matched to its own cluster and poorly matched to neighboring clusters.
+  • The Davies-Bouldin Index (DBI) evaluates clustering quality by computing the ratio of within-cluster scatter to between-cluster separation. For a set of $k$ clusters, it is defined as:
+    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (14)
+    where $\sigma_{i}$ is the average distance of all points in cluster $i$ to their centroid $c_{i}$, and $d(c_{i},c_{j})$ is the distance between the centroids of clusters $i$ and $j$.
+Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
+A.5 Downstream Task: Canopy height estimation
+Study Areas and Ground Truth Datasets
+[…]
+Modelling methods
+To predict the canopy height from the different embeddings, we used a standard U-Net architecture [76].
+It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations.
+For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube.
+The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values.
+The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches.
+At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation.
+In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training.
+Model training was done in Python using PyTorch.
+To ensure spatial generalization and avoid overfitting local patterns, the reference data were divided into training, validation, and test sets that are spatially disjoint.
+Each study region was divided into four spatial folds based on cardinal directions (bottom, top, left, right), each fold covering 50% of the area.
+For each fold, the remaining half of the region was used for training and validation.
+To account for inherent model randomness, we trained three independent models per fold, resulting in 12 trained models for each foundation model.
+This strategy captures variability due to both spatial heterogeneity and training randomness.
+The final performance of the model was evaluated on the test data held using the following metrics: coefficient of determination (R²), root mean square error (RMSE), and mean bias.
+Baselines
+TESSERA, PRESTO [19], and GSE [39] embeddings were compared by training identical models using each representation as input, with all other settings held constant.
+Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
+To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al.
+[2021][59] (GLAD), Lang et al.
+[2023][58] (ETH), and Tolan et al.
+[2024][60] (Meta) as well as a regional map produced by Lang et al.
+[2021][68].
+These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
+Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
+Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
+This approach retained the same model architecture, patch size, and evaluation setup, but replaced embeddings with cloud-free annual median composites of Sentinel-2 imagery.
+Cloud masking and compositing were performed in GEE, resulting in 12 channel input images, with one channel for each Sentinel-2 spectral band.
+Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
+A.6 Downstream Task: Burned area detection
+Study Areas and Ground Truth Datasets
+For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [114], referred to as Burn Area A and Burn Area B.
+Each Burn Area contains over two million 1 m pixels.
+Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure.
+In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
+Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land.
+In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
+To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [115].
+The dataset from the Monitoring Trends in Burn Severity (MTBS) program [95] was used to identify burned areas and severity of the burn.
+Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
+Burn severity, defined as the measure of ecological damage caused by fire [116], was derived from categorical and continuous severity metrics provided by MTBS.
+To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire.
+MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
+Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [116].
+Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations.
+The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [102].
+For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high).
+The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
+Dimensionality Reduction
+We used UMAP (Uniform Manifold Approximation and Projection) [45] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires.
+We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections.
+To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs.
+unburned), land cover class, or burn severity.
+These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
+To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
+UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space.
+UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
+Label Efficiency Analysis and Baselines
+We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available.
+The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set.
+We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%.
+The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance.
+Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
+We compared three types of input embeddings for the classification task.
+The first was TESSERA, using embeddings for 2020.
+The second was a task-specific composite baseline, constructed using prior knowledge of the fire event.
+This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector.
+The third type of input was Google AlphaEarth Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [39].
+GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
+A.7 Downstream Task: Above-ground biomass estimation
+Dataset
+For the AGB regression task, we used the ground truth data from the BioMassters dataset [72].
+This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution.
+Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre.
+Patch coordinates and collection years–provided by the authors of Ref.
+[72] upon request–were used to retrieve satellite imagery to generate TESSERA and Google AlphaEarth Satellite Embeddings (GSE) [39].
+UNet model architecture
+To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [76], using TESSERA or GSE embeddings  [39] as input.
+The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
+The network followed a standard UNet structure with two encoder and decoder stages.
+The encoder comprised two convolutional blocks with 256 and 512 channels, each consisting of two 3×3 convolutions followed by batch normalization, ReLU activation, and 2×2 max pooling.
+A 1024-channel bottleneck block was used prior to upsampling.
+The decoder mirrored the encoder, using transposed convolutions for upsampling and skip connections from the encoder, followed by double-convolution blocks.
+A final 1×1 convolution produced the single-channel regression output.
+The model had approximately 30 million trainable parameters.
+The loss function was defined as the root mean squared error (RMSE) averaged over individual patches.
+Following the evaluation protocol in Ref.
+[72], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch.
+For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20).
+The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
+Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
+Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
+Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
+Baselines
+In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [72] and the SpectralGPT foundation model [18] combined with a UPerNet [78].
+For SpectralGPT, we used the implementation provided in Ref.
+[79], with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE [77] for temporal aggregation, and a UPerNet as the regression head.
+This setup was found in [79] to outperform other tested foundation models on the BioMassters dataset.
+To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation.
+Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [79] because it resulted in slightly better performance.
+Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
+For evaluating the supervised baseline, we downloaded the pre-trained weights of the BioMassters winning model and ran inference on the test set, applying a post-processing step to clip predictions to a non-negative range.
+Although we attempted to retrain the model from scratch using two H200 GPUs and a different CUDA version than in the original setup, this configuration led to inferior performance compared to the original pre-trained model.
+According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
+A.8 Downstream Task: Stocking indices in Voluntary Carbon Markets
+Relative height modelling
+Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events.
+The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
+To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings.
+This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
+To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [53] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil.
+From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables.
+A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
+The TerraClass Amazonia 2022 land cover map [117] was used to distinguish between forest and non-forest.
+To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
+Stocking index baselines
+For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
+Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
+For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [58], CTrees canopy tree height (5 m) [118], ESA-CCI AGB (100 m) [119], Meta CHM (1 m)  [60], and GTDX AGB (25 m) [120].
+Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
+To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
+After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and $R^{2}$) directly comparable.
+The rescaling is as follows:
+  1. 1.
+    TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 $\times$ + 160.5
+  2. 2.
+[…]
+  3. 3.
+    CTrees canopy height converted using 1.806 $\times$ + 44.9
+  4. 4.
+    ESA AGB converted using 0.407 $\times$ + 34.0
+  5. 5.
+    Meta canopy height converted using 3.723 $\times$ + 46.6
+  6. 6.
+    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
+A.9 Interactive Habitat Mapping Tool
+Figure 14: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
+We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
+Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure 14.
+The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles.
+To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
+The core of the tool is the interactive labelling workflow.
+A user can define custom habitat classes and provide training data by clicking on the map to place labelled points.
+These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier.
+The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view.
+This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise.
+We also provide a batch processing tool for larger regions of interest.
+The tools and their open source code are freely available in our github repository [37].
+Table 5: TESSERA provides unprecedented ease of use, scale, and accuracy.
 Unlike other remote sensing foundation models, TESSERA provides analysis-ready and gap-free outputs (R), is open (O), and provides global (G) annual (A) coverage at 10 m resolution (10 m).
@@
 TESSERA (2025) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-GeoKR (2021)  (?) $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-CVPRW (2021)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
-GASSL (2021)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
-SeCo (2021)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-DINO-MM (2022)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
-SatMAE (2022)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
-RS-BYOL (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
-GeCo (2022)  (?) $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
-RingMo (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-RVSA (2022)  (?) $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
-RSP (2022)  (?) $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
-MATTER (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-CSPT (2022)  (?) $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-CVPRW (2022)  (?) $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-BFM (2023)  (?) $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-TOV (2023)  (?) ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
-CMID (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-RingMo-Sense (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
-lal-SimCLR (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-CACo (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SatLas (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-GFM (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Scale-MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-DINO-MC (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-CROMA (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Cross-Scale MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-DeCUR (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
-Presto (2023)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-CtxMIM (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-FG-MAE (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Prithvi (2023)  (?) $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-RingMo-Lite (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-IGARSS (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-EarthPTi (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-USat (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-AIEarth (2023)  (?) p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Clay (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Hydro (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-U-Barn (2023)  (?) $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
-GeRSP (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SwiMDiff (2023)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-OFA-Net (2023)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-SML-FR (2023)  (?) p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-Spectral-GP (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-S2MAE (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SatMAE++ (2024)  (?) $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-msGFM (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SkySense (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-MTP (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-DOFA (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-MMEarth (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-LeMeViT (2024)  (?) $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
-SoftCon (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-RS-DFM (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-A2MAE (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-HyperSIGM (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SelectiveMAE (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-OmniSat (2024)  (?) $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
-MM-VSE (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
-MA3E (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SpectralEarth (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SenPa-MAE (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-RingMo-Ae (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SAR-JEPA (2024)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
-PIS (2024)  (?) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
-OReole-FM (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-PIEVIT (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SatVisionTOA (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-RS-vHeat (2024)  (?) $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
-Prithvi-EO-2.0 (2024)  (?) $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-AnySat (2024)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-WildSAT (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SeaMo (2024)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-FoMo (2025)  (?) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-SatMamba (2025)  (?) $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Galileo (2025)  (?) p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
-SatDiFuser (2025) (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-RoMA (2025)  (?) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Panopticon (2025)  (?) $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-FedSense (2025)  (?) $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Copernicus-FM (2025)  (?) $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-HyperFree (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-DynamicVis (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-FlexiMo (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Google Satellite (2025) (?) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-RingMoE (2025)  (?) $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+GeoKR (2021) [121] $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+CVPRW (2021) [122] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
+GASSL (2021) [23] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
+SeCo (2021) [123] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+MOSAIKS (2021) [124] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ ?
+DINO-MM (2022) [125] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
+SatMAE (2022) [16] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
+RS-BYOL (2022) [126] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
+GeCo (2022) [127] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
+RingMo (2022) [128] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+RVSA (2022) [129] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
+RSP (2022) [14] $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
+MATTER (2022) [130] $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+CSPT (2022) [131] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+CVPRW (2022) [132] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+BFM (2023) [133] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+TOV (2023) [134] ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
+CMID (2023) [135] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
+RingMo-Sense (2023) [136] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
+lal-SimCLR (2023) [137] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+CACo (2023) [138] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SatLas (2023) [15] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+GFM (2023) [139] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Scale-MAE (2023) [140] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+DINO-MC (2023) [141] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+CROMA (2023) [142] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Cross-Scale MAE (2023) [143] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+DeCUR (2023) [144] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
+Presto (2023) [19] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+CtxMIM (2023) [145] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+FG-MAE (2023) [146] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Prithvi (2023) [17] $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+RingMo-Lite (2023) [147] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+IGARSS (2023) [148] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+EarthPTi (2023) [149] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+USat (2023) [150] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+AIEarth (2023) [151] p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Clay (2023) [152] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Hydro (2023) [153] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+U-Barn (2023) [154] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
+GeRSP (2023) [155] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SwiMDiff (2023) [156] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+OFA-Net (2023) [157] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+SML-FR (2023) [158] p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+Spectral-GP (2024) [18] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+S2MAE (2024) [159] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SatMAE++ (2024) [160] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+msGFM (2024) [161] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+SkySense (2024) [24] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+MTP (2024) [162] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+DOFA (2024) [163] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+MMEarth (2024) [164] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+LeMeViT (2024) [165] $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
+SoftCon (2024) [166] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+RS-DFM (2024) [167] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
+A2MAE (2024) [168] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+HyperSIGM (2024) [169] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SelectiveMAE (2024) [170] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+OmniSat (2024) [171] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
+MM-VSE (2024) [172] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
+MA3E (2024) [173] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SpectralEarth (2024) [174] $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SenPa-MAE (2024) [175] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+RingMo-Ae (2024) [176] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SAR-JEPA (2024) [177] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
+PIS (2024) [178] $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
+OReole-FM (2024) [179] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+PIEVIT (2024) [180] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SatVisionTOA (2024) [181] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
+RS-vHeat (2024) [182] $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
+Prithvi-EO-2.0 (2024) [110] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
+AnySat (2024) [183] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+WildSAT (2024) [184] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+SeaMo (2024) [185] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+FoMo (2025) [186] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+SatMamba (2025) [187] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Galileo (2025) [188] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
+SatDiFuser (2025)[189] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+RoMA (2025) [190] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Panopticon (2025) [191] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+FedSense (2025) [192] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
+Copernicus-FM (2025) [193] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+HyperFree (2025) [194] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
+DynamicVis (2025) [195] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+FlexiMo (2025) [196] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
+Google Satellite (2025)[39] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+SkySense++ (2025) [197] $\times$ $\checkmark$ $\checkmark$ p p $\checkmark$ $\times$ p $\checkmark$ $\checkmark$ $\checkmark$
+RingMoE (2025) [198] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$

### v6

@@
-TESSERA: Precomputed FAIR Global Pixel Embeddings for Earth Representation and Analysis
+TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis
 Abstract
-Petabytes of satellite Earth Observation (EO) data are freely available and can address critical global challenges.
-However, EO data quality is poor due to clouds and variable lighting conditions.
-To address this, practitioners typically use compositing, but this critically removes the temporal phenological signal.
-Moreover, supervised machine learning to map composited pixels to task-specific classes requires accurately labelled data that are rarely available.
-We present TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings requiring only a few labels for task-specific training to achieve state-of-the-art performance across diverse complex tasks.
-TESSERA uses two encoders that combine optical data with synthetic aperture radar backscatter coefficients at 10m resolution, creating embeddings fused with a multilayer perceptron to generate annual global embedding maps.
-TESSERA closely matches or outperforms state-of-the-art task-specific models and other foundation models across five diverse downstream tasks.
-It is unprecedented in ease of use, scale, and accuracy: no other open foundation model provides precomputed outputs with global, annual coverage at 10m resolution.
-keywords:
-Satellite Remote Sensing, Earth Observation, Foundation Model
+Satellite Earth-observation (EO) time series in the optical and microwave ranges of the electromagnetic spectrum are often irregular due to orbital patterns and cloud obstruction.
+Compositing addresses these issues but loses information with respect to vegetation phenology, which is critical for many downstream tasks.
+Instead, we present TESSERA, a pixel-wise foundation model for multi-modal (Sentinel-1/2) EO time series that learns robust, label-efficient embeddings.
+During model training, TESSERA uses Barlow Twins and sparse random temporal sampling to enforce invariance to the selection of valid observations.
+We employ two key regularizers: global shuffling to decorrelate spatial neighborhoods and mix-based regulation to improve invariance under extreme sparsity.
+We find that for diverse classification, segmentation, and regression tasks, TESSERA embeddings deliver state-of-the-art accuracy with high label efficiency, often requiring only a small task head and minimal computation.
+To democratize access, adhere to FAIR principles, and simplify use, we release global, annual, 10m, pixel-wise int8 embeddings together with open weights/code and lightweight adaptation heads, thus providing practical tooling for large-scale retrieval and inference at planetary scale.
+The model training/inference code, downstream task code, and pre-generated embeddings can be accessed at https://github.com/ucam-eo.
+**footnotetext: Corresponding author.
 1 Introduction
-Monitoring Earth’s dynamic systems through satellite Earth Observation (EO) is critical to addressing global challenges including food security, forest and biodiversity loss, climate change, and disaster mitigation [1, 2, 3].
-Although petabytes of EO data are available from sensors in various modalities [4], practitioners face fundamental bottlenecks: the scarcity of large, accurately labelled datasets required for supervised training [5], and poor data quality due to sensor noise, cloud cover, atmospheric effects, and non-uniform temporal sampling [6, 7, 8].
-To mitigate these issues, users typically employ annual or seasonal compositing [9], then apply supervised machine learning to map from composited images to task-specific classification [10].
-However, compositing removes the critical temporal signal arising from phenology, and training task-specific ML models is computationally onerous.
-Recent EO approaches address the labelling problem using foundation models: general-purpose models pre-trained on broad EO datasets [11].
-These include supervised pre-training on large labelled datasets [12, 13, 14, 15], and self-supervised learning (SSL) strategies that learn from data structure itself.
-Common SSL techniques include masked autoencoders [16, 17, 18, 19], diffusion-based models [20, 21, 22], and contrastive methods [23, 24, 25, 26].
-These foundation models require fine-tuning for specific downstream tasks such as habitat mapping and semantic segmentation [27, 28, 29, 30, 31].
-Although existing SSL approaches create rich multidimensional embeddings from unlabelled data [32, 33], they still do not address temporal signal loss during compositing.
-Lisaius et al.
-[5] demonstrated that the Barlow Twins loss [34], grounded in redundancy reduction principles [35], can preserve spectral-temporal signals in latent embeddings when processing dense EO time series.
-Building on this insight, we introduce TESSERA, a pixel-oriented foundation model for EO time series that creates 128-dimensional latent embeddings at 10 m resolution and global scale.
-TESSERA outperforms state-of-the-art methods across diverse tasks including estimating canopy height in Borneo rainforests, detecting crop types in Austria, mapping above-ground biomass in Finland, identifying fire scars in California and evaluating agroforestry stocking indices in Brazil.
-Critically, TESSERA provides precomputed outputs that are fully compliant with FAIR (Findable, Accessible, Interoperable, and Reusable) data principles [36].
-The embeddings are findable through a per-tile version-controlled registry, openly accessible without requiring inference, interoperable via standard numpy arrays and GeoTIFF metadata, and reusable under an open license with fully reproducible model training from open Sentinel-1/2 data.
-Users need neither fine-tune the model, download voluminous satellite data, nor perform inferencing beyond downstream tasks.
-This computational frugality, data privacy preservation, and FAIR compliance removes crucial barriers to use.
-TESSERA is label-efficient, requiring only small numbers of labels for task-specific training to achieve state-of-the-art performance, and is open source for extensibility and reproducibility.
-We provide a gap-free, global, annual, 10 m resolution embedding map for 2024 through the Geotessera Python library [37], with plans for global maps covering 2017-2024.
-We have also released interactive and batch tools for habitat mapping [38].
-TESSERA is unprecedented in combining ease of use, scale, and accuracy: no other foundation model provides precomputed outputs while being open and offering global, annual coverage at 10 m resolution using pixel-level spectral-temporal features (Supplementary Text Table 5).
-The closest work is Google’s AlphaEarth Satellite Embedding foundation model (GSE) [39], which provides similar global coverage but is neither open nor reproducible.
-We demonstrate that TESSERA embeddings match or outperform AlphaEarth embeddings in several downstream tasks.
-2 Methods
-Figure 1: Data processing pipeline.
-a, Overview.
-Time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are extracted for each pixel location to form mode-specific ‘d-pixels’ ( vertical rectangles; cloud-corrupted rows in gray).
-Two augmentations are created from each d-pixel as cloud-free random samples from the corresponding d-pixel.
-Augmentations are processed through a dual-branch architecture with weights shared between the branches.
-The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings.
-During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network.
-A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked days in the Sentinel-2 d-pixel.
-b, Detailed structure of a single encoder branch.
-Each d-pixel is embedded using a single-layer MLP (not shown) and a Day-of-Year-based temporal encoding is added.
-The resulting vector is processed by an 4-block Transformer Encoder and then a Gated Recurrent Unit (GRU) pooling layer; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
-TESSERA uses self-supervised learning to create per-pixel embeddings from a global dataset of unlabelled time-series of observations from Sentinel-1 Synthetic Aperture Radar (SAR) and Sentinel-2 Multispectral Instruments (MSI), learning complementary information from the optical and microwave data streams.
-It is trained by optimizing a modified Barlow Twins loss function [34].
-The original Barlow Twins approach seeks to create embeddings that are a) invariant to augmentations of its input, where these augmentations are typically rotations, reflections, and cropping of an image and b) whose dimensions are as decorrelated as possible.
-Inspired by the work by Lisaius et al.
-[5], we use the Barlow Twins approach to learn embeddings that are invariant to the choice of random subsets of cloud-free observations in the Sentinel-2 optical time series, fused with the corresponding Sentinel-1 data, while minimizing redundancy between the learned feature dimensions.
-This is achieved by randomly sampling non-cloudy days from the Sentinel-2 annual observation time series so that model then learns to reconcile these partial views, interpolating missing data.
-The loss function includes the standard Barlow Twins terms for invariance and redundancy reduction, complemented with an additional mix-up regularization term to enhance robustness and inter-sample interaction during training (see Supplementary Text Equation 11).
-We now discuss the system architecture.
-TESSERA comprises two main components: a dual-branch encoder and a projector network (see Fig.
-1 for details).
-TESSERA models its input as a set of ‘d-pixels.’ Each d-pixel represents the annual time series of Sentinel-1 and Sentinel-2 observations at a specific 10 m pixel on Earth as a two-dimensional array ordered in time (that is, channels $\times$ timesteps).
-This format preserves the temporal evolution of the spectral/backscatter signatures inherent in the satellite data, while allowing observational gaps and irregularities, such as those due to cloud cover or orbit overlap.
-Rows that are cloud-corrupted or otherwise invalid are indicated by a binary mask and ignored when creating augmentations.
-Model training and inferencing is done by random temporal sampling followed by flattening of these d-pixels (see Fig.
-1 and Supplementary Text for details).
-In practice, we randomly sub-sample each d-pixel three times, to create a total of three pairs of augmentations.
-For each 10 m pixel, the fused TESSERA encoders generate a compact (8 bit), 128-dimensional embedding that encapsulates that pixel’s annual temporal and spectral characteristics.
-Under typical mid-latitude observation conditions, this compresses the input data by a factor of $\approx$ 24, with higher compression ratios at higher latitudes.
-TESSERA uses two separate, parallel Transformer-based encoders: one for Sentinel-1 SAR backscatter data in the C-band (VV and VH polarizations) and another for Sentinel-2 MSI data (10 spectral bands covering the visible, near, and shortwave infrared).
-Each encoder processes a d-pixel, utilizing multi-head self-attention mechanisms to capture the temporal dependencies and patterns specific to each data modality.
-Positional encodings derived from the Day-of-Year (DoY) are incorporated to provide temporal context.
-An attention pooling layer within each encoder aggregates the temporal features into a fixed-size vector that represents the annual signature for that modality [38].
-See Figure 1 and Supplementary Text for architectural details.
-Modality-specific embeddings derived from dual encoders are fused using a multilayer perceptron (MLP) which is trained alongside the encoders.
-This fused embedding, which serves as the 128-dimensional pixel embedding for downstream tasks, is then quantized before being fed into a high-dimensional projector network.
-This projector, a significantly larger MLP, expands the dimensionality of the fused representation to facilitate redundancy reduction during the self-supervised learning phase.
-A detailed breakdown of the model’s architectural parameters is provided in the Supplementary table 2.
-We have studied how the model’s performance on a typical downstream task improves as we increase the projector dimension (responsible for more than 95% of the model’s parameters) and the amount of computing as well as carried out an ablation study to understand which model components and design decisions are mainly responsible for performance  [38].
-2.1 Model Training
-The pre-training phase utilizes a large multi-year global dataset of approximately 800 million d-pixels, down-sampled 400:1 from Sentinel-1 and Sentinel-2 time series from the time period 2017-2024.
-To ensure successful generalization, a multilevel data sampling, shuffling, and batching strategy was developed.
-This ensured that each training mini-batch contained a highly diverse set of d-pixels from various geographic locations and acquisition conditions, preventing model overfitting and promoting the learning of universally applicable features.
-The model was trained for one epoch in this dataset using distributed training techniques (see [38] for further details on data preparation, the shuffling strategy, and optimization).
-The key hyperparameters used during pre-training are summarized in Supplementary table 4.
-Figure 2: TESSERA embeddings capture fine-grained geospatial patterns across multiple scales.
-The first three dimensions of the 128-dimensional embeddings are mapped to RGB channels for visualization.
-a, Examples from diverse global locations, showing a variety of landforms and vegetation patterns.
-b, A multi-scale view of a complex agricultural landscape.
-b.1 shows the overall region, with subsequent panels (b.2, b.3) zooming into the white-boxed areas, revealing progressively finer details down to individual crop fields.
-c, Continental-scale visualizations demonstrating the richness and consistency of the embeddings over vast areas, including (c.1) the Continental United States, (c.2) Europe, and (c.3) the Amazon basin.
-2.2 Model Inference and Global Embeddings Maps
-Following pre-training, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally.
+Figure 1: TESSERA advances the embedding-as-data approach for Earth Observation.
+It delivers (a) analysis-ready global-scale products, (b) pixel-wise, high-fidelity representations, and has (c) state-of-the-art downstream accuracy.
+Earth Observation (EO) satellites continuously observe our planet, yet cloud cover, irregular revisit intervals depending on latitude, sensors that differ in repeat period, spatial and spectral resolutions, and noise complicate the resulting time series.
+For many EO applications, such as agriculture, forestry, and environmental monitoring, clouds frequently obscure the evolution of target phenomena.
+Although compositing can yield cloud-free mosaics and reduce noise in microwave measurements, it flattens phenological dynamics and transient events that downstream models require [114].
+Recent remote-sensing foundation models (RSFMs) have made substantial progress in large-scale pretraining, yet most assume heavily preprocessed cloud-filtered inputs.
+These predominantly patch-based models train on composites or temporal averages, a simplification that discards partly-clouded observations that contain valuable temporal information.
+Consequently, learned embeddings often bias toward idealized, cloud-free conditions, limiting their robustness to the irregular, incomplete sampling that characterizes operational EO time series [107].
+We approach learning EO-based embeddings differently: rather than filtering imperfect observations, we use Barlow Twins [113, 52] to create embeddings that are robust to temporal sampling variability.
+Specifically, by forcing embeddings to be invariant to randomly selected subsets of cloud-free observations from the same location, our embeddings model consistent physical processes, which results in their generalization across sensors, seasons, and regions.
+Building on work by Lisaius et al.
+[52], which describes the use of Barlow Twins for EO for optical data streams, we introduce TESSERA, a pixel-wise RSFM for multi-modal (Sentinel-1/2) sequences.
+Our work differs from that of Lisaius et al.
+[52] in four additional ways.
+First, their work created embeddings from Sentinel 2 alone, whereas TESSERA combines observations from Sentinel 1 and 2 in a novel architecture.
+Second, we introduce two complementary regularizers: global shuffling to disrupt short-range spatial correlations, while mix-based regularization blends observations across time and samples to stabilize learning under severe sparsity.
+Third, unlike nearly all prior work, open-source global, annual, pixel-wise int8 embeddings for each 10m $\times$ 10m location on land, along with open-source model weights, code, and the GeoTessera Python library***https://github.com/ucam-eo/geotessera to efficiently retrieve embeddings from our data store for large-scale inference.
+TESSERA embeddings can be paired with lightweight adaptation heads (small MLPs or shallow convolutional models) to achieve high performance with minimal computation, enabling rapid prototyping and straightforward deployment.
+TESSERA thus emphasizes usability through the FAIR principles (Findable, Accessible, Interoperable, and Reusable).
+Finally, we evaluate TESSERA on a wide range of classification, segmentation, and regression benchmarks spanning multiple regions and label regimes.
+We find that our embeddings remain consistent when calendars shift, seasons are partially missing, or cloud cover is dense.
+Moreover, we achieve state-of-the-art accuracy with strong label efficiency, particularly under low-label conditions, and the TESSERA embeddings degrade gracefully under heavy cloud cover.
+Ablations confirm that temporal sampling invariance, global shuffling, and mix-based regulation contribute each to robustness and generalization.
+Our contributions are:
+  1. 1.
+    Global, pixel-wise, label-efficient embeddings. We use a novel self-supervised architecture to train a multi-modal (S1/S2) pixel-wise RSFM.
+  2. 2.
+    FAIR embeddings-as-data. We are releasing global, annual, 10 m, pixel-wise int8 embeddings with open-source weights, code, and a Python library—advancing FAIR deployment-ready resources.
+  3. 3.
+    Comprehensive evaluation.
+We compare our embeddings with a large number of other state-of-the-art foundation models and find that, for diverse classification, segmentation, and regression tasks, TESSERA embeddings deliver outstanding performance.
+2 Related Work
+2.1 Foundation Models for Earth Observation
+Remote-sensing foundation models (RSFMs) have rapidly advanced along several methodological lines.
+A large body of work has used self-supervised pretraining using either contrastive objectives (e.g., SeCo [59], SkySense [31]) or masked image modeling (e.g., SatMAE [15], Scale-MAE [71], Cross-Scale MAE [84], S2MAE [47], SatMAE++ [65]).
+Other variants leverage global–local view alignment or distilled objectives (e.g., DINO-MM/DINO-MC [98, 105]).
+Model scopes have expanded from single-modality encoders toward broader families such as RingMo and RingMo-Lite [81, 100], Satlas [8], Prithvi and Prithvi-EO-2.0 [76, 82], GFM/msGFM [61, 32], Skysense series [31, 106, 56, 121, 119].
+Work on multi-sensor fusion explicitly integrates optical and SAR (e.g., CROMA [29], SkySense [31], RingMo-Sense [112]).
+In parallel, early efforts like MOSAIKS [72] explored lightweight universal features to lower adoption barriers.
+Despite these advances, most RSFMs are trained on patch-based inputs and implicitly assume that the data have been preprocessed (e.g., aggressive cloud filtering or temporal averaging).
+Such assumptions simplify pretraining but discard partially cloudy images that still carry useful temporal signals, biasing representations toward idealized, cloud-free conditions.
+Moreover, strong performance typically requires fine-tuning the backbone for each downstream task, which can be compute- and label-intensive for many EO users.
+2.2 Embedding-Based Foundation Models
+A complementary line of work delivers embeddings-as-data: precomputed, generic, task-agnostic features that can be easily adapted with lightweight heads (shown in Figure 2 c).
+Representative models include Presto [88], which learns pixel-level temporal representations, and Google Deepmind’s AlphaEarth [11], which provides precomputed embeddings as a ready-for-analysis product.
+This paradigm reduces computational costs and democratizes access to high performance; however, existing approaches are often not fully open-sourced, lack pixel-level granularity at a global scale, or do not explicitly account for irregular temporal sampling.
+Our work follows the embedding-based paradigm while addressing these gaps.
+TESSERA introduces temporal sampling invariance to handle cloud- and orbital pattern induced irregularity and combines redundancy reduction with global shuffling and mix-based regulation for robustness.
+In contrast to prior embedding releases, we provide FAIR global, annual, 10 m pixel-wise embeddings, together with open weights, code, and GeoTessera python library, enabling analysis-ready use without backbone fine-tuning.
+Figure 2: Downstream adaptation paradigms.
+(a–b) Most RSFMs require fine-tuning the encoder or training task-specific decoders and often do not expose intermediate embeddings.
+(b-c) TESSERA produces fixed, task-agnostic embeddings that plug into lightweight heads, avoiding backbone fine-tuning.
+3 TESSERA
+Figure 3: Overview of the TESSERA processing pipeline.
+(a) Overall architecture: multi-temporal Sentinel-1/2 observations are converted into modality-specific d-pixels, augmented twice, and encoded by dual branches with shared weights to produce compact 128-D embeddings.
+(b) Zoom-in on the Sentinel-1 branch: temporal sequences are processed by a 4-block Transformer followed by GRU pooling to capture dynamic backscatter patterns before fusion into the joint TESSERA embedding.
+3.1 d-pixel: A Representation for Temporal Data
+We consider remote sensing data [25, 26] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
+$R_{t}\in\mathbb{R}^{W\times H\times C}$
+where $W$ is the width (longitude dimension), $H$ is the height (latitude dimension), and $C$ is the number of spectral channels.
+Each tile is accompanied by a corresponding binary mask $V_{t}$ of dimensions:
+$V_{t}\in\{0,1\}^{W\times H}$
+where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
+Temporal Data Stacking
+We stack resampled and spatially-aligned tiles over a time period that spans $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as
+$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$
+$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$
+Time Series Extraction and d-pixel Definition
+For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
+$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$
+We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
+$P_{i,j}(c)=S(i,j,c)$
+In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time.
+Note that d-pixels are usually sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there are valid data, with a value of 1 indicating that the corresponding row in $P_{i,j}$ is valid.
+The d-pixel formulation preserves the full temporal phenology while gracefully accommodating irregular sampling.
+3.2 Model Architecture
+TESSERA employs a simple yet efficient dual-branch encoder architecture to process optical and radar modalities separately before fusion.
+Modality Encoders.
+Each encoder receives a masked time series $\{(P_{i,j}^{(t)},m_{i,j}^{(t)})\}_{t=1}^{T}$.
+Valid observations are embedded via a linear projection $\phi:\mathbb{R}^{C}\to\mathbb{R}^{d}$, and temporal context is injected via learnable Day-of-Year (DoY) positional encodings:
+$e_{t}=\phi(P_{i,j}^{(t)})+\psi(\text{DoY}(t)),$
+where $\psi(\cdot)$ maps the normalized day of the year to a $d$-dimensional embedding space.
+The sequence $\{e_{t}\}_{t:m_{i,j}^{(t)}=1}$ is padded to a fixed length $L=40$ via sampling with replacement if necessary, then processed by a 4-layer Transformer encoder with multi-head self-attention:
+$H=\text{TransformerEncoder}([e_{t_{1}},\dots,e_{t_{L}}]).$
+A GRU-based pooling layer aggregates the temporal sequence into a fixed-size vector:
+$z_{\text{mod}}=\text{GRU}(H)\in\mathbb{R}^{128}.$
+Feature Fusion and Quantization.
+The optical embeddings ($z_{\text{S2}}$) and radar embeddings ($z_{\text{S1}}$) are concatenated and passed through a 2-layer MLP to produce the fused 128-dimensional embedding:
+$z=\text{MLP}([z_{\text{S2}};z_{\text{S1}}])\in\mathbb{R}^{128}.$
+[…]
+Projector Network.
+For training, $z$ is expanded to 16,384 dimensions using a deep projector MLP:
+$\hat{z}=\text{Projector}(z)\in\mathbb{R}^{16384},$
+which is discarded during inference.
+3.3 Pretraining
+TESSERA is pretrained on $\sim$800 million d-pixels sampled from 3,012 global MGRS tiles (2017–2024), using 16 AMD MI300X GPUs (192GB GPU RAM each).
+For each d-pixel, two augmented views $(Y_{A},Y_{B})$ are generated by independently sampling 40 valid time steps from the Sentinel-1 and Sentinel-2 time series.
+The model is trained for one epoch with a global batch size of 32,768 using AdamW ($\eta=0.002$, weight decay $10^{-6}$), with a linear warmup followed by cosine decay.
+The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
+The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as [113]:
+$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$
+Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
+The indices $i$ and $j$ iterate over the dimensions of the embedding vectors.
+The first term (invariance term) encourages similar embeddings for different views of the same input ($C_{ii}\to 1$).
+The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
+We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [113].
+To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al.
+[7].
+For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$.
+The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution in the unit interval $\alpha_{mix}\sim U(0,1)$.
+The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views.
+The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
+$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$
+$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$
+$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$
+where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model output. The total loss function optimized during training is a weighted sum:
+$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$
+where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [7]) to be effective based on our experiments.
+Crucially, we implement a global shuffling strategy that randomizes d-pixels across all geographic tiles before batching, smoothing the loss curve, preventing spatial autocorrelation, and improving generalization (see the supplementary materials).
+3.4 Inference
+Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10m pixel globally.
 This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders.
@@
 The primary output of this inference stage is a set of annual, global, 10m resolution TESSERA embedding maps.
-These maps are designed as readily usable multichannel geospatial data layers.
-This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Supplementary Text).
-We are currently generating global embeddings for 2024 (approx.
-250 TB) and will progressively extend coverage backward year by year until 2017.
-Visualizing the first three dimensions of the embeddings as R, G, and B channels reveals rich, detailed textures corresponding to diverse land cover and land use patterns.
-These visualizations highlight the model’s ability to capture information across multiple scales, from individual agricultural fields to entire continents (Fig.
-2).
-3 Downstream Tasks
-The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks.
-These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation.
-(See Supplementary fig.
-12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings.
-Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling.
-This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
-3.1 Crop Type Classification
-Figure 3: TESSERA embeddings achieve state-of-the-art performance in crop type classification.
-The evaluation was conducted on the 2022 Austrian crop dataset.
-a, Weighted and Macro F1 scores for pixel-wise classification as a function of the training data ratio.
-TESSERA embeddings, coupled with a simple MLP, consistently outperform a Random Forest model and other foundation models like PRESTO and Google AlphaEarth Satellite Embedding (GSE).
-b, Patch-based semantic segmentation performance comparison.
-TESSERA embeddings, used with a UNet model, achieve higher mean Intersection over Union (mIoU) and Macro F1 scores than other foundation models across various patch sizes and training data ratios.
-c, Weighted F1 score in few-shot learning scenarios.
-TESSERA maintains a significant performance advantage even with very few training samples per class.
-Error bars represent the standard deviation over multiple runs.
-d, UMAP visualization of the embedding spaces for TESSERA, PRESTO, and GSE for 17 crop classes.
-TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
-Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [40].
-The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [7].
-However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [7].
-In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [7, 41].
-Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [42, 43].
-Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
-To evaluate performance, we used the INVEKOS Austrian crop dataset [44] from the 2021–2022 growing season.
-This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [44].
-We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [19], and (3) Google AlphaEarth Satellite Embedding (GSE) [39].
-For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison.
-We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning.
-We also evaluated performance on a patch-based semantic segmentation task against several other foundation models.
-Details on data processing, model architectures, and training protocols are provided in the Supplementary Text.
-TESSERA consistently surpasses both baselines in all training regimes, especially in settings where labelled data are scarce (Fig.
-3).
-In pixel-wise classification tasks with training data splits from 1% to 30%, a simple MLP trained on TESSERA embeddings achieves higher weighted and macro F1 scores than an identical MLP trained on PRESTO or GSE embeddings, and substantially outperforms the RF baseline (Fig.
-3a).
-TESSERA also maintains statistically significant advantages in one-shot and few-shot scenarios (Fig.
-3c).
-Furthermore, TESSERA’s utility extends to patch-based semantic segmentation, where it consistently achieves higher mIoU and macro F1 scores compared to other leading foundation models, including PRESTO and GSE, across various patch sizes (Fig.
-3b).
-These performance gains are due to the high quality of the learned embeddings.
-A UMAP analysis [45] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig.
-3d).
-This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
-3.2 Canopy Height Estimation
-Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation.
-Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo.
-Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google AlphaEarth Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure.
-TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies.
-Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM.
-All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations.
-These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration.
-A 1:1 reference line (blue) is shown in each panel for comparison.
-Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [46, 47].
-It is correlated with ecosystem function [48], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [49, 50, 51].
-Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
-Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability.
-Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [52].
-GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [53], and does not produce unreliable estimates in mountainous terrain [54].
-Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [55].
-Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [56, 57] - yet none of the three most widely used global CHM models currently implement sensor fusion [58, 59, 60].
-Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500 features in Potapov et al.
-[59]), which may generalize poorly across ecosystems.
-These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [61], and inconsistencies in input quality.
-As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [62].
-Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration.
-These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [63, 64].
-We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [65] with openly available ground-truth data [66].
-To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation.
-A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input.
-To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set.
-The accuracy metrics for the foundation models are reported as the average across the 12 runs.
-TESSERA achieved a coefficient of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62 m.
-We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig.
-4).
-First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.
-[59], the ETH map by Lang et al.
-[58], and the Meta map by Tolan et al.
-[60].
-Although local regression models, such as ours, generally outperform global products in regional evaluations [67], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [59, 58, 60]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
-We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.
-[68], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass.
-Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [19] and Google AlphaEarth Satellite Embedding (GSE) [39], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models.
-Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
-In all comparisons, TESSERA outperformed competing methods.
-Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility.
-Instead, only the region and year of interest are needed to replicate the input data pipeline.
-By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
-3.3 Above-ground Forest Biomass Estimation
-Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings.
-a, TESSERA achieves lower RMSE values than SpectralGPT+UPerNet and results similar to the state-of-the-art supervised baseline and Google Satellite Embedding foundation model (top row).
-The coefficient of determination ($R^{2}$) and mean bias (MB) are shown in the middle and bottom row as functions of label availability.
-b, Mean predicted AGB values show increasing bias at higher ground-truth AGB values, which, however, are sparse in the dataset (the y-axis in the upper panel is on a logarithmic scale).
-Mean predictions are calculated for ground-truth AGB values grouped into 5 t/ha intervals.
-c, Predicted AGB map for an example patch using 4% of labels in training.
-d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (left) and the supervised baseline (right).
-The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
-Forests play a central role in the global carbon cycle [69].
-This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [70].
-A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions.
-Thus, there has been a long-standing effort to link field measurements to global satellite observations [71].
-Recently, with the creation of the BioMassters benchmark dataset [72] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [73].
-We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98.
-In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [18] and GSE [39].
-First, a summary of the BioMassters dataset.
-This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [74, 75] informed by extensive field plots.
-Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel.
-The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig.
-5b) that would not appear in coarser resolution data.
-To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
-We compared a UNet-based regression model [76] trained on TESSERA embeddings with SpectralGPT embeddings [18] combined with a lightweight spatial-temporal encoder [77] for temporal aggregation and a UPerNet [78] regression head, because this combination has been found in the Pangaea benchmark [79] to outperform other foundation models [72, 79].
-Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase.
-Thus, we adopted the Pangaea implementation of their model [79], which uses preprocessed satellite imagery from the BioMassters dataset.
-We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline.
-This model uses a UNet with a temporal attention encoder [72].
-Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input.
-However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
-Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
-Our performance metrics are the mean RMSE, the coefficient of determination ($R^{2}$), and the mean bias (MB) in AGB across the test patches as a function of the fraction of labelled data used to train the regressor (Fig.
-5a).
-We make three significant findings.
-First, although custom models developed using supervised learning are typically difficult to outperform, even with a small fraction of training labels, our regressor nearly matches the performance of the BioMassters competition winner, which uses a supervised learning approach.
-Second, our regressor trained on 1% of the labelled data outperforms a UPerNet-based regressor trained on the entire training set using SpectralGPT embeddings.
-Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
-3.4 Burned Area Detection
-Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California.
-a, UMAP projection of $20$ TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events.
-A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance.
-b, Burned vs.
-unburned classification performance using random forest classifiers trained on different input representations.
-TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google AlphaEarth Satellite Embeddings (GSE).
-c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas.
-As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
-The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [80, 81, 82].
-Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [83], predicting wildfire occurrence [84], assessing ecosystem recovery [85], and establishing effective fire management strategies [86].
-Remote sensing is widely used to detect burned areas and assess severity regionally and globally [87].
-Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [87, 88].
-These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [87].
-Most existing methods use optical sensors [87], which can detect fire-specific spectral signatures but are susceptible to cloud cover [89], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [90].
-SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [87, 91, 92].
-Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [87, 93].
-Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [91, 92, 94].
-TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig.
-6).
-We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Fig.
-6).
-Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [95].
-UMAP projections [45] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters.
-In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Fig.
-6a).
-To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels.
-We compared three input representations: (1) TESSERA embeddings, (2) Google AlphaEarth Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post- fire Sentinel-2 composites.
-The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Fig.
-6b).
-Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency.
-These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
-To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps.
-These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Fig.
-6c).
-The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
-3.5 Stocking Indices in Voluntary Carbon Markets
-Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market (data has been linearly rescaled).
-Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products.
-Note that the different products were originally provided in different units and converted to AGB (see  [38] for details).
-The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [96].
-To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas.
-This is a costly and inefficient undertaking if it is done only by field sampling.
-To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [97]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB.
-Upon acceptance, project developers can use the index to assess the performance of the project.
-Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM.
-To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.
-[98].
-Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [38] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil.
-This rescaling ensures a fair comparison between the various products.
-A comparison of R2 and RMSE demonstrates that the TESSERA-derived stocking index consistently outperforms alternative products (Fig.
-7).
-As the alternative stocking indices all use similar (optical and microwave) inputs to derive canopy height, respectively, AGB, the superior performance of the TESSERA-derived stocking index must be attributed to the efficiency of the TESSERA embeddings.
+These maps are designed as readily usable multidimensional geospatial data layers.
+This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing, running the TESSERA model itself (see supplementary materials).
+Inference is performed on 96 Intel Data Center GPU Max 1550s (128 GB GPU RAM each), processing $\sim$5 PB of raw Sentinel data per year from Microsoft Planetary Computer.
+Each year yields $\sim$300 TB of int8-quantized embeddings, with 2024 released first, followed by 2017–2023.
+We release embeddings through the pip-installable GeoTessera Python library, enabling easy access and integration into downstream workflows.
+4 Experiments
+Table 1: Results across tasks.
+(a) Classification: F1 $\uparrow$ on TreeSatAI-TS and Austrian Crop (left column; TreeSatAI-TS shows original/finetuned).
+(b) Segmentation: mIoU $\uparrow$ on PASTIS-R and Austrian Crop (original/finetuned for CROMA–ViT Baseline; single evaluation for Presto, AlphaEarth, TESSERA).
+(c) Regression: RMSE $\downarrow$ on Biomassters and Borneo CHM (original/finetuned for baselines; single evaluation otherwise).
+best; second.
+(a1) Classification — TreeSatAI-TS (F1 $\uparrow$)
+Model TreeSatAI-TS 1% 30% All CROMA[29] 39.73/43.14 62.44/63.24 70.21/72.09 DOFA[109] 35.04/40.15 59.23/63.46 66.88/67.20 Prithvi[76] 37.21/41.89 60.01/60.66 65.86/67.04 RemoteCLIP[53] 34.94/37.77 57.52/59.70 68.32/68.03 SatlasNet[8] 33.86/37.83 56.80/61.31 66.27/67.85 Scale-MAE[71] 36.20/40.96 59.26/61.29 70.20/73.12 SpectralGPT[34] 41.82/45.66 59.77/60.64 71.04/72.23 Galileo[89] 39.99/42.38 59.42/61.94 69.44/72.17 Skysense[31] 39.72/41.39 61.88/64.80 69.35/72.24 UNet Baseline[73] 39.88/40.43 60.03/62.22 71.39/73.30 ViT Baseline[22] 40.52/41.70 61.79/65.62 72.24/72.60 Presto[88] 46.02 61.45 67.81 AlphaEarth[11] 52.79 71.94 76.90 TESSERA (Ours) 60.58 75.42 77.96
+(a2) Classification — Austrian Crop (F1 $\uparrow$)
+Model Austrian Crop 1% 3% 5% 7% 10% 15% 20% 30% Presto[88] 32.74 37.63 40.26 41.50 43.94 46.01 49.74 57.89 AlphaEarth[11] 37.22 40.06 42.73 45.08 48.01 50.47 52.47 56.36 TESSERA (Ours) 66.15 71.62 74.32 77.22 78.69 80.12 80.79 82.09
+(b) Segmentation — PASTIS-R / Austrian Crop (mIoU $\uparrow$)
+Model PASTIS-R Austrian Crop 1% 30% All 1% 30% All CROMA[29] 7.61/12.33 21.62/23.78 36.24/41.29 5.64/5.19 12.89/15.76 17.10/18.28 DOFA[109] 4.82/5.53 19.28/21.74 25.94/29.55 4.90/4.92 10.05/14.00 17.14/17.39 Prithvi[76] 8.24/11.01 24.86/29.51 34.09/39.52 4.36/3.82 11.12/14.35 16.25/18.60 RemoteCLIP[53] 5.50/7.55 11.72/16.40 20.17/21.03 3.52/3.37 7.76/10.12 11.16/15.15 SatlasNet[8] 9.54/11.58 17.86/21.48 21.02/31.38 4.27/4.05 8.45/9.64 15.61/18.97 Scale-MAE[71] 5.57/7.86 16.34/18.48 23.96/28.51 4.48/4.93 7.42/10.31 13.75/15.22 SpectralGPT[34] 11.25/15.95 26.23/29.33 35.10/40.28 4.85/4.17 13.77/15.20 20.58/22.84 Galileo[89] 9.78/11.83 15.63/17.91 27.92/35.28 9.95/9.03 14.24/15.89 22.80/26.16 Skysense[31] 13.42/14.47 28.54/32.54 32.87/36.41 8.93/9.38 15.29/16.12 26.82/31.08 UNet Baseline[73] 13.52/16.44 25.75/30.08 30.16/37.56 10.64/11.40 14.95/15.10 27.00/30.57 ViT Baseline[22] 12.58/15.72 29.43/34.07 37.37/42.57 10.88/10.03 15.26/16.75 30.72/31.77 Presto[88] - - - 14.25 27.93 34.04 AlphaEarth[11] 27.12 46.42 51.08 14.64 21.75 25.70 TESSERA (Ours) 27.54 46.04 50.68 28.20 48.29 53.12
+(c) Regression — Biomassters / Borneo CHM (RMSE $\downarrow$)
+Model Biomassters Borneo CHM 1% 30% All 1% 30% All CROMA[29] 46.58/44.28 39.24/36.80 37.14/35.07 44.52/41.89 31.25/29.67 20.56/19.48 DOFA[109] 51.86/47.84 45.72/43.59 43.05/41.10 43.96/40.68 28.97/27.93 17.63/17.51 Prithvi[76] 43.89/39.37 41.49/38.68 41.12/38.88 43.50/44.27 29.42/27.80 19.71/19.41 RemoteCLIP[53] 61.70/61.01 52.13/51.32 49.21/47.92 43.24/43.10 32.78/30.78 28.30/27.59 SatlasNet[8] 49.51/44.64 44.24/43.23 41.93/41.36 45.85/43.01 34.01/32.26 22.03/21.02 Scale-MAE[71] 56.41/51.51 51.02/49.34 47.05/45.17 47.51/46.72 32.00/31.11 21.12/20.41 SpectralGPT[34] 64.81/62.28 38.54/36.12 36.87/36.12 43.91/43.65 26.03/25.08 16.53/15.89 Galileo[89] 52.72/49.15 39.55/36.82 38.52/36.92 44.82/42.78 32.18/31.92 20.48/20.27 Skysense[31] 47.52/42.28 37.51/37.10 32.52/30.78 42.74/40.96 27.42/26.65 16.56/15.58 UNet Baseline[73] 44.62/41.70 37.92/36.16 36.41/35.17 41.81/38.61 28.01/26.02 17.79/16.87 ViT Baseline[22] 48.24/49.42 38.56/38.27 37.98/37.52 42.02/39.95 27.87/26.61 16.48/15.77 Presto[88] - - - 36.43 20.95 17.88 AlphaEarth[11] 35.45 32.33 29.59 36.03 20.62 16.11 TESSERA (Ours) 32.50 29.15 27.43 34.82 15.18 12.21
+4.1 Experimental Setup
+Datasets.
+We evaluate TESSERA on six representative benchmarks that span classification, segmentation, and regression tasks. All datasets use multi-temporal Sentinel-1/2 imagery:
+  • TreeSatAI-TS: Tree species classification in Germany.
+  • PASTIS-R: Parcel segmentation in France.
+  • Austrian Crop (Classification): Crop type classification in Austria [1].
+  • Biomassters: Above-Ground Biomass (AGB) regression in Finland [63].
+  • Borneo Canopy Height: Canopy height regression in Malaysia [16].
+  • Austrian Crop (Segmentation): Crop semantic segmentation in Austria [1].
+To ensure comprehensive evaluation, we select two datasets per task whenever possible: one large-scale (e.g., continental or national coverage) and one fine-grained regional dataset.
+This design validates both geographic transferability and fine-detail modeling capability.
+High-resolution multi-temporal S1+S2 datasets with dense labels are scarce; hence, we additionally contribute two new benchmarks: Austrian Crop (parcel-level crop mapping) and Borneo Canopy Height (LiDAR-calibrated forest structure in Southeast Asia).
+Baselines.
+We compare against a comprehensive suite of representative remote-sensing foundation models (RSFMs): CROMA [29], DOFA [109], Prithvi [76], RemoteCLIP [53], SatlasNet [8], Scale-MAE [71], SpectralGPT [34], Galileo [89], Skysense [31], Presto [88], and AlphaEarth  [11].
+We also include standard UNet [73] and ViT [22] baselines for reference.
+Table 2: Model specification. Multi-temporal and multi-modal (multiple satellite sources) flags, venue, and parameter size. Params (M) indicates encoder frozen / fully fine-tuned.
+Model Multi-temporal Multi-modal Publication Params (M)
+CROMA $\times$ NeurIPS’23 49.1 / 704.9
+DOFA $\times$ arXiv’24 40.6 / 151.7
+Prithvi arXiv’24 40.6 / 126.8
+RemoteCLIP $\times$ TGRS’24 40.6 / 128.1
+SatlasNet ICCV’23 33.3 / 121.2
+Scale-MAE $\times$ $\times$ ICCV’23 49.1 / 352.2
+SpectralGPT $\times$ TPAMI’24 183.5 / 268.9
+Galileo ICML’25 39.4 / 125.9
+Skysense CVPR’24 36.5 / 341.5
+UNet $\times$ $\times$ MICCAI’15 14.8 / 14.8
+ViT-B/16 $\times$ $\times$ ICLR’21 125.1 / 125.1
+Presto arXiv’23 30.2 / 30.2
+AlphaEarth arXiv’25 30.1 / 30.1
+TESSERA (Ours) 30.2 / 30.2
+Implementation.
+For embedding-based models (TESSERA, Presto, AlphaEarth), we train a 2-layer MLP for pixel-wise classification, and a lightweight UNet for patch-based classification, segmentation, and regression.
+Patch-based models are fine-tuned using UPerNet heads.
+All models are evaluated at 1%, 30% and 100% label ratios to study label efficiency.
+Additional implementation details are in the supplementary material.
+4.2 Classification Results
+Patch-based classification: TreeSatAI-TS.
+Table 1 (a1) reports results on TreeSatAI-TS, a national-scale forest species benchmark.
+TESSERA achieves a state-of-the-art F1-score of 77.96 with full supervision, significantly outperforming all other models, including the next-best competitor, AlphaEarth (76.90)***The TESSERA encoder (45.7M parameters) used during these tasks is also significantly smaller than AlphaEarth’s (480M parameters).
+and the strongest fine-tuned baseline, the UNet Baseline (73.30).
+This highlights how our pixel-level temporal modeling effectively captures phenological variations, which improve class separability [37].
+Notably, TESSERA’s superiority is even more pronounced in the few-shot setting: at 1% labels, it achieves an F1 score of 60.58 , surpassing the second-best model, AlphaEarth (52.79), by nearly 8 points.
+Pixel-wise crop classification: Austrian Crop.
+On the Austrian Crop dataset (Table 1 (a2)), TESSERA demonstrates strong performance.
+As visualized in Figure 4 (a), it consistently surpasses Presto and AlphaEarth in all standard label ratios (1%-30%).
+The performance gap is most pronounced in the low-data regime: at just 1% labeled data, TESSERA achieves a 66.15 F1 score.
+This outperforms AlphaEarth (37.22) by more than 28.9 points and Presto (32.74) by more than 33.4 points.
+We further investigate extreme label scarcity in Figure 4 (b), evaluating performance with only 1 to 20 labeled samples per class.
+In this challenging few-shot scenario, TESSERA’s performance remains robust and scales gracefully (e.g., $\sim$0.5 F1 with just 4 samples/class), while the F1 scores of other models are below 0.4.
+Figure 4: Label-efficient crop classification on Austrian Crop.
+(a) Weighted F1 vs.
+training label ratio using a small head on frozen embeddings.
+(b) Few-shot performance.
+TESSERA attains strong accuracy with very few labels; error bars denote variation over runs.
+4.3 Segmentation Results
+Field parcel segmentation: PASTIS-R.
+On the large-scale PASTIS-R benchmark (Table 1 (b)), TESSERA achieves competitive performance.
+With full labels, it obtains 50.68 mIoU, ranking second and just behind AlphaEarth (51.08).
+However, in the low-label regime (1% labels), TESSERA (27.54) surpasses AlphaEarth (27.12) and significantly outperforms all fine-tuned baselines, such as the UNet Baseline (16.44) and Skysense (14.47).
+Unlike some RSFMs that require compute-heavy decoders (>>100M parameters), our compact 30M-parameter UNet suffices.
+This demonstrates that spatial context can be learned efficiently using TESSERA’s pre-trained temporal embeddings.
+Regional crop segmentation: Austrian Crop.
+For the fine-grained Austrian Crop segmentation task, TESSERA achieves 53.12 mIoU, establishing a new SOTA and significantly improving over both AlphaEarth (25.70) by +27.4 points and Presto (34.04) by +19.0 points.
+This performance gain is consistent across all label regimes, especially at 1% labels (TESSERA: 28.20 vs AlphaEarth: 14.64), confirming the effective transferability of temporally coherent embeddings.
+TESSERA yields sharper parcel edges and reduces confusion among visually similar crops (see supplementary materials), indicating its ability to maintain semantic consistency over time.
+4.4 Regression Results
+Above-ground biomass estimation: Biomassters.
+As shown in Table 1 (c), TESSERA achieves an RMSE of 27.43 t/ha with full labels – outperforming all other foundation models, including the next-best, AlphaEarth (29.59 t/ha).
+This result surpasses the strongest fine-tuned RSFM baseline, Skysense (30.78 t/ha).
+The results demonstrate that temporal embeddings capture biomass dynamics effectively, even under limited supervision.
+Figure 5 provides a detailed comparison for this task after clipping implausibly high biomass values above 500 t/ha to this threshold.
+In this analysis, TESSERA (26.61 t/ha) nearly matches the best-performing supervised task-specific model (25.90 t/ha), which won the Biomassters competition of more than 1000 submissions.
+The figure further illustrates that TESSERA maintains consistent performance in all label ratios, shows minimal bias in all biomass ranges, and generates spatially coherent AGB maps with only 4% of total labels.
+Canopy height regression: Borneo Canopy Height.
+In the Borneo CHM dataset, TESSERA again demonstrates superior performance, achieving the lowest RMSE of 12.2 m.
+This result significantly outperforms all other models, including AlphaEarth (16.1 m), Presto (17.9 m), and the best-performing fine-tuned RSFM such as Skysense (15.6 m) and the ViT Baseline (15.8 m).
+We also compare with DINOv2-derived and Sentinel-2 composite height maps, where TESSERA demonstrates the closest agreement with LiDAR ground truth (see supplementary materials).
+[…]
+5 Discussion & Ablation Studies
+Figure 6: Robustness and sensitivity analyses.
+(a) Cloud robustness: performance remains stable until Sentinel-2 valid observations per year drop below $\sim$20, after which accuracy degrades markedly.
+(b) Finetuning impact: evaluating the effect of three different learning rate strategies during finetuning.
+(c) Embedding specificity: comparing the performance of a global embedding versus an embedding retrained on region-specific data.
+(d) Temporal sampling uncertainty: selecting L=40L{=}40 timesteps within a year produces consistently similar results across random choices.
+Figure 7: Scaling behavior.
+(a) Projector depth/width sweep: validation F1 improves with capacity; † marks the max-F1 setup under memory constraints; * denotes the chosen TESSERA v1.0.
+(b) Validation F1 scales predictably with training FLOPs and parameters, with v1.0 near the compute-accuracy frontier.
+We validate our design choices on the Austrian Crop classification task.
+Table 3 summarizes the results; in the following we organize the findings with brief Q&A headings and defer all methodological details to the supplementary material.
+Table 3: Ablation on the Austrian Crop dataset. We report Validation F1 and RankMe (higher $\uparrow$ is better). Parentheses show change vs. the Baseline.
+Configuration Val. F1 ($\uparrow$) RankMe ($\uparrow$)
+Baseline (Full Model) 77.3 0.963
+w/o Global Data Shuffling 68.1 (-9.2) 0.847 (-0.116)
+w/o Mixup Regularization 66.2 (-11.1) 0.857 (-0.106)
+w/o Sentinel-1 Data 74.2 (-3.1) 0.931 (-0.032)
+w/o Shuffling & Mixup 62.6 (-14.7) 0.867 (-0.096)
+w/o Embedding Quantization 77.9 (+0.6) 0.972 (+0.009)
+w/o Pretraining 43.8(-33.5) -
+Do shuffling and mixup help?
+Yes.
+Removing global shuffling or mixup drops validation F1 by 9.2-9.2/11.1-11.1 and RankMe score by 0.116-0.116/0.106-0.106.
+Removing both leads to the largest fall (14.7-14.7 F1; see Table 3).
+The two strategies are complementary: shuffling smoothes the loss curve, mitigates spatial autocorrelation while mixup helps to prevent feature overfitting [7].
+How robust are the embeddings to cloudiness?
+We find that when the annual number of non-cloudy valid Sentinel-2 observations drops $\leq\!10$, Macro-F1 performance drops sharply (see Figure 6).
+At a global scale, we find that embeddings are quite robust to cloudiness since there are more than 70 observations per year for each orbit.
+Is fine-tuning ec weights necessary?
+Self-supervised pretraining of the encoder yields large performance gains compared to randomly initialized weights, a +33.5 F1 improvement in the Austrian Crop classification task.
+Pretraining, somewhat surprisingly, makes subsequent finetuning unnecessary: Figure 6 (b) shows that finetuning the encoder provides no benefit.
+This suggests that the embeddings encapsulate a highly generalized representation.
+Are regional embeddings necessary?
+We find that retraining with region-specific data offers negligible gains (Figure 6 c).
+A single global model is sufficient for high performance across diverse geographies, reinforcing the Embeddings-as-Data paradigm and making costly regional retraining unnecessary.
+Can a small spatial head recover spatial context?
+Yes.
+Although pretraining is pixel-wise (with no explicit spatial structure), adding a very small spatial head (e.g., a small UNet) consistently improves segmentation and regression tasks, with larger gains at higher label budgets.
+This suggests that spatial cues can be injected downstream efficiently without complicating pretraining or creating purely pixel-based embeddings.
+How sensitive are embeddings to random temporal sampling?
+Not much.
+With L=40L{=}40 temporal samples, repeated random draws produce tightly clustered results (low variance), consistent with our time-sampling invariance objective (shown in Figure 6 c).
+We also compare $L\!\in\!\{20,40,96,365\}$ and find that L=40L{=}40 is the accuracy–efficiency sweet spot (see supplementary material).
+Scaling laws.
+Figure 7 shows that the performance scales predictably with the depth/width of the projector, FLOPs, and the parameters; TESSERA v1.0 (star) lies near the optimal compute–accuracy frontier, while the dagger (†) marks the maximum-F1 configuration constrained by memory/batch size.
+Flexibility of the inference time window.
+In the default setting, TESSERA is trained using a one-year temporal window.
+However, this one-year window is not a strict requirement.
+We find that applying the model to shorter seasonal windows during inference (e.g., only from spring to autumn) still yields effective embeddings (see supplementary materials for seasonal window analyses on the Austrian crop dataset).
+Does int8 quantization significantly degrade performance?
+No: there is only a small drop in performance.
+Removing quantization slightly improves metrics (e.g., +0.6+0.6 F1 and +0.009+0.009 RankMe in Table 3), indicating that int8 introduces minor information loss.
+However, using int8 compresses the embeddings to $\sim\!25\%$ of the fp32 storage footprint (and reduces I/O), which is attractive for global-scale products and for use on an edge device.
 6 Conclusion
-In this work, we introduce TESSERA, an innovative remote sensing foundation model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data.
-TESSERA embeddings establish new performance standards, while our open-source methodology, FAIR-compliant data distribution through version-controlled registries, standard geospatial formats, and the GeoTessera library permit reproducible science and enable use of our work even by non-specialists who lack access to high-performance computing.
-We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models.
-Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
-5 Acknowledgments
-We gratefully acknowledge support from Mantle Labs and help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
-Funding:
-ZF and SK were funded by a charitable donation made by Dr Robert Sansom to the University of Cambridge.
-SJ was funded by a charitable donation from John Bernstein.
-JB received support from the UKRI CRCRM scheme (grant MR/Y024354).
-SS received funding from the Jenny and Antti Wihuri Foundation and the Helsinki Institute for Information Technology.
-DAC and SK were funded by the NERC Centre for Landscape Regeneration (NE/W00495X/1).
-TJ was funded by NERC grant NE/S010750/1.
-MCL was funded by UKRI and Mantle Labs.
-Author contributions:
-Initial conceptualization: CA
-System design: CA, ZF, SJ, SK
-System implementation: ZF, RY, MCL, AM
-Data collection: TJ
-Data analysis: ZF, RY, JK, SJ, AM, TJ, DAC, SK
-Downstream tasks: MCL, CA, SS, DAC, CA, MI, ZF, JK
-Paper conceptualization: SK, AM, DAC, AB, SJ
-Paper draft: ZF, MCL, RY, SS, SK, JK
-Paper review and refinement: All
-Competing interests:
-There are no competing interests to declare.
-Data and materials availability:
-Our framework integrates openly available data from the Copernicus Sentinel-2 (S2) [99] optical and Sentinel-1 (S1) [100] SAR missions, leveraging the strengths of each sensor.
-We used S2 Level-2A surface reflectance products, providing detailed spectral information across 10 bands (Red, Blue, Green, NIR, NIR8, Red-edge1, Red-edge2, Red-edge3, SWIR16, and SWIR22).
-To remove cloud contamination, we filtered the time series using the associated Sen2Cor scene classification map to mask cloudy observations.
-For SAR data, we processed Sentinel-1 Ground Range Detected imagery to derive calibrated backscatter coefficients in VV and VH polarizations, applying standard corrections using the ESA SNAP toolbox [101].
-The Google Satellite Embedding V1 dataset [39] is licensed under CC-BY 4.0 and is produced by Google and Google DeepMind.
-PRESTO embeddings can be created on demand in Google Earth Engine [102] using VertexAI [103].
-Our global embedding map–initially for 2024 and eventually one for each year from 2017 to 2024–is $\sim$250 TB and is freely available for download, along with the parameters of the TESSERA model, the code to retrieve Sentinel-1 and Sentinel-2 data, and the training and inferencing code [37].
-To allow for the seamless use of our embeddings in existing geospatial workflows, we have released the  geotessera Python library into the PyPI repository [104], which is the official third-party software repository for Python.
-This library automates the retrieval of TESSERA embeddings for a given year and geographical region, merging them into a numpy buffer [105] that can be passed directly into downstream tasks.
-The embeddings are indexed using Pooch [106] with per-tile checksums stored in an open Git repository, allowing users to pin specific versions for reproducibility or automatically refresh to the latest embeddings as needed.
-Additionally, a Parquet index is published daily for rapid spatial queries across the entire dataset.
-The crop type identification downstream task uses the open INVEKOS Austrian crop dataset [44].
-The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset [95].
-The canopy height downstream task uses open access airborne laser scan data collected in the Danum valley and Sepilok [66].
-The AGB downstream task uses the open BioMassters dataset [72].
-The code for the winning model of the BioMassters competition is described in [72] and freely available on the corresponding GitHub page.
-The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark [79].
-The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners; hence, data access will require a Data Transfer Agreement.
-Appendix A Materials and Methods
-Table 1: Variable definitions for remote sensing data structure
-Variable Dimensions Description
-$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
-$V_{t}$ $W\times H$ Binary mask for tile at time $t$ (1=valid, 0=invalid/clouded)
-$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
-$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
-$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
-$P_{i,j}$ $C\times T$ d-pixel: all spectral channels by timesteps at location $(i,j)$
-$m_{i,j}$ $T\times 1$ d-pixel mask indicating valid timesteps at location $(i,j)$
-$W,H$ Scalar Spatial dimensions (width, height)
-$C$ Scalar Number of spectral channels + polarizations
-$T$ Scalar Number of time steps
-A.1 Model details
-Our input data consists of Sentinel-2 optical satellite imagery (Level-2A bottom-of-atmosphere) and Sentinel-1 Synthetic Aperture Radar (SAR, Radiometrically Terrain Corrected).
-Data Representation
-We consider remote sensing data [100, 99] with $C$ channels (spectral bands or polarizations). Each data tile $R_{t}$ at time $t$ is represented as a 3D array with dimensions:
-$R_{t}\in\mathbb{R}^{W\times H\times C}$ (1)
-where $W$ is the width (longitude dimension), $H$ is the height (latitude dimension), and $C$ is the number of spectral channels.
-Each tile is accompanied by a corresponding binary mask $V_{t}$ of dimensions:
-$V_{t}\in\{0,1\}^{W\times H}$ (2)
-where $V_{t}(i,j)=0$ indicates clouding or missing data for the pixel at spatial coordinates $(i,j)$, and $V_{t}(i,j)=1$ indicates valid data.
-Temporal Data Stacking
-We stack spatially aligned tiles over a time period that spans $T$ time steps ($t=0,1,\ldots,T-1$). The temporal data stack is defined as:
-$\mathbf{D}=[R_{0},R_{1},\ldots,R_{T-1}]$ (3)
-$\mathbf{M}=[V_{0},V_{1},\ldots,V_{T-1}]$ (4)
-Time Series Extraction and d-pixel Definition
-For a given spatial location $(i,j)$ and spectral channel $c$, the time series $S_{i,j,c}$ represents all channel $c$ values at coordinates $(i,j)$ over the entire time period:
-$S_{i,j,c}=[R_{0}(i,j,c),R_{1}(i,j,c),\ldots,R_{T-1}(i,j,c)]$ (5)
-We define a d-pixel $P_{i,j}$ as the collection of all spectral channels by timesteps at a given spatial location $(i,j)$:
-$P_{i,j}(c)=S(i,j,c)$ (6)
-In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time.
-Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there are valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
-The d-pixel Representation
-We represent each 10 m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig.
-1 a).
-In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization.
-This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities.
-Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored in the temporal sampling process.
+We introduced TESSERA, a pixel-wise foundation model for Earth Observation that learns robust representations from sparse, multi-modal time series via temporal sampling invariance.
+Extensive experiments show that TESSERA establishes a new state-of-the-art across diverse tasks with unparalleled label efficiency.
+We pioneer the “Embeddings-as-Data” paradigm, releasing analysis-ready global embeddings via the GeoTessera library to lower barriers for geospatial analysis and empower future research.
+Appendix A Supplementary Material
+In this supplementary material, we provide additional details on downstream tasks, model architecture, and training procedures that were omitted from the main paper due to space constraints.
+A.1 Taxonomy of Remote Sensing Foundation Models
+The current landscape of RSFMs, as illustrated in Figure 8, is vibrant and rapidly expanding, primarily along two avenues: Visual Foundation Models (VFMs) and Visual Language Models (VLMs).
+While these models have pushed the boundaries of what is possible in terms of performance, they have largely overlooked the critical dimension of usability, creating a significant bottleneck for the broader scientific community.
+Figure 8: Taxonomy of RSFMs.
+All RSFMs are categorized into two primary classes: visual foundation models and visual language models, which are further subdivided according to different methodologies.
+The final column of the Sankey diagram illustrates the data modalities used by each foundation model.
+Dashed rectangles indicate areas with minimal or virtually no existing research.
+The checkerboard pattern denotes that, in addition to providing model parameters, these foundation models also offer globally generated, analysis-ready representation products derived from the model itself.
+VFMs are engineered to interpret a diverse array of visual remote sensing data.
+Early and influential models like SatMAE [15] and its successors have successfully adapted the masked autoencoder (MAE) [33] paradigm for temporal MSI and RGB imagery, demonstrating strong capabilities in learning spatial patterns.
+For instance, Scale-MAE [71] introduced a scale-aware pre-training strategy to handle multi-scale geospatial data, while SatMAE++ [65] explored multi-scale reconstructions to better capture scale variability in RS images.
+To address the need for multi-modal fusion, models like MMEarth [64] and the recently proposed Galileo [89] have emerged, integrating an extensive range of inputs including MSI, SAR, and DEM through sophisticated fusion frameworks.
+Concurrently, models focused on temporal analysis have also been developed.
+Presto [88], for instance, is a pixel-level Transformer pre-trained on a global time-series dataset, showing excellent performance in tasks requiring temporal understanding.
+However, a unifying weakness across almost all these VFMs is their profound usability challenge.
+They are released as massive models, placing the onus of data acquisition, complex preprocessing, and computationally intensive fine-tuning squarely on the end-user.
+This approach requires users to possess significant GPU resources and deep technical expertise, creating a high barrier to entry.
+While Google’s AlphaEarth makes a notable exception by providing pre-generated representation maps, it is not fully open-source.
+VLMs for remote sensing aim to bridge the gap between visual data and natural language, partially addressing cross-domain integration.
+Models like SatCLIP  [42], RemoteClip  [53], and GeoRSCLIP [120] have adapted contrastive learning (e.g., CLIP [69]) to align satellite images with textual descriptions, enabling powerful zero-shot classification and text-based image retrieval.
+More advanced models like EarthGPT [118] and SkySenseGPT [56] build upon frameworks like LLaVA [54] to facilitate conversational interaction and multi-sensor (e.g., SAR, infrared) image comprehension.
+While these models open up exciting new avenues for human-computer interaction with geospatial data, they inherit the same fundamental usability limitations as VFMs.
+They are typically released as large models requiring expert handling and significant computational power, and none provide large-scale, pre-generated products that would lower the barrier to entry for non-specialist users.
+In summary, the field has made tremendous strides in improving the performance of RSFMs.
+Yet, this progress has been largely confined to a model-centric paradigm, inadvertently creating a high barrier to entry that contradicts the goal of democratizing science.
+The critical gap, therefore, is not a lack of powerful models, but a lack of a user-centric framework that delivers their power in an accessible and ready-to-use format.
+A.2 Model Details
 Dual-Encoder Architecture
@@
 Thus, our augmentations are fundamentally about sampling from the available, inherently incomplete information streams, rather than artificially distorting a complete input.
-The network processes these two views ($Y_{A},Y_{B}$) through the dual-encoder and the projector to produce batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
-The standard Barlow Twins loss function, $\mathcal{L}_{BT}$, is defined as [34]:
-$\mathcal{L}_{BT}=\sum_{i}(1-C_{ii})^{2}+\lambda_{BT}\sum_{i}\sum_{j\neq i}C_{ij}^{2}$ (7)
-Here, $C$ is the cross-correlation matrix computed between the batch-normalized embeddings $Z_{A}$ and $Z_{B}$.
-The indices $i$ and $j$ iterate over the dimensions of the embedding vectors.
-The first term (invariance term) encourages similar embeddings for different views of the same input ($C_{ii}\to 1$).
-The second term (redundancy reduction term) promotes information efficiency by minimizing the correlation between different embedding dimensions ($C_{ij}\to 0$ for $i\neq j$), weighted by $\lambda_{BT}$.
-We used the value of $\lambda_{BT}=5\times 10^{-3}$ as recommended in Reference [34].
-To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al.
-[108].
-For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$.
-The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution on the unit interval, $\alpha_{mix}\sim U(0,1)$.
-The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views.
-The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
-$C_{target}^{MA}$ $=\alpha_{mix}(Z_{A})^{T}Z_{A}+(1-\alpha_{mix})(Z_{S})^{T}Z_{A}$ (8)
-$C_{target}^{MS}$ $=\alpha_{mix}(Z_{A})^{T}Z_{S}+(1-\alpha_{mix})(Z_{S})^{T}Z_{S}$ (9)
-$\mathcal{L}_{MIX}$ $=\frac{1}{2}(\|C^{MA}-C_{target}^{MA}\|_{F}^{2}+\|C^{MS}-C_{target}^{MS}\|_{F}^{2})$ (10)
-where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model output. The total loss function optimized during training is a weighted sum:
-$\mathcal{L}_{total}=\mathcal{L}_{BT}+\lambda_{mix}\mathcal{L}_{MIX}$ (11)
-where $\lambda_{mix}$ controls the strength of the mix-up regularization. We found $\lambda_{mix}=1.0$ (as recommended by [108]) to be effective based on our experiments.
+A.3 Pre-training Details
 The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles chosen from the years 2017-2024.
@@
 Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig.
-9.
+10.
 A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
@@
 This pipeline is conceptually illustrated in Fig.
-10a.
+11a.
 The impact of this approach on training stability is empirically demonstrated in Fig.
-10b.
+11b.
 Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
@@
 Following this global shuffle, the data augmentation required by the Barlow Twins framework is applied.
-As detailed in the Supplementary Text, this involves generating two distinct augmented views (e.g., $Y^{A}$ and $Y^{B}$) for each d-pixel.
 To manage the significant I/O demands of shuffling and augmenting such a large volume of data, we developed a custom high-performance pipeline that handles the reading of raw d-pixel data, executes the global shuffle, and prepares the data for augmentation.
@@
 D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool.
-A custom Rust binary [37] performs a global shuffle on this multi-terabyte dataset before applying augmentations.
+A custom Rust binary performs a global shuffle on this multi-terabyte dataset before applying augmentations.
 The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches.
@@
 The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
-Global Embedding Map Generation
+A.4 Global Embedding Map Generation
 A primary output of the TESSERA project is the generation of annual global embedding maps with 10 meter resolution for the years 2017-2024.
@@
 This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10 meter grid.
-A.2 Scaling with data and network size
+A.5 Scaling with data and network size
 To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task.
 Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
-11).
+7).
 Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
@@
 We varied the number of these hidden layers (referred to as ’depth’ in Fig.
-11a) and their width (number of neurons per layer), with results shown in Fig.
-11a.
+7a) and their width (number of neurons per layer), with results shown in Fig.
+7a.
 As discussed by Zbontar et al.
@@
 For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
-Figure 11: Model performance improves with increasing projector size and additional computation.
-a, Hyperparameter sweep for the projector network’s architecture.
-Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis).
-The colour and size of the point indicate the validation F1 score and the number of parameters, respectively.
-Performance generally increases with both width and depth.
-The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints.
-The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency.
-b, Scaling laws for TESSERA.
-The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters).
-A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used.
-The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
 Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
-11b.
+7b.
 The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
@@
 The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
-Figure 12: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications.
-Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks.
-Users can then fine-tune the entire model or only the head.
-Another approach (b) is where both the encoder and the projector (or decoder) are typically updated.
-In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input.
-This results in lightweight downstream models that consist only of a task-specific head.
-Ablation study
-Table 3: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance.
-We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset.
-The baseline represents the full model.
-Each subsequent row shows the result of removing the specified component.
-Performance is measured by validation F1 score and the RankMe score.
-Arrows ($\uparrow$) indicate that higher values are better.
-The results demonstrate that all components contribute positively to the model’s overall performance.
-Configuration Validation F1 Score ($\uparrow$) RankMe Score ($\uparrow$)
-Baseline (Full Model) 0.773 0.963
-w/o Global Data Shuffling 0.681 0.847
-w/o Mixup Regularization 0.662 0.857
-w/o Sentinel-1 Data 0.742 0.931
-w/o Shuffling & Mixup 0.626 0.867
-w/o Embedding Quantization 0.779 0.972
+A.5.1 Ablation study details
 To validate the contribution of each key component within the TESSERA framework, we performed an ablation study.
@@
 A score closer to 1 indicates that the embeddings are of higher effective rank, utilizing the full dimensionality of the feature space more effectively to create richer, less redundant representations.
-The results of the ablation study are summarized in Table 3.
 Our analysis reveals several key insights into the model’s architecture and training strategy.
@@
 This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
-A.3 Downstream Task Application Methodology
+A.5.2 Justification for the Number of Sampled Timesteps ($L$)
+Table 5: Parameter sweep of the number of sampled timesteps ($L$). We evaluate the trade-off between F1-score on the Austrian crop classification dataset and inference speed.
+$L$
+Inference Speed
+(samples/sec)
+Encoder Param. Training Samples F1 Score (%)
+15 1509 $\sim$36.4M 22M 65.04
+20 1452 66.58
+30 1276 71.43
+40 1068 73.34
+50 724 73.69
+60 497 73.77
+As detailed in the main paper, TESSERA is trained using sparse random temporal sampling to enforce invariance to the selection of valid observations.
+A critical hyperparameter in this process is $L$, the number of valid timesteps sampled.
+We finalized our choice at $L=40$.
+This decision is supported by two primary justifications:
+  1. 1.
+    Performance vs.
+Efficiency Trade-off: We conducted a comprehensive parameter sweep on the Austrian crop classification dataset to determine the optimal balance between model performance and computational cost (i.e., inference speed).
+As demonstrated in Table 5, setting $L=40$ achieves the best trade-off.
+While further increasing $L$ (e.g., to 50 or 60) results in only marginal or negligible performance gains, it increases the computational overhead during inference.
+  2. 2.
+    Global Data Representativeness: Our choice is also empirically grounded in global-scale data statistics.
+We analyzed the global distribution of valid Sentinel-2 observations acquired between 2017 and 2024 and found that the global average number of valid observations per pixel is approximately 37.6.
+This confirms that $L=40$ is not only efficient but also a robust and representative value for global-scale modeling, closely aligning with the average valid Sentinel-2 data availability worldwide (see Figure 12).
+Figure 12: A distribution map of the number of valid Sentinel-2 observations from 2017 to 2024. Valid observations are defined as those that are free of clouds, water bodies, and snow.
+Furthermore, we observed that downstream task performance is largely insensitive to the number of Sentinel-1 observations sampled (as illustrated in Figure LABEL:fig:ablation_s1_timesteps), likely due to the higher temporal density and all-weather capabilities of SAR.
+Therefore, for the sake of architectural consistency and to simplify the multi-modal training pipeline, we also set the number of sampled timesteps for Sentinel-1 to $L=40$.
+A.6 Downstream Task Application Methodology
+A.6.1 TESSERA
 A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data.
@@
 The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
-The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. 12):
   1. 1.
-    Download Embeddings for the Region of Interest: The Geotessera Python library [37] allows users to download embeddings for a desired region and year in the form of a numpy array.
+    Download Embeddings for the Region of Interest: The GeoTessera Python library allows users to download embeddings for a desired region and year in the form of a numpy array.
   2. 2.
@@
 To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
-A.4 Downstream Task: Austrian Crop Classification
+A.6.2 Presto & AlphaEarth
+For Presto and AlphaEarth, we adopted a methodology conceptually similar to TESSERA’s, focusing on evaluating their publicly available embeddings.
+For Presto, we used the official Google Earth Engine (GEE) scripts provided by the authors to generate embeddings for the regions of interest.
+The data preprocessing was executed on GEE, and inference was subsequently performed on Google Vertex AI, loading the officially released pre-trained model weights.
+For AlphaEarth, we directly used the publicly available Google AlphaEarth Satellite Embedding V1 dataset [11].
+In both cases, the pre-computed embeddings served as the input features.
+The visualization of embeddings for some downstream tasks of TESSERA, AlphaEarth, and Presto are shown in Figure 13.
+We then designed and trained lightweight task-specific heads (e.g., MLPs) for each downstream task, following the same principle as described for TESSERA.
+Figure 13: The visualization of embeddings for some downstream tasks of TESSERA, AlphaEarth, and Presto
+A.6.3 Other RSFMs
+To evaluate the performance of other RSFMs, we used the standardized Pangaea evaluation framework.
+This framework provides a consistent protocol for benchmarking RSFMs across a suite of diverse downstream tasks.
+We evaluated these models on the Austrian Crop (Segmentation), PASTIS-R, Biomassters, and Borneo CHM datasets.
+Unlike our embedding-focused approach for TESSERA, the Pangaea protocol typically involves training task-specific decoders on top of the pre-trained encoder features.
+For classification tasks, a simple LinearClassifier head was attached and trained.
+For the regression and segmentation tasks, heavier decoders such as UperNet+LTAE or a baseline Unet were employed, as specified in the benchmark’s guidelines.
+The exact downstream task heads used for these models are detailed in Table 6.
+Table 6: Downstream task head specification.
+This table lists the decoders (heads) used for different tasks across all models.
+Note that for each model listed, the Regression and Segmentation tasks use the same head.
+Model Classification Head Regression & Segmentation Head
+CROMA LinearClassifier UperNet+LTAE
+DOFA LinearClassifier UperNet+LTAE
+Prithvi LinearClassifier UperNet+LTAE
+RemoteCLIP LinearClassifier UperNet+LTAE
+SatlasNet LinearClassifier UperNet+LTAE
+Scale-MAE LinearClassifier UperNet+LTAE
+SpectralGPT LinearClassifier UperNet+LTAE
+Galileo LinearClassifier UperNet+LTAE
+Skysense LinearClassifier UperNet+LTAE
+ViT Baseline LinearClassifier UperNet+LTAE
+UNet Baseline LinearClassifier Unet
+Presto LinearClassifier Unet
+AlphaEarth LinearClassifier Unet
+TESSERA LinearClassifier Unet
+A.7 Downstream Task: Austrian Crop Classification
 Dataset and Preprocessing
@@
 The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation.
-For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
+When constructing the pixel-level classification dataset, we use fields as the unit for data splitting.
+Each field is assigned a unique identifier along with its corresponding area.
+We allocate pixels from fields constituting X% of the total area to the training set, while the remaining (1-X)% of the total field area is divided such that 1/7 is used for validation and the rest for testing.
+For training, validation, and test splits, we ensure that all 17 crop types are represented in each split.
+For the patch-based dataset designed for segmentation tasks, we first divide the region of interest vertically into five equal parts.
+From the topmost section, we extract 1000 patches of 640m × 640m as the test set.
+An equal number of patches with the same dimensions are cropped from the section immediately below for the validation set.
+From the remaining three sections, 3000 patches of the same size are sampled to form the training set.
+All patches maintain a spatial resolution of 10 meters.
 Pixel-wise Classification Baselines
@@
 These embeddings were then fed into an MLP head identical to the one used for TESSERA.
-  • GSE + MLP: To include another key foundation model, we used the publicly available Google AphaEarth Satellite Embedding (GSE) V1 dataset [39].
+  • AlphaEarth + MLP: To include another key foundation model, we used the publicly available Google AphaEarth Satellite Embedding V1 dataset [11].
 These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
   • Random Forest: As a traditional baseline, we trained a Random Forest (RF) classifier directly on raw time-series data.
-For each pixel, we utilized all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year.
+For each pixel, we used all available Sentinel-1 (2 polarizations) and Sentinel-2 (10 spectral bands) observations throughout the year.
 The temporal and spectral/polarization dimensions were flattened and concatenated to form a single 1256-dimensional feature vector.
@@
 To assess spatial-contextual performance, we conducted a semantic segmentation experiment. The approach varied based on the foundation model’s architecture:
-  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [19], and GSE [39], a distinct workflow was used.
+  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, Presto [88], and AlphaEarth [11], a distinct workflow was used.
 We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$).
 These patches were then fed into a standard UNet architecture [73], which learns to model the spatial relationships between the embeddings to produce a segmentation map.
-It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do.
+It is noteworthy that while TESSERA and Presto’s base embeddings do not explicitly model spatial context, AlphaEarth’s do.
 However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
@@
 The first column shows the ground truth segmentation.
-For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
+For each model (TESSERA, Presto, and AlphaEarth), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
 Embedding Space Analysis
-The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [45] algorithm to the embeddings of every pixel within the Austrian study area.
+Figure 15: UMAP visualization of the embedding spaces for TESSERA, Presto, and AlphaEarth for 17 crop classes. TESSERA’s embeddings exhibit clearer separation and more coherent clustering visually.
+The 2D visualizations shown in Figure 15 were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [60] algorithm to the embeddings of every pixel within the Austrian study area.
 UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology.
@@
 UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space.
-In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
+In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, Presto, and AlphaEarth embeddings.
 To quantitatively measure the quality of the clustering in the original 128-dimensional space, we calculated the Silhouette score and the Davies-Bouldin Index.
@@
 Lower DBI values indicate better clustering, with a score of 0 representing the ideal case where the clusters are compact and well separated.
-A.5 Downstream Task: Canopy height estimation
+Impact of Seasonal Windows
+We analyze the impact of using different temporal windows on the model’s performance for the downstream task of crop classification on the Austrian dataset.
+As shown in Table 8, we conduct an ablation by systematically excluding one of the four meteorological seasons (Spring, Summer, Autumn, Winter) from the input time series.
+The results reveal that, for this specific regional dataset, excluding time steps from the winter season improves the classification F1 score from 78.94% (all seasons included) to $\mathbf{79.83\%}$.
+This suggests that the dormant winter period, which is typically characterized by low vegetation phenological activity and is prone to cloud cover, snow, or reduced solar zenith angle effects, introduces a substantial amount of non-discriminative noise.
+By excluding the winter data, the model can focus its attention and resources on the most informative growing seasons, which allows for the learning of a more concise and effective temporal representation for crop type identification.
+Table 8: Ablation on the Austrian Crop dataset: Seasonal Data Inclusion.
+We report the Validation F1 score (higher $\uparrow$ is better) for the crop classification task.
+The columns for Spring to Winter use  to indicate inclusion and  to indicate exclusion.
+Parentheses show change vs.
+the Full Model.
+Configuration Spring Summer Autumn Winter F1 ($\uparrow$)
+Full Model (Baseline) 78.94
+w/o Spring 78.62 (-0.32)
+w/o Summer 78.10 (-0.84)
+w/o Autumn 78.22 (-0.72)
+w/o Winter 79.83 (+0.89)
+A.8 Downstream Task: Canopy height estimation
 Study Areas and Ground Truth Datasets
@@
 Baselines
-TESSERA, PRESTO [19], and GSE [39] embeddings were compared by training identical models using each representation as input, with all other settings held constant.
+TESSERA, Presto [88], and AlphaEarth [11] embeddings were compared by training identical models using each representation as input, with all other settings held constant.
 Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
-To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al.
-[2021][59] (GLAD), Lang et al.
-[2023][58] (ETH), and Tolan et al.
-[2024][60] (Meta) as well as a regional map produced by Lang et al.
-[2021][68].
+To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against a global canopy height product: Tolan et al.
+[2024][87] (Meta).
 These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS.
 Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
+Figure 16: Comparison of canopy height estimation performance across four models.
+Predicted canopy height (y-axis) is compared against airborne LiDAR reference measurements (x-axis) in Danum Valley, Borneo.
+The figure shows scatter-density plots for four model variants trained using different input representations: TESSERA, Presto, AlphaEarth, and the DINOv2 Global CHM product (renamed from the original Meta Global CHM).
+All models were evaluated using the same LiDAR reference dataset for consistency.
+TESSERA achieves the highest coefficient of determination and lowest RMSE, demonstrating superior performance in structurally complex tropical canopies.
+In contrast, the DINOv2 Global CHM product exhibits strong underestimation, large negative bias, and a near-zero or negative correlation, reflecting substantial saturation at lower canopy heights.
+A 1:1 reference line (blue) is included in each subplot for comparison.
 Finally, we trained the U-Net model using Sentinel-2 input as a baseline.
@@
 Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
-A.6 Downstream Task: Burned area detection
-Study Areas and Ground Truth Datasets
-For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [114], referred to as Burn Area A and Burn Area B.
-Each Burn Area contains over two million 1 m pixels.
-Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure.
-In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha).
-Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land.
-In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland.
-To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [115].
-The dataset from the Monitoring Trends in Burn Severity (MTBS) program [95] was used to identify burned areas and severity of the burn.
-Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset.
-Burn severity, defined as the measure of ecological damage caused by fire [116], was derived from categorical and continuous severity metrics provided by MTBS.
-To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire.
-MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images.
-Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [116].
-Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations.
-The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [102].
-For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high).
-The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
-Dimensionality Reduction
-We used UMAP (Uniform Manifold Approximation and Projection) [45] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires.
-We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections.
-To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs.
-unburned), land cover class, or burn severity.
-These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest.
-To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction.
-UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space.
-UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
-Label Efficiency Analysis and Baselines
-We conducted a label efficiency experiment to evaluate the performance of a random forest classifier when only a limited amount of labelled data is available.
-The valid pixels from both study areas were combined and split into a fixed test set (20%) and a training set.
-We then trained multiple classifiers on progressively smaller stratified random subsamples of the full training set, with ratios ranging from 1.0 to 0.001%.
-The classifier used 100 estimators, a maximum depth of 12, minimum leaf samples of 20, and class_weight=‘balanced’ to account for class imbalance.
-Performance was measured using the macro-averaged F1 score between burned and unburned classes on the fixed test set.
-We compared three types of input embeddings for the classification task.
-The first was TESSERA, using embeddings for 2020.
-The second was a task-specific composite baseline, constructed using prior knowledge of the fire event.
-This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector.
-The third type of input was Google AlphaEarth Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [39].
-GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
-A.7 Downstream Task: Above-ground biomass estimation
+A.9 Downstream Task: Above-ground biomass estimation
 Dataset
@@
 Patch coordinates and collection years–provided by the authors of Ref.
-[72] upon request–were used to retrieve satellite imagery to generate TESSERA and Google AlphaEarth Satellite Embeddings (GSE) [39].
+[63] upon request–were used to retrieve satellite imagery to generate TESSERA and Google AlphaEarth Satellite Embeddings (AlphaEarth) [11].
 UNet model architecture
-To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [76], using TESSERA or GSE embeddings  [39] as input.
-The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
+To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [73], using TESSERA or AlphaEarth embeddings  [11] as input.
+The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and AlphaEarth, respectively, where the dimension denotes the number of channels.
 The network followed a standard UNet structure with two encoder and decoder stages.
@@
 Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
-Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
+Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the AlphaEarth embeddings.
 Baselines
-In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [72] and the SpectralGPT foundation model [18] combined with a UPerNet [78].
+In addition to AlphaEarth, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [63] and the SpectralGPT foundation model [34] combined with a UPerNet [108].
 For SpectralGPT, we used the implementation provided in Ref.
@@
 According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
-A.8 Downstream Task: Stocking indices in Voluntary Carbon Markets
-Relative height modelling
-Participants in the voluntary carbon market can use a remotely sensed “stocking index” to quantify project additionality, that is, the degree to which the carbon sequestration in the project exceeds what would have occurred in the natural course of events.
-The chosen stocking index must be correlated with the above-ground biomass (AGB) of the observed objects, and the strength of the correlation must be demonstrated using a suitable in situ reference dataset.
-To demonstrate the superior quality of the TESSERA embeddings compared to existing state-of-the-art canopy height or AGB products, a simple stocking index was derived from GEDI relative heights and TESSERA embeddings.
-This permits comparing state-of-the-art models often used in the voluntary carbon market with a TESSERA-derived stocking index.
-To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [53] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil.
-From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables.
-A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index.
-The TerraClass Amazonia 2022 land cover map [117] was used to distinguish between forest and non-forest.
-To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
-Stocking index baselines
-For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC).
-Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA.
-For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [58], CTrees canopy tree height (5 m) [118], ESA-CCI AGB (100 m) [119], Meta CHM (1 m)  [60], and GTDX AGB (25 m) [120].
-Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
-To ensure a fair comparison between canopy height, AGB and relative height products, the extracted averages were linearly rescaled to best correlate with in situ (AGB) data.
-After rescaling, all products are in units of t/ha and have an intercept of zero and a slope of one with respect to the in situ data, making their RMSE (and $R^{2}$) directly comparable.
-The rescaling is as follows:
-  1. 1.
-    TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 $\times$ + 160.5
-  2. 2.
-[…]
-  3. 3.
-    CTrees canopy height converted using 1.806 $\times$ + 44.9
-  4. 4.
-    ESA AGB converted using 0.407 $\times$ + 34.0
-  5. 5.
-    Meta canopy height converted using 3.723 $\times$ + 46.6
-  6. 6.
-    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
-A.9 Interactive Habitat Mapping Tool
+A.10 Interactive Habitat Mapping Tool
 Figure 17: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
@@
 We also provide a batch processing tool for larger regions of interest.
-The tools and their open source code are freely available in our github repository [37].
+The tools and their open source code are freely available in the github repository.
+A.11 Pseudo Code
+Algorithm 1 Mixup Consistency Regularization for Multimodal Barlow Twins
+1:def MixupBT(s2_aug1, s2_aug2, s1_aug1, s1_aug2, model, barlow_lambda, lambd_mix):
+2:# s2_* and s1_* are batches of augmented inputs
+3:# model returns projection z and representation repr
+4:# standard Barlow Twins branch
+5:z1, repr1 = model(s2_aug1, s1_aug1)
+6:z2, repr2 = model(s2_aug2, s1_aug2)
+7:loss_bt, bar_bt, off_bt = BarlowTwins(z1, z2)
+8:# mixup branch
+9:B = s2_aug1.shape[0]
+10:idx = randperm(B)
+11:alpha = Beta(beta_alpha, beta_beta).sample()
+12:# mix S2 and S1 views
+13:y_m_s2 = alpha * s2_aug1 + (1 - alpha) * s2_aug2[idx]
+14:y_m_s1 = alpha * s1_aug1 + (1 - alpha) * s1_aug2[idx]
+15:z_m, repr_m = model(y_m_s2, y_m_s1)
+16:# cross-correlation of mixed and original views
+17:cc_m_a = Corr(z_m, z1)
+18:cc_m_b = Corr(z_m, z2)
+19:cc_11 = Corr(z1, z1)
+20:cc_2idx1 = Corr(z2[idx], z1)
+21:cc_12 = Corr(z1, z2)
+22:cc_2idx2 = Corr(z2[idx], z2)
+23:# mixup targets and consistency loss
+24:cc_m_a_gt = alpha * cc_11 + (1 - alpha) * cc_2idx1
+25:cc_m_b_gt = alpha * cc_12 + (1 - alpha) * cc_2idx2
+26:diff_a = (cc_m_a - cc_m_a_gt).pow(2).sum()
+27:diff_b = (cc_m_b - cc_m_b_gt).pow(2).sum()
+28:loss_mix = lambd_mix * barlow_lambda * (diff_a + diff_b)
+29:# final SSL loss with mixup consistency
+30:loss = loss_bt + loss_mix
+31:return loss
+Algorithm 2 Global Shuffling and Multimodal Data Construction
+1:def GlobalShuffling(tile_roots, batch_size, T):
+2:# tile_roots: list of paths to MGRS tile directories
+3:# batch_size: number of tiles to process in memory simultaneously
+4:# T: number of time steps for sparse sampling
+5:tiles = shuffle(tile_roots) # Randomize tile processing order
+6:num_batches = len(tiles) // batch_size
+7:for b in range(num_batches):
+8:    current_tiles = tiles[b*batch_size : (b+1)*batch_size]
+9:    pixel_buffer_aug1 = []
+10:    pixel_buffer_aug2 = []
+11:    # Parallel processing of tiles to extract valid pixels
+12:    parfor tile in current_tiles:
+13:      s2, s1, masks = load_data(tile)
+14:      valid_locs = filter_valid_pixels(masks, s1, s2)
+15:      if len(valid_locs) == 0: continue
+16:      # Sparse random temporal sampling per pixel
+17:      for (y, x) in valid_locs:
+18:        idx = get_valid_time_indices(masks, y, x)
+19:        v1 = Sample(s2, s1, idx, T) # View 1
+20:        v2 = Sample(s2, s1, idx, T) # View 2
+21:        pixel_buffer_aug1.append(v1)
+22:        pixel_buffer_aug2.append(v2)
+23:    # Aggregate pixels from spatially disjoint tiles
+24:    stack_aug1 = Stack(pixel_buffer_aug1) # Shape: (N_total, T, C)
+25:    stack_aug2 = Stack(pixel_buffer_aug2)
+26:    # Global Shuffling: Decorrelate spatial neighborhoods
+27:    N_total = stack_aug1.shape[0]
+28:    perm_idx = randperm(N_total)
+29:    shuffled_aug1 = stack_aug1[perm_idx]
+30:    shuffled_aug2 = stack_aug2[perm_idx]
+31:    # Save as chunks for pre-training dataloader
+32:    save_chunks(shuffled_aug1, shuffled_aug2, out_dir)
 Table 9: TESSERA provides unprecedented ease of use, scale, and accuracy.
@@
 It uses spectral-temporal features (ST) at pixel level (P), is invariant to clouds (CI), provides a compressed representation (CR), is oriented to land features (LO), and includes both multispectral and radar information (MS).
-Codes: $\checkmark$= Yes, $\times$= No, ?
+Codes: Y= Yes, N= No, ?
 = Unknown, P= Potentially true, ..
@@
 Foundation Model R O G A 10 m ST P CI CR LO MS
-TESSERA (2025) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-GeoKR (2021) [121] $\times$ ? $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-CVPRW (2021) [122] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
-GASSL (2021) [23] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ .. $\checkmark$ $\times$
-SeCo (2021) [123] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-MOSAIKS (2021) [124] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ ?
-DINO-MM (2022) [125] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ .. $\checkmark$ $\checkmark$
-SatMAE (2022) [16] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ p .. p $\times$
-RS-BYOL (2022) [126] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ .. $\times$ $\checkmark$
-GeCo (2022) [127] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. p $\times$
-RingMo (2022) [128] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-RVSA (2022) [129] $\times$ .. .. $\times$ .. $\times$ $\times$ $\times$ .. $\checkmark$ $\times$
-RSP (2022) [14] $\times$ .. .. $\times$ p $\times$ $\times$ p $\checkmark$ $\checkmark$ $\times$
-MATTER (2022) [130] $\times$ .. .. $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-CSPT (2022) [131] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-CVPRW (2022) [132] $\times$ .. .. $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-BFM (2023) [133] $\times$ .. .. $\times$ ? $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-TOV (2023) [134] ? .. $\times$ $\times$ $\checkmark$ $\times$ ? $\times$ ? $\checkmark$ $\times$
-CMID (2023) [135] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ ? $\checkmark$ $\times$
-RingMo-Sense (2023) [136] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ ? $\checkmark$ $\checkmark$
-lal-SimCLR (2023) [137] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-CACo (2023) [138] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SatLas (2023) [15] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-GFM (2023) [139] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Scale-MAE (2023) [140] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-DINO-MC (2023) [141] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-CROMA (2023) [142] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Cross-Scale MAE (2023) [143] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-DeCUR (2023) [144] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
-Presto (2023) [19] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-CtxMIM (2023) [145] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-FG-MAE (2023) [146] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Prithvi (2023) [17] $\times$ .. $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-RingMo-Lite (2023) [147] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-IGARSS (2023) [148] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-EarthPTi (2023) [149] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-USat (2023) [150] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-AIEarth (2023) [151] p ? $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Clay (2023) [152] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Hydro (2023) [153] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-U-Barn (2023) [154] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
-GeRSP (2023) [155] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SwiMDiff (2023) [156] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-OFA-Net (2023) [157] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-SML-FR (2023) [158] p p $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
-Spectral-GP (2024) [18] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-S2MAE (2024) [159] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SatMAE++ (2024) [160] $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-msGFM (2024) [161] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-SkySense (2024) [24] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-MTP (2024) [162] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-DOFA (2024) [163] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-MMEarth (2024) [164] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-LeMeViT (2024) [165] $\times$ .. $\times$ $\times$ ? $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
-SoftCon (2024) [166] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-RS-DFM (2024) [167] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
-A2MAE (2024) [168] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-HyperSIGM (2024) [169] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SelectiveMAE (2024) [170] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-OmniSat (2024) [171] $\times$ .. $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
-MM-VSE (2024) [172] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$
-MA3E (2024) [173] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SpectralEarth (2024) [174] $\times$ .. $\checkmark$ $\checkmark$ $\times$ ? $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SenPa-MAE (2024) [175] $\times$ .. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-RingMo-Ae (2024) [176] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SAR-JEPA (2024) [177] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
-PIS (2024) [178] $\times$ ? $\checkmark$ $\checkmark$ $\checkmark$ .. .. .. .. .. ..
-OReole-FM (2024) [179] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-PIEVIT (2024) [180] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SatVisionTOA (2024) [181] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
-RS-vHeat (2024) [182] $\times$ .. ? ? ? ? $\checkmark$ ? $\checkmark$ ? $\checkmark$
-Prithvi-EO-2.0 (2024) [110] $\times$ .. $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
-AnySat (2024) [183] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-WildSAT (2024) [184] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-SeaMo (2024) [185] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-FoMo (2025) [186] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-SatMamba (2025) [187] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Galileo (2025) [188] p p $\checkmark$ ? $\checkmark$ $\times$ ? ? $\checkmark$ $\checkmark$ $\checkmark$
-SatDiFuser (2025)[189] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-RoMA (2025) [190] $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Panopticon (2025) [191] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-FedSense (2025) [192] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
-Copernicus-FM (2025) [193] $\times$ .. $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-HyperFree (2025) [194] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
-DynamicVis (2025) [195] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-FlexiMo (2025) [196] $\times$ .. $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
-Google Satellite (2025)[39] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-SkySense++ (2025) [197] $\times$ $\checkmark$ $\checkmark$ p p $\checkmark$ $\times$ p $\checkmark$ $\checkmark$ $\checkmark$
-RingMoE (2025) [198] $\times$ .. $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+TESSERA (2025) Y Y Y Y Y Y Y Y Y Y Y
+GeoKR (2021) [44] N ? N N Y N Y N N Y N
+CVPRW (2021) [80] N .. N N Y N N N .. Y N
+GASSL (2021) [6] N .. N N Y N Y N .. Y N
+SeCo (2021) [59] N N N N Y N N N ? Y N
+MOSAIKS (2021) [72] N Y Y N N N N N Y Y ?
+DINO-MM (2022) [98] N .. N N Y N N N .. Y Y
+SatMAE (2022) [15] N .. N N Y Y N P .. P N
+RS-BYOL (2022) [39] N .. .. N Y N Y Y .. N Y
+GeCo (2022) [45] N .. .. N .. N N N .. P N
+RingMo (2022) [81] N .. .. N Y N N N ? Y N
+RVSA (2022) [93] N .. .. N .. N N N .. Y N
+RSP (2022) [92] N .. .. N P N N P Y Y N
+MATTER (2022) [2] N .. .. N Y Y N N Y Y N
+CSPT (2022) [117] N .. .. N ? N N N Y Y N
+CVPRW (2022) [75] N .. .. N Y N N N Y Y Y
+BFM (2023) [12] N .. .. N ? N N N ? Y N
+TOV (2023) [85] ? .. N N Y N ? N ? Y N
+CMID (2023) [62] N .. N N Y N N N ? Y N
+RingMo-Sense (2023) [112] N .. N N Y Y N N ? Y Y
+lal-SimCLR (2023) [67] N .. N N Y N N N Y Y Y
+CACo (2023) [57] N .. N N Y Y N N Y Y N
+SatLas (2023) [8] N .. Y N Y Y N N Y Y N
+GFM (2023) [61] N .. N N Y N N N N Y N
+Scale-MAE (2023) [71] N .. N N Y N N N N Y N
+DINO-MC (2023) [105] N .. N N Y N N N N Y N
+CROMA (2023) [29] N .. N N Y N N N N Y Y
+Cross-Scale MAE (2023) [84] N .. N N Y N N N N Y N
+DeCUR (2023) [101] N .. N N Y N N Y N Y Y
+Presto (2023) [88] N .. Y Y Y Y Y Y Y Y Y
+CtxMIM (2023) [116] N .. N N Y N N N N Y N
+FG-MAE (2023) [99] N .. N N Y N N N N Y N
+Prithvi (2023) [76] N .. N N N Y N Y Y Y N
+RingMo-Lite (2023) [100] N .. N N Y N N N N Y N
+IGARSS (2023) [27] N .. N N Y N N N N Y N
+EarthPTi (2023) [78] N .. N N Y Y Y Y Y Y N
+USat (2023) [38] N .. N N Y N N N N Y N
+AIEarth (2023) [111] P ? Y N Y N N N N Y N
+Clay (2023) [14] N .. Y N Y N N N N Y N
+Hydro (2023) [17] N .. N N Y N N N N N N
+U-Barn (2023) [24] N .. N Y Y Y N Y N Y N
+GeRSP (2023) [36] N .. N N Y N N N N Y N
+SwiMDiff (2023) [86] N .. N N Y N N N N Y N
+OFA-Net (2023) [110] N .. Y N Y N N N N Y Y
+SML-FR (2023) [21] P P Y N Y N N N N N N
+Spectral-GP (2024) [34] N N Y N Y N N N N Y N
+S2MAE (2024) [47] N N Y N Y N N N N Y N
+SatMAE++ (2024) [65] N N Y N Y N N N N Y N
+msGFM (2024) [32] N N N N Y N N N N Y N
+SkySense (2024) [31] N N N N Y Y N N N Y Y
+MTP (2024) [94] N .. Y N Y N Y N N Y N
+DOFA (2024) [109] N .. Y N Y N N N Y Y Y
+MMEarth (2024) [64] N .. Y N Y N N N Y Y Y
+LeMeViT (2024) [41] N .. N N ? N N N Y N N
+SoftCon (2024) [102] N .. Y N Y N Y N N Y N
+RS-DFM (2024) [104] N .. N N Y N Y N N Y N
+A2MAE (2024) [115] N .. Y N Y N N N Y Y N
+HyperSIGM (2024) [95] N .. Y N Y Y Y N Y Y N
+SelectiveMAE (2024) [96] N .. Y N Y N N N Y Y N
+OmniSat (2024) [4] N .. N Y Y Y N ? Y Y Y
+MM-VSE (2024) [70] N .. Y Y Y Y N ? Y Y Y
+MA3E (2024) [51] N .. N N Y N N N Y Y N
+SpectralEarth (2024) [50] N .. Y Y N ? Y N Y Y N
+SenPa-MAE (2024) [68] N .. Y Y Y Y Y N Y Y N
+RingMo-Ae (2024) [19] N .. N N Y N Y N Y Y N
+SAR-JEPA (2024) [46] N .. N N Y N N Y N N N
+PIS (2024) [3] N ? Y Y Y .. .. .. .. .. ..
+OReole-FM (2024) [20] N .. Y N Y N Y N Y Y N
+PIEVIT (2024) [55] N .. Y N Y N Y N Y Y N
+SatVisionTOA (2024) [79] N .. Y Y N Y Y N Y Y N
+RS-vHeat (2024) [35] N .. ? ? ? ? Y ? Y ? Y
+Prithvi-EO-2.0 (2024) [82] N .. Y Y N Y N Y Y Y N
+AnySat (2024) [5] N .. Y N Y N Y N Y Y Y
+WildSAT (2024) [18] N N N N N N N N Y Y N
+SeaMo (2024) [49] N N N N Y N N N Y Y N
+FoMo (2025) [10] Y Y Y N Y N Y N Y Y Y
+SatMamba (2025) [23] Y Y N N N N N N N Y N
+Galileo (2025) [89] P P Y ? Y N ? ? Y Y Y
+SatDiFuser (2025)[40] N N N N Y N N N N Y N
+RoMA (2025) [97] N N N N Y N N N N Y N
+Panopticon (2025) [91] N Y N N Y Y N N Y Y Y
+FedSense (2025) [83] N N N N N N N N N Y N
+Copernicus-FM (2025) [103] N .. Y N Y Y N N Y Y Y
+HyperFree (2025) [43] N .. N N Y N N N Y Y N
+DynamicVis (2025) [13] N .. N N Y N N N N Y Y
+FlexiMo (2025) [48] N .. N N Y N N N N Y Y
+Google Satellite (2025)[11] Y N Y Y Y Y N Y Y Y Y
+SkySense++ (2025) [106] N Y Y P P Y N P Y Y Y
+RingMoE (2025) [9] N .. N N Y Y N N Y Y Y

### v7

@@
 To democratize access, adhere to FAIR principles, and simplify use, we release global, annual, 10m, pixel-wise int8 embeddings together with open weights/code and lightweight adaptation heads, thus providing practical tooling for large-scale retrieval and inference at planetary scale.
-The model training/inference code, downstream task code, and pre-generated embeddings can be accessed at https://github.com/ucam-eo.
+All code and data are available at https://github.com/ucam-eo/tessera.
 **footnotetext: Corresponding author.
@@
 Second, we introduce two complementary regularizers: global shuffling to disrupt short-range spatial correlations, while mix-based regularization blends observations across time and samples to stabilize learning under severe sparsity.
-Third, unlike nearly all prior work, open-source global, annual, pixel-wise int8 embeddings for each 10m $\times$ 10m location on land, along with open-source model weights, code, and the GeoTessera Python library***https://github.com/ucam-eo/geotessera to efficiently retrieve embeddings from our data store for large-scale inference.
+Third, unlike nearly all prior work, open-source global, annual, pixel-wise int8 embeddings for each 10m $\times$ 10m location on land, along with open-source model weights, code, and the GeoTessera Python library [58] to efficiently retrieve embeddings from our data store for large-scale inference.
 TESSERA embeddings can be paired with lightweight adaptation heads (small MLPs or shallow convolutional models) to achieve high performance with minimal computation, enabling rapid prototyping and straightforward deployment.
@@
 In parallel, early efforts like MOSAIKS [74] explored lightweight universal features to lower adoption barriers.
-Despite these advances, most RSFMs are trained on patch-based inputs and implicitly assume that the data have been preprocessed (e.g., aggressive cloud filtering or temporal averaging).
-Such assumptions simplify pretraining but discard partially cloudy images that still carry useful temporal signals, biasing representations toward idealized, cloud-free conditions.
-Moreover, strong performance typically requires fine-tuning the backbone for each downstream task, which can be compute- and label-intensive for many EO users.
+Despite these advances, most RSFMs are trained on patch-based input and implicitly assume that the data has been preprocessed (e.g., aggressive cloud filtering or temporal averaging).
+Such assumptions simplify pretraining, but discard partially cloudy images that still carry useful temporal signals, biasing representations toward idealized, cloud-free conditions.
+Moreover, strong performance typically requires fine-tuning the backbone for each downstream task, which can be compute- and label-intensive for EO users.
 2.2 Embedding-Based Foundation Models
 A complementary line of work delivers embeddings-as-data: precomputed, generic, task-agnostic features that can be easily adapted with lightweight heads (shown in Figure 2 c).
-Representative models include Presto [88], which learns pixel-level temporal representations, and Google Deepmind’s AlphaEarth [11], which provides precomputed embeddings as a ready-for-analysis product.
+Representative models include Presto [91], which learns pixel-level temporal representations, and Google’s AlphaEarth [13], which provides precomputed embeddings as a ready-for-analysis product.
 This paradigm reduces computational costs and democratizes access to high performance; however, existing approaches are often not fully open-sourced, lack pixel-level granularity at a global scale, or do not explicitly account for irregular temporal sampling.
@@
 TESSERA introduces temporal sampling invariance to handle cloud- and orbital pattern induced irregularity and combines redundancy reduction with global shuffling and mix-based regulation for robustness.
-In contrast to prior embedding releases, we provide FAIR global, annual, 10 m pixel-wise embeddings, together with open weights, code, and GeoTessera python library, enabling analysis-ready use without backbone fine-tuning.
+In contrast to prior embedding releases, we provide FAIR global, annual, 10 m pixel-wise embeddings, together with open weights, code, and the GeoTessera library [58] enabling analysis-ready use without backbone fine-tuning.
 Figure 2: Downstream adaptation paradigms.
@@
 This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing, running the TESSERA model itself (see supplementary materials).
-Inference is performed on 96 Intel Data Center GPU Max 1550s (128 GB GPU RAM each), processing $\sim$5 PB of raw Sentinel data per year from Microsoft Planetary Computer.
-Each year yields $\sim$300 TB of int8-quantized embeddings, with 2024 released first, followed by 2017–2023.
 We release embeddings through the pip-installable GeoTessera Python library, enabling easy access and integration into downstream workflows.
@@
 At a global scale, we find that embeddings are quite robust to cloudiness since there are more than 70 observations per year for each orbit.
-Is fine-tuning ec weights necessary?
+Is fine-tuning weights necessary?
 Self-supervised pretraining of the encoder yields large performance gains compared to randomly initialized weights, a +33.5 F1 improvement in the Austrian Crop classification task.
@@
 A single global model is sufficient for high performance across diverse geographies, reinforcing the Embeddings-as-Data paradigm and making costly regional retraining unnecessary.
-Can a small spatial head recover spatial context?
+Can a spatial head recover spatial context?
 Yes.
-Although pretraining is pixel-wise (with no explicit spatial structure), adding a very small spatial head (e.g., a small UNet) consistently improves segmentation and regression tasks, with larger gains at higher label budgets.
+Although pretraining is pixel-wise (with no explicit spatial structure), adding a spatial head (e.g., a UNet) consistently improves segmentation and regression tasks, with larger gains at higher label budgets.
 This suggests that spatial cues can be injected downstream efficiently without complicating pretraining or creating purely pixel-based embeddings.
@@
 We pioneer the “Embeddings-as-Data” paradigm, releasing analysis-ready global embeddings via the GeoTessera library to lower barriers for geospatial analysis and empower future research.
+7 Acknowledgments
+We thank UKRI (MR/Z505456/1), DAWN, AMD, Vultr, the STFC Durham DiRAC HPC Facility (ST/P002293/1, ST/R002371/1, ST/S002502/1, ST/R000832/1) and the Microsoft AI For Good Lab for resources.
+Donations from Tarides, Jane Street, Google, Dr.
+Robert Sansom and John Bernstein also financially supported this research.
 Appendix A Supplementary Material
@@
 The global batch size was 32,768.
-Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary fig.
-10.
+Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary  Figure 10.
 A crucial aspect of our training methodology is a data shuffling strategy, essential for learning globally representative features from a representative sample of a vast and geographically diverse dataset.
@@
 To overcome this, we developed a custom data processing pipeline to implement a truly global shuffle across all $\approx$800 million training samples, which constituted more than 2 TB of initial d-pixel data.
-This pipeline is conceptually illustrated in Fig.
-11a.
-The impact of this approach on training stability is empirically demonstrated in Fig.
-11b.
+This pipeline is conceptually illustrated in Figure 11 a.
+The impact of this approach on training stability is empirically demonstrated in  Figure 11 b.
 Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot).
@@
 This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10 meter grid.
-A.5 Scaling with data and network size
+A.5 Spatial Context: Pixel-wise vs. Patch-based Approaches
+In this section, we elaborate on the rationale behind TESSERA’s pixel-wise design compared to traditional patch-based architectures.
+By maintaining pixel-wise independence, TESSERA avoids embedding fixed spatial priors, which preserves the integrity of raw spectral signatures and ensures downstream performance is driven by fundamental signal characteristics rather than pre-determined structural assumptions.
+Introducing spatial context via fixed patches (Spatial variant) introduces significant vulnerabilities, particularly in regions with high cloud frequency such as Borneo.
+As the quantitative comparisons demonstrate (Figure 12a), introducing patch-based spatial priors can negatively impact the robustness of the representations.
+Furthermore, filtering for strictly clear spatial patches results in an exponential decrease in data availability, leading to severe embedding artifacts.
+This vulnerability is clearly visible in the embedding visualizations of the Borneo region (Figure 12b), where imposing a spatial patch requirement drastically reduces valid data and creates massive spatial voids.
+By extracting features at the pixel level, TESSERA allows researchers to incorporate domain-specific spatial constraints post hoc without compromising the foundational embeddings.
+Figure 12: Spatial Context Ablation.
+(a) Quantitative comparison detailing performance differences when introducing patch-based spatial priors.
+(b) Visualization of the generated embeddings in the Borneo region.
+A.6 Geographic Generalization
+To demonstrate TESSERA’s ability to generalize to diverse phenological regions and smallholder agricultural settings, we evaluate our model on the CropHarvest Togo dataset [90] and the Ethiopia Crop classification dataset  [10].
+Despite using only a continuous DoY sinusoidal encoding as a weak temporal prior, TESSERA correctly handles local seasonality driven primarily by spectral trends.
+Table 5 compares TESSERA against other baseline foundation models.
+TESSERA significantly outperforms CROMA and AlphaEarth on both benchmarks, demonstrating superior generalization in challenging zero-shot and low-data transfer scenarios characteristic of Sub-Saharan Africa.
+Table 5: Downstream Performance on Sub-Saharan African Benchmarks.
+Evaluation using F1 scores (%) for the Togo (CropHarvest) and Accuary (%) for the Ethiopia Crop dataset.
+TESSERA demonstrates superior adaptability to local seasonalities compared to existing models.
+Model Togo (F1 %) Ethiopia (Acc %)
+CROMA 79.15 /
+AlphaEarth / $\sim$41
+TESSERA (Ours) 83.01 44.92
+A.7 Scaling with data and network size
 To identify an optimal and efficient architecture for TESSERA, and to understand the scaling properties of our model, we conducted a series of experiments using the Austrian crop classification task as the downstream task.
-Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Fig.
-7).
+Our investigation focused on two main areas: the impact of the projector network’s dimensions and the overall model scaling laws concerning performance versus computational cost (Figure 7).
 Our experiments revealed that the performance of the final embeddings on the chosen downstream task of crop classification in Austria was relatively insensitive to the size of the Transformer encoders.
@@
 Our parameter sweep focused on this computationally intensive hidden block.
-We varied the number of these hidden layers (referred to as ’depth’ in Fig.
-7a) and their width (number of neurons per layer), with results shown in Fig.
-7a.
+We varied the number of these hidden layers (referred to as ’depth’ in  Figure 7 a) and their width (number of neurons per layer), with results shown in  Figure 7 a.
 As discussed by Zbontar et al.
@@
 For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
-Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig.
-7b.
+Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in  Figure 7 b.
 The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget.
@@
 The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
-A.5.1 Ablation study details
+A.7.1 Ablation study details
 To validate the contribution of each key component within the TESSERA framework, we performed an ablation study.
@@
 This underlies our ”Embeddings-as-Data” approach, making global-scale high-resolution analysis practical: Without this, each year’s embeddings would consume  1PB, making the hosting of 8 years of embedding costly and likely unachievable.
-A.5.2 Justification for the Number of Sampled Timesteps ($L$)
+A.7.2 Justification for the Number of Sampled Timesteps ($L$)
 Table 6: Parameter sweep of the number of sampled timesteps ($L$). We evaluate the trade-off between F1-score on the Austrian crop classification dataset and inference speed.
@@
 Figure 13: A distribution map of the number of valid Sentinel-2 observations from 2017 to 2024. Valid observations are defined as those that are free of clouds, water bodies, and snow.
-Furthermore, we observed that downstream task performance is largely insensitive to the number of Sentinel-1 observations sampled (as illustrated in Figure LABEL:fig:ablation_s1_timesteps), likely due to the higher temporal density and all-weather capabilities of SAR.
+Furthermore, we observed that downstream task performance is largely insensitive to the number of Sentinel-1 observations sampled, likely due to the higher temporal density and all-weather capabilities of SAR.
 Therefore, for the sake of architectural consistency and to simplify the multi-modal training pipeline, we also set the number of sampled timesteps for Sentinel-1 to $L=40$.
-A.6 Downstream Task Application Methodology
-A.6.1 TESSERA
+A.8 Downstream Task Application Methodology
+A.8.1 TESSERA
 A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data.
@@
 To enhance stability, multiple embeddings can be created with different random temporal samples and averaged.
-A.6.2 Presto & AlphaEarth
+A.8.2 Presto & AlphaEarth
 For Presto and AlphaEarth, we adopted a methodology conceptually similar to TESSERA’s, focusing on evaluating their publicly available embeddings.
@@
 Figure 14: The visualization of embeddings for some downstream tasks of TESSERA, AlphaEarth, and Presto
-A.6.3 Other RSFMs
+A.8.3 Other RSFMs
 To evaluate the performance of other RSFMs, we used the standardized Pangaea evaluation framework.
@@
 TESSERA LinearClassifier Unet
-A.7 Downstream Task: Austrian Crop Classification
+A.9 Downstream Task: Austrian Crop Classification
 Dataset and Preprocessing
@@
 We did not freeze the encoders during the fine-tuning.
-Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 14.
+Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary  Figure 15.
 Figure 15: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation.
@@
 w/o Winter 79.83 (+0.89)
-A.8 Downstream Task: Canopy height estimation
+A.10 Downstream Task: Canopy height estimation
 Study Areas and Ground Truth Datasets
@@
 Predicted canopy height (y-axis) is compared against airborne LiDAR reference measurements (x-axis) in Danum Valley, Borneo.
-The figure shows scatter-density plots for four model variants trained using different input representations: TESSERA, Presto, AlphaEarth, and the DINOv2 Global CHM product (renamed from the original Meta Global CHM).
+The figure shows scatter-density plots for four model variants trained using different input representations: TESSERA, Presto, AlphaEarth, and the DINOv2 Global CHM product.
 All models were evaluated using the same LiDAR reference dataset for consistency.
@@
 Although different Sentinel-2 bands have different spatial resolutions, all were resampled to 10 m to match the FM pipelines.
-A.9 Downstream Task: Above-ground biomass estimation
+A.11 Downstream Task: Above-ground biomass estimation
 Dataset
@@
 The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model.
-Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
+Figure 5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot.
 Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4).
@@
 According to the BioMassters codebase, the original model was trained on two A100 40GB GPUs over the course of eight days, highlighting the greater time efficiency of TESSERA, which required only five hours.
-A.10 Interactive Habitat Mapping Tool
+A.12 Interactive Habitat Mapping Tool
 Figure 18: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes.
@@
 The tools and their open source code are freely available in the github repository.
-A.11 Pseudo Code
+A.13 Pseudo Code
+This section provides the pseudocode for mixup regularization and global shuffling.
 Algorithm 1 Mixup Consistency Regularization for Multimodal Barlow Twins